### v4

# Accounting for multiplicity in machine learning benchmark performance

Accounting for multiplicity in machine learning benchmark performance
Kajsa Møllersen1 and Einar Holsbø2
(1Department of Community Medicine
2Department of Computer Science
UiT - The Arctic University of Norway
1 Introduction: Population vs sample state-of-the-art performance
Machine learning methods are commonly evaluated and compared by their performance on data sets from public repositories, where methods can be evaluated under identical conditions and across time. Theoretical results play an important role in machine learning research, but for the last couple of decades the success of increasingly complex methods has pushed the proof towards the pudding: performance on benchmark data sets.
The numbers don’t lie: if method A performs better than method B on a specific data set, that is what it did. But the truth that the numbers tell is limited. We think of a data set, for example a collection of images, as a random sample from an underlying population. In science we are interested in the performance on the population, and the observed performance acts as an estimate. A good estimator gives a reliable result, but overfitting to the data set, originating from wrong use of cross-validation, data leakage and data reuse, can give overly optimistic estimates. An independent test set, often referred to as a hold-out set, is a sample that is not used for development of the method, and gives an unbiased estimator. Competitions and challenges with prize money typically withhold the test set so that it becomes impossible for the participants to use it for development.
Although we know these things we have a tendency to extrapolate the performance metric and interpret it as something about the world rather than something about a data set. If we always appended “on MNIST” when we said so-and-so method had accuracy so-and-so nobody would care. Who cares about MNIST? We hold out data precisely because we want to estimate some sort of long-term behavior on similar data.
An unbiased estimator solves only half of the problem; the variance tells just as much as the bias regarding the quality of a method. For machine learning methods, the sources of variance are plentiful. One is the random nature of the test set, the training set and the validation set. In addition comes the setting of hyperparameters, which can lead to substantial variation, and minor variance sources such as random seeds and stochastic dropouts. The sources of variance that are part of the training are often referred to as influencing the method’s stability or robustness. With knowledge about the behaviour of both the bias and the variance of an estimator, the error can be predicted, often in terms of a confidence interval. It is certainly more useful to know what the lower limit of the 95% confidence interval is, than the observed performance on a specific test set.
The highest ranked performance on a problem is referred to as state-of-the-art (SOTA) performance, and is used, among other things, as a reference point for publication of new methods. As a specific method’s performance on a test set is just an estimate of its true underlying performance, the reported SOTA performances are estimates, and hence come with bias and variance.
State-of-the-art performance increasingly relies on benchmark data sets, often in combination with a challenge or a competition. The practice of hold-out test sets in competitions safe-guards against overoptimistic estimates for the individual classifiers. Data is often released after the competition has ended, and new methods can be compared to the state-of-the-art performance.
As we will show in the subsequent sections, using the highest-ranked performance as an estimate for state-of-the-art is a biased estimator, giving overoptimistic results. The mechanisms at play are those of multiplicity, a topic that is well-studied in the context of multiple comparisons and multiple testing, but has, as far as the authors are aware of, been nearly absent from the discussion regarding state-of-the-art estimates.
Biased state-of-the-art has consequences both for practical implementations and for research itself. A health care provider might want to use some machine learning method for tumour detection, but only if the method correctly classifies at least 80%. The health care provider closely follows updates on state-of-the-art performance, and are ready to invest time and money in a pilot project once the lower bound surpasses 80%. If the state-of-the-art performance is an optimistic estimate, the health care provider has now lost time and money, and perhaps also some faith in research results.
For the research community, the situation is perhaps even more dramatic. The optimistic state-of-the-art estimate is used as a standard for evaluating new methods. Not to say that a new method necessarily needs to beat the current state-of-the-art to catch interest from reviewers and editors, but it certainly helps if its performance is comparable. It has been pointed out (see Section 2) that the incremental increase in performance can obstruct the spread of new ideas that have substantial potential, but just haven’t reached their peak yet.
In this article, we provide a probability distribution for the case of multiple classifiers. We demonstrate the impact of multiplicity through simulated examples. We show how classifier dependency impacts the variance, but also that the impact is limited when the accuracy is high. Finally, we discuss a real-world example; a Kaggle competition from 2020. All code is available at https://github.com/3inar/ninety-nine.
2 Related work
Multiple comparisons is a well-studied subject within statistics, see, e.g., Tukey (1991) for an interesting discussion on some topics of multiple comparisons, among them the confidence intervals. For a direct demonstration of how multiple testing can lead to false discoveries, Bennett et al. (2009) gives an example where brain activity is detected in a (dead) salmon “looking” at photos of people. In the machine learning context, with multiple data sets and classifiers, Demšar (2006) suggests adjusting for multiple testing, according to well-known statistical theory. Our article focuses on the estimate itself and its uncertainty, and not comparison of methods.
The re-use of benchmark datasets has been addressed in several publications. Salzberg (1997) point to the problem that arises because these datasets are static, and therefore run the risk of being exhausted by overfitting. Exhausted public data sets still serve important functions, e.g., a new idea can show to perform satisfactory, even when it is not superior. Thompson et al. (2020) distinguishing between simultaneous and sequential multiple correction procedures. In sequential testing, which is the reality of public datasets, the number of tests is constantly updated. An interesting topic they bring up is the conflict between sharing incentive, open access and stable false positive rate. There are several demonstrations of the vulnerability of re-use of benchmark datasets. Teney et al. (2020) performed a successful attack on the VQA-CP benchmark dataset, that is, they are able to attain high performance with a non-generalisable classifier. They point towards the risk of ML methods capturing idiosyncrasies of a dataset. Torralba and Efros (2011) introduced the game Name That Dataset! motivating their concern of methods being tailored to a dataset rather than a problem. The lack of statistically significant differences among top performing methods is demonstrated by, e.g., Everingham et al. (2015) and Fernández-Delgado et al. (2014).
The same mechanisms of overfitting are present in challenges through multiple submissions to the validation set. Blum and Hardt (2015) introduced an algorithm that releases the public leaderboard score only if there has been significant improvement compared to the previous submission. This will prevent overfitting by hindering adjustments to random fluctuations in the score, and not true improvements of the method. Dwork et al. (2017) use the principles of differential privacy to prevent overfitting to the validation set.
Several publications show that the overfitting is not as bad as expected, and provide possible explanations. Recht et al. (2019) constructed new, independent test sets of CIFAR-10 and ImageNet. They found that the accuracy drop is dramatic, around 10%. However, the order of the best performing methods is preserved, which can be explained by the fact that multiple classes (Feldman et al. (2019)) and model similarity (Mania et al. (2019)) slows down overfitting. Roelofs et al. (2019) compared the results from the private leaderboard (one submission per team) and the public leaderboard (several submissions per team) in Kaggle challenges. Several of their analyses indicate overfitting even though their main conclusion is that there is little evidence of overfitting.
Overfitting and its consequences are avoided in most challenges by withholding the test set so that a method can only be evaluated once and not adapted to the test set performance. Hold-out datasets also prevents wrong use of cross-validation, see, e.g., Fernández-Delgado et al. (2014) and the seminal paper of Dietterich (1998). Another contribution from public challenges is the requirement of reproducibility to collect the prize money, which is a well-known concern in science, see, e.g., Gundersen and Kjensmo (2018). Public challenges typically rank the participants by a (single) performance measure, and can fall victims of Goodhart’s law, paraphrased as “When a measure becomes a target, it ceases to be a good measure.” by Teney et al. (2020). Ma et al. (2021) offers a framework and platform for evaluation of NLP methods, where other aspects such as memory use and robustness are evaluated.
The robustness/stability of a method, and the variability in performance are crucial aspects that the hold-out test set of public challenges does not offer an immediate solution to. Bouthillier et al. (2021) showed empirically that hyperparameter choice and the random nature of the data are two large contributors to variance, whereas data augmentation, dropouts, and weights initialisations are minor contributors. Choice of hyperparameters are often adjusted to validation set results, and the strategies of Blum and Hardt (2015) and Dwork et al. (2017) can hinder that, if implemented by the challenge hosts. Our contribution is a focus on better estimate for state-of-the-art performance under the current conditions, where the approach of Bouthillier et al. (2021) does not apply. Bousquet and Elisseeff (2002) investigate how sampling randomness influences accuracy, both for regression algorithms and classification algorithms, relying on the work of Talagrand (1996) who stated that “A random variable that depends (in a “smooth” way) on the influence of many independent variables (but not too much of any of them) is essentially constant.”
There is a rich literature on various aspects regarding the use of public challenges as standards for scientific publications, see, e.g., Varoquaux and Cheplygina (2022) for a much broader overview than the one given here. Concerns and solutions regarding multiplicity and better performance estimates are active research fields. We contribute to this conversation by shining the light on how multiplicity of classifiers’ performance on hold-out test sets create optimistic estimates of state-of-the-art performance.
3 Multiple classifiers and biased state-of-the-art estimation
3.1 Notation
[…]
$\,\,\,\,Y\sim\mathcal{B}ern(\theta)$
[…]
$\,\,\,\,X\sim\mathcal{B}(n,1\text{-}\theta)$
$\hat{\theta}(X)=\frac{n-X}{n}$ $:\text{estimator for }\theta,\text{ unbiased}$
$P(X=x)$ $:\text{probability of $x$ failures in one experiment}$
$P_{x}=P(X\leq x)$ $:\text{probability of at most $x$ failures in one experiment}$
[…]
$\,\,\,\,C_{x}\sim\mathcal{B}(m,P_{x})$
[…]
$\hat{\theta}^{\,max}_{SOTA}(\mathbf{X})=\max_{j}\hat{\theta}_{j}$ $:\text{estimator for }\theta_{SOTA},\text{ biased}$
[…]
3.2 Two coin-flip examples
Multiplicity is, simply put, how the probability of an outcome changes when an experiment is performed multiple times. Consider $n=20$ flips of a fair coin where the outcome is the number of heads, referred to as the number of successes in a binomial distribution, $\mathcal{B}(n,\theta)$, where the probability of success is $\theta=0.5$. Let the random variable $X$ denote the number of failures, $X\sim\mathcal{B}(n,1\text{-}\theta)$. The probability of $15$ heads or more corresponds to observing $x=n-15$ tails or less, that is, the cumulative distribution $P(X\leq x|n,1\text{-}\theta)$. This corresponds to the cumulative distribution for the estimator $\hat{\theta}(X)=\frac{n\text{-}X}{n}$, and we have that $P(\hat{\theta}(X)\geq\frac{n\text{-}x}{n})=P(X\leq x)=0.02069$ for $\hat{\theta}\geq 0.75$. In other words, if $\theta=0.5$, there is a probability of $0.02069$ that the estimate is $0.75$ or above in a series of $n=20$ trials. The probability of exactly $x=5$ can be calculated from the probability mass function (pmf) $P(X=x|n,1\text{-}\theta)=0.01479$, but the interest usually lies with the cumulative distribution. We focus on the case where $X\leq x$, but it is straightforward to substitute $P(X\leq x)$ for $P(X=x)$ in the following.
If we fix $x$ we can define the experiment described above as a Bernoulli trial, where success is defined as $X\leq x$, with corresponding probability $P(X\leq x)$, which we from now on write as $P_{x}$ for ease of notation. If we perform the coin-flip experiment $m$ times, and success is defined as $X\leq x$, we have a series of $m$ Bernoulli trials, and the number of successes follow the binomial distribution $\mathcal{B}(m,P_{x})$. Let $C_{x}$ be the random variable that denotes the number of successful experiments. The probability of, for example, $C_{x}=10$ experiments with at most $x=5$ tails, is calculated from the pmf $P(C_{x}=10|m,P_{x})$. Of special interest is the case where at least one experiment succeeds, $P(C_{x}>0)=1-P(C_{x}=0)$. In our example, if the coin-flip experiment is performed $m=100$ times, the probability of observing $15$ heads or more at least once is $P(C_{x}>0|m,P_{x})=1-P(C_{x}=0|m,P_{x})$, with $P_{x}=0.02069$ from above. In the estimation context, we have that although the probability of $\hat{\theta}\geq 0.75$ is only $0.02069$ for the individual experiments, the probability of at least one $\hat{\theta}\geq 0.75$ is $0.8765$ when the experiment is performed $m=100$ times.
Consider an experiment where a coin is flipped $n=20$ times, and the number of successes is the number of times a classifier correctly predicts the outcome. An unbiased estimator for a classifier’s probability of correct prediction, $\theta$, is ${\hat{\theta}=(n-X)/n}$, where $X$ is the number of failures. The corresponding estimate is referred to as accuracy. The true state-of-the-art for coin-flip classifiers is $\theta_{SOTA}=0.5$, since it is impossible to train a classifier to predict coin flips. If $m=1,000$ classifiers independently give their predictions for $n=20$ flips, we can consider this as a binomial experiment with $1,000$ Bernoulli trials, where the probability of success is the individual classifier’s probability of $n-X\leq n-x$ correct predictions. The probability of at least one classifier achieving an accuracy of at least $0.90$ is $0.18231-P(C_{x}=0|m,P_{x})=$ 0.1823, where $P_{x}=P(X\leq 18|n,1\text{-}\theta)=0.00020$. This means that using the top-ranked method as an estimate and ignoring multiplicity there is a substantial risk of overestimating the state-of-the-art.
3.3 The probability distribution of $\hat{\theta}^{\,max}_{SOTA}$
We define the SOTA probability of correct prediction, $\theta_{SOTA}$, as the highest probability of correct prediction among all classifiers applied to a test set, in accordance with the common set-up for public competition. Of interest is $P(\hat{\theta}^{\,max}_{SOTA}|m,n,\theta)$, the probability of $\hat{\theta}^{\,max}_{SOTA}$, being the estimated state of the art from taking the highest ranked classifier, given the number of classifiers, the size of the test set, and the probabilities of correct prediction, $\theta=\{\theta_{1},\ldots,\theta_{m}\}$.
We will in the following assume independent classifiers, and discuss dependency in Section 4. We demonstrate the impact of multiplicity in the simple case of identical probabilities of correct prediction, $\theta_{j}=\theta$ for all classifiers, $j=1,\ldots,m$, and where $\theta_{SOTA}=\theta$. We discuss non-identical probabilities in Section 4.2.
Let $n$ denote the size of the test set, $m$ be the number of classifiers, and $\theta$ be the probability of correct prediction. Let $X$ be the random variable that denotes the number of failures (wrong predictions) for a given classifier. We have that $X\sim\mathcal{B}(n,1\text{-}\theta)$. We denote by $P_{x}$ the cumulative distribution $P(X\leq x)$. The random sample, $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ consists of the number of wrong predictions for each classifier $j=1,\ldots,m$ on a test set. Let $C_{x}$ be the random variable that denotes the number of classifiers with at most $x$ failures. We have that $C_{x}\sim\mathcal{B}(m,P_{x})$, because each classifier will independently have a probability $P_{x}$ of having at most $x$ wrong predictions.
We are interested in the probability of at least one classifier having at most $x$ wrong predictions, because it corresponds to the probability function of the $\hat{\theta}^{\,max}_{SOTA}$ estimator, $P(\hat{\theta}^{\,max}_{SOTA}>\frac{n-x}{n})$, in other words, at least one $\hat{\theta}_{j}>\frac{n-x}{n}$. We can derive this from $C_{x}$:
$P(C_{x}>0|m,n,\theta).$ (1)
Note that this corresponds to the cumulative distribution function (cdf) $F(z)$ for the random variable $Z$, denoting the most wrong predictions among all classifier. We have cdf $F(z)$, the probability that the most wrong predictions among all classifiers is less than or equal to a number $z$, can easily be derived from the binomial distribution
$F(z)$ $=P(C_{z}>0|m,n,\theta)=1-P(C_{z}=0|m,P_{z})$
$=1-(1-P_{z})^{m}.$ (2)
Let $f(z)$ denote the probability mass function of $Z$. We then have
$f(z)$ $=F(z)-F(z-1)=1-(1-P_{z})^{m}-\{1-(1-P_{z-1})^{m}\}$
$=(1-P_{z-1})^{m}-(1-P_{z})^{m}=$
$=(1-\sum_{\ell=0}^{z-1}P(Z=\ell))^{m}-(1-\sum_{\ell=0}^{z-1}P(Z=\ell)-P(Z=z))^{m}$
$=-\sum_{k=1}^{m}(-1)^{k}\binom{m}{k}(1-\sum_{\ell=0}^{z-1}P(Z=\ell))^{m-k}P(Z=z)^{k}.$ (3)
The last line is derived by binomial expansion. The practical relevance of the binomial expansion is limited by the need to approximate binomial coefficient for large $m$. From Eq. 3.3, the expected value and variance can easily be calculated. An example of $F(z)$ can be seen in Fig. 2, and the corresponding $f(z)$ in Fig. 3. The pmf $f(z)=P(Z=z)$ denotes the probability of at least one classifier having exactly $z$ wrong predictions, and no classifiers having fewer than $z$ wrong predictions.
3.4 A simulated public competition example
The multiplicity mechanisms described above apply to public competitions where several teams submit their predictions on a test set. Winning teams are those with highest accuracy, $\max_{j}\hat{\theta}_{j}$, where $\hat{\theta}_{j}=(n-x_{j})/n$ is the accuracy of a classifier $j$ on the test set. This is often referred to as the state of the art, and hence, we have that $\hat{\theta}^{\,max}_{SOTA}(\mathbf{x})=\max_{j}\hat{\theta}_{j}$. As exemplified in Section 3.2, multiplicity biases the estimation of $\theta_{SOTA}$, and Section 3.3 gives us the tools to analyze its impact. In the following, we will use a significance level of $\alpha=0.05$ and two-sided confidence intervals.
Consider a classification problem with a test set of size $n=3,000$, and a classifier with probability of correct prediction $\theta=0.90$. The pmf of a single classifier is shown as a histogram in Fig. 1. The 95% confidence interval for $\theta$ if we have observed $\hat{\theta}=0.90$ is $(0.8887,0.9105)$, and we denote it by $(\theta_{\alpha/2},\theta_{1\text{-}\alpha/2})$.
If there are $m=1,000$ classifiers, each with probability of correct prediction $\theta=0.90$, the probability of at least one team achieving at least $\theta_{1\text{-}\alpha/2}$ is close to 1. This comes as no surprise given the definition of a confidence interval. It means that we can be almost certain that the accuracy of the top ranked classifier, which is the default SOTA, is above the upper limit of the confidence interval of the true performance.
On the other hand, it is straightforward to show that if each individual classifier has a probability of $P_{x}=0.00002532$ for at most $x$ failures, then, with a probability of $\alpha/2$, at least one team will achieve at least $\hat{\theta}=0.9213$. We denote this by $\theta^{m}_{1\text{-}\alpha/2}$, the multiplicity-adjusted upper limit of the $(1\text{-}\alpha)$ confidence interval.
Similarly, the expected value of $\hat{\theta}^{\,max}_{SOTA}$ can easily be calculated from Eq. 3.3, and we have $E(\hat{\theta}^{\,max}_{SOTA})=0.9173$. Both are optimistic estimates, knowing that $\theta_{SOTA}$ is $0.90$.
Within the competition context, this is not necessarily a problem, because each team has the same probability of achieving the highest accuracy given that they have the same $\theta$. In a scientific context, using $\max_{j}\hat{\theta}_{j}$ as an estimate for $\theta_{SOTA}$ without multiplicity corrections is problematic. Consider a new classifier with probability of correct prediction $\theta_{m+1}=\theta_{1\text{-}\alpha/2}$ from above, displayed as the red curve in Fig. 1. By common scientific standards, the expected performance, $E(\hat{\theta}_{m+1})=\theta_{m+1}$ is considered significantly better than the expected performance of any of the $1,000$ classifiers with $\theta=0.90$, and is of interest to the research community. But the probability of this classifier beating the upper limit of the multiplicity adjusted confidence interval is very low: $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})=0.0190$, and can be seen as the red area in Fig. 1. Similarly, $P(\hat{\theta}_{m+1}\geq E(\hat{\theta}^{\,max}_{SOTA}))=0.0996$, seen as the blue plus red area in Fig. 1. This means that by using the maximum accuracy as SOTA, we can expect that 9 out of 10 classifiers that are in fact better will still be naively considered worse.
The probability mass functions of single $\hat{\theta}$.
Figure 1: The grey histogram shows the pmf of a classifier with probability of correct prediction $\theta$. The red curve is the pmf of a classifier with probability of correct prediction $\theta_{1\text{-}\alpha/2}$, the upper limit of the $95\%$ CI for $\hat{\theta}(\mathbf{x})=\theta$. $E(\hat{\theta}^{\,max}_{SOTA})$ denotes the expected value of the highest ranked among $m=1,000$ classifiers. $\theta^{m}_{1\text{-}\alpha/2}$ is the multiplicity corrected upper limit of the $95\%$ CI. The shaded areas corresponds to $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})$ (red) and $P\big{(}\hat{\theta}_{m+1}\geq E(\hat{\theta}^{\,max}_{SOTA})\big{)}$ (blue+red).
The cumulative distribution function, $F(z)$ from Eq. 3.3, is displayed in Fig. 2, and describes the cumulative probability of at least one classifier having at most $z$ failures. Or, in other words, the cumulative probability of maximum accuracy. The number of failures are displayed on the top horizontal axis, and the corresponding accuracy on the bottom horizontal axis. Note that the horizontal axis is reversed compared to Fig. 1.
Cumulative distribution function, $F(z|m,n,\theta)$, for at least one classifier having at most $z$ failures.
Figure 2: The figure shows the cumulative distribution function for at least one classifier having at most $z$ failures. The horizontal axis denotes the number of failures, $z$, on top and the accuracy at the bottom.
The corresponding pmf, $f(z)$ from Eq. 3.3, is displayed as a histogram in Fig. 3.
Probability mass function, $f(z|m,n,\theta)$, for the smallest number of failures, $z$, among all classifiers.
Figure 3: The figure shows the corresponding pmf for Fig. 2. The horizontal axis denotes the number of failures, $z$, on top and the corresponding accuracy at the bottom.
From Eq. 3.3 we get the expected value and variance. For $m=1,000$, $n=3,000$ and $\theta=0.9$ this corresponds to $E(\hat{\theta}^{\,max}_{SOTA})=0.9173$ with a standard deviation of $0.001817$. We can see how different values for $m$, $n$ and $\theta$ affects the $\hat{\theta}^{\,max}_{SOTA}$ distribution in Table 1.
Expected values and standard deviations for $\hat{\theta}^{\,max}_{SOTA}$
m n $\theta$ $E(\hat{\theta}^{\,max}_{SOTA})$ $\sigma_{\hat{\theta}^{\,max}_{SOTA}}$
1,000 3,000 0.85 0.8707 0.002197
1,000 3,000 0.90 0.9173 0.001817
1,000 3,000 0.95 0.9624 0.001277
100 3,000 0.90 0.9135 0.002250
500 3,000 0.90 0.9163 0.001923
1,000 1,000 0.90 0.9294 0.003007
1,000 10,000 0.90 0.9096 0.001022
Table 1: The expectation and standard deviation for the SOTA estimator with varying test set size, number of classifiers and probability of correct prediction. The bold line is the example used in Fig. 3.
4 Non-i.i.d.
The probability distribution in Section 3.4 is based on independent classifiers with identical $\theta$s. This assumption is rarely true in practice. A classifier’s ability to perform better than random guesses is due to correlation between features and labels in a training set. The teams will pick up on many of the same features and correlations, regardless of which method they choose. In addition, many teams will use different versions of related methods. We investigate dependency and non-identical classifiers, and the impact on the bias. Table 2 gives an overview, details are in the following sections.
Upper bounds, expected values and standard deviations for $\hat{\theta}^{\,max}_{SOTA}$
section independent identical upper 95% CI expected value standard deviation
3.4 YES YES 0.9213 0.9173 0.001817
4.1 NO YES 0.9207 0.9140 0.003481
4.1 NO YES 0.9173 0.9140 0.001484
4.2 YES NO 0.9177 0.9130 0.002129
4.3 NO NO 0.9173 0.9101 0.003649
Table 2: The upper bound of the 95% CI, the expectation and standard deviation for the SOTA estimator for different combinations of independency and identicality. The true state-of-the-art is $\theta=0.90$.
4.1 Dependent, identical classifiers
Let $Y_{j}$ be the Bernoulli variable with value $1$ if classifier $j$ predicts a data point correctly. We have $P(Y_{j}=1)=\theta_{j}$, and $\theta_{j}=\theta$ for classifiers with equal probability of success. Mania et al. (2019) calculated the similarity, defined as the fraction of two models giving the same output, which is equivalent to
$P\big{(}\{Y_{j},Y_{k}\}=\{0,0\}\big{)}+P(\{Y_{j},Y_{k}\}=\{1,1\})$ (4)
$=(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}$
for independent models. For dependent models with correlation $\rho_{jk}$, this becomes
$(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}+2\rho_{jk}\sqrt{\theta_{j}\theta_{k}(1-\theta_{j})(1-\theta_{k})}.$ (5)
By setting $\theta_{j}=\theta$ and $\rho_{jk}=\rho$, $\rho$ can easily be estimated from the estimations of $\theta$, $P\big{(}\{Y_{j},Y_{k}\}=\{0,0\}\big{)}$ and $P(\{Y_{j},Y_{k}\}=\{1,1\})$. In Mania et al. (2019), they used the average accuracy for top-performing ImageNet models, $\hat{\theta}=0.756$. For details, see Mania et al. (2019). Note that these are multi-class problems, so a lower accuracy than for binary classification problems is to be expected. From their calculated similarity of $0.85$, the estimated correlation is $0.36$.
Boland et al. (1989) describes dependency in the majority systems context, with $\rho_{0j}=\text{corr}(Y_{0},Y_{j}),\,j=1,\ldots m$, and the $Y_{j}$s are independent given $Y_{0}$. This gives $\rho_{jk}=\rho_{0j}^{2}$. The observations from $Y_{0}$ does not contribute directly to $\hat{\theta}^{\,max}_{SOTA}$, but can be seen as, e.g., the influence that a common training set has on the classifiers.
The binomial distribution for dependent Bernoulli trials is not an easy problem, see, e.g., Ladd (1975) and Hisakado et al. (2006). We here demonstrate by simulation how dependency reduces the bias. We use the estimated correlation from Mania et al. (2019), which corresponds to $\rho_{0j}=0.6$ in the multiplicity set-up with an external $Y_{0}$.
We use the set-up from Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, and probability of correct prediction $\theta=0.90$. We simulate two scenarios; one where $Y_{0}$ is random, and one where the $y_{0}$’s are fixed with exactly $(n-x)/n=0.90$ correct predictions. The fixed scenario can correspond to fixed training conditions, e.g., models pre-trained on ImageNet. The random scenario picks up the randomness in, e.g., a training set. Although all teams have the same training set, the training set is a random sample.
The upper bound of the $95\%$ confidence interval for the dependent classifiers is at $0.9173$ for the fixed $y_{0}$ and $0.9207$ for the random $Y_{0}$, but they have the same expected value.
Minimum number of failures in $m=1,000$ classifiers.
(a) $\rho_{0j}=0.6$, random $Y$
(b) $\rho_{0j}=0.6$, fixed $y$
Figure 4: The minimum number of failures in $100,000$ simulations, with probability of success $\theta=0.9$.
Note that, with $\theta=0.90$, the similarity as described by Mania et al. (2019) is $0.82$ for two independent classifiers. When the probability of correct prediction is high, there is little room for variation between classifiers, and the inherent similarity is high.
4.2 Non-identical, independent classifiers
The Poisson Binomial distribution, see, e.g., Wang (1993), describes the sum of $n$ non-identical independent Bernoulli trials with success probabilities $p_{1},p_{2},\ldots,p_{n}$. The pmf for $K$ number of successes is
$P(K=k)=\sum\limits_{{A\in G_{k}}}\prod\limits_{{i\in A}}p_{i}\prod\limits_{{j\in A^{c}}}(1-p_{j}),$ (6)
where $G_{k}$ is all subsets that can be selected from {1, 2, …, n}.
In the multiple classifier context, we are interested in generalising Eq. 3.3 to the non-identical case.
$F(z)=P(C_{z}>0|m,n,\mathbf{\theta})=1-P(C_{z}=0|m,n,\mathbf{\theta}),$ (7)
where $P(C_{z}=0)$ corresponds to 00 classifiers with non-identical probabilities of success, $P(Z\leq z|\theta_{j})$. The cdf is
$F(z)$ $=1-\sum\limits_{{A\in G_{0}}}\prod\limits_{{i\in A}}P(Z\leq z|\theta_{i})\prod\limits_{{j\in A^{c}}}(1-P(Z\leq z|\theta_{j}))$
$=1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}.$ (8)
The pmf is
$f(z)$ $=F(z)-F(z-1)$
$=1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}-\{1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z\text{-}1|\theta_{j})\big{)}\}$
$=\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z\text{-}1|\theta_{j})\big{)}-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}$ (9)
$=\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z\text{-1}}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}$
Analytically, it is more complex than the case of identical classifiers, but it is easy to simulate.
We use a similar set-up as in Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, and probabilities of correct prediction $\theta=\{\theta_{1},\ldots,\theta_{m}\}$, where $\min\theta=0.875,\,\max\theta=0.90$, and they are equally spaced.
The simulated non-identical upper bound of the 95% confidence interval is $0.9177$, with $100,000$ repetitions. The $E(\hat{\theta}^{\,max}_{SOTA})$ is $0.9129$, with standard deviation of $0.0021$. The cdf and pmf are displayed in Fig. 5 and Fig. 6.
Cumulative distribution function for non-identical classifiers.
Figure 5:
Probability mass function for non-identical classifiers.
Figure 6:
4.3 Non-identical, dependent classifiers
The most complex, but also the most realistic case, is where classifiers are non-identical and dependent. An analytical solution is not known Hisakado et al. (2006), but simulations are easy. The results of Boland et al. (1989) are easily extended, and we have that
$P(Y_{j}|Y_{0}=1)$ $=\frac{\rho\sqrt{\theta_{j}(1-\theta_{j})\theta_{0}(1-\theta_{0})}+\theta_{j}\theta_{0}}{\theta_{0}}$ (10)
$P(Y_{j}|Y_{0}=0)$ $=\frac{-\rho\sqrt{\theta_{j}(1-\theta_{j})\theta_{0}(1-\theta_{0})}+\theta_{j}(1-\theta_{0})}{(1-\theta_{0})}.$ (11)
Note that, since $P\leq 1$, it follows from 11 that
$\theta_{j}\geq\frac{\rho^{2}\frac{\theta_{0}}{1-\theta_{0}}}{1+\rho^{2}\frac{\theta_{0}}{1-\theta_{0}}},$ (12)
and similarly for $0\leq P$ and 10.
We get the a the simulated dependent upper bound of the 0.95 confidence interval of $0.91733$, with $100,000$ repetitions.
5 Real world examples
As we have shown above with simulated data, the observed maximum is an overoptimistic estimate for SOTA. These results are based on known state-of-the-art, but faced with real world data, the problem to be solved is of opposite nature: The observed $\hat{\theta}^{\,max}$ is known, but the true SOTA must be estimated. In the public competition or public data set context there are many factors that come in to play, making the bias calculations everything but straightforward.
We here demonstrate different aspects of real world analysis through SOTA estimation of three different cases: Obesity Risk (5.1.1), Cassava Leaf (5.1.2, and Melanoma (5.2). Obesity Risk and Cassava Leaf have accuracy, defined as $\hat{\theta}$, as their ranking criterion, whereas Melanoma has area under the receiver operating characteristic (ROC) curve (AUC) as ranking criterion. To estimate multiplicity effect on the AUC scores is not as straightforward as with accuracy, but the mechanisms are the same and AUC is a more common ranking criterion than accuracy.
In a public competition with $m$ participating teams, let $\hat{\theta}_{j}$, $j=1,\ldots,m$ be the estimated probability of correct prediction for each classifier based on the performance on an independent test set of size $n$. We will refer to these as the competition observations. $\hat{\theta}^{\,max}$ is the observed performance of the winning team. Among the different scenarios presented in Section 4, the non-identical dependent is the most realistic model. Whereas in Table 2 we simulated $E(\hat{\theta}^{\,max})$ from known $\theta_{1},\ldots,\theta_{m}$, the task is now to estimate $\theta_{SOTA}$ based on $\hat{\theta}_{1},\ldots,\hat{\theta}_{m}$. A ’natural’ approach is to find $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ that gives $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$. That is, find a set of $\theta$’s so that the expected value of the observed maximum on a test set of size $n$ is equal to the competition observed maximum. We use a straightforward set-up, where the simulated $\hat{\tilde{\theta}}_{j}$ is the mean value of $n$ realisations of $Y_{j}$, where $P(Y_{j})=\tilde{\theta}_{j}$. The simulation is repeated $10,000$ times, and for each repetition the observed maximum value is recorded. $E(\hat{\tilde{\theta}}^{\,\max})$ is the mean value over the repetitions.
In the studies in Section 4, the $\theta_{1},\ldots,\theta_{m}$ were chosen to be equally spaced, but this does not align with the competition observations, as seen in Figures 7(a), 8(a) and 9. We base our suggested $\tilde{\theta}_{j}$’s on cropping the competition observations, as seen, e.g., in Figure 7(c), but this is far from the only approach. Shrinking, which we also investigated, might give a more realistic set of $\tilde{\theta}_{j}$’s, but it also requires an extra parameter. For the examples presented here shrinking gave similar results to cropping, and are therefore not included.
An alternative to the expected value of the observed maximum is the upper limit of the 95% CI. The latter is more in line with common interpretation of the uncertainty of observed results for research, but which to choose will depend on the aim of the analysis. We demonstrate the impact in the Obesity Risk example, reported in Table 3.
Not every $\hat{\theta}^{\,max}$ is a result of multiplicity, and we demonstrate this in the Cassava Leaf example.
5.1 Estimating $\theta_{SOTA}$
5.1.1 Multi-Class Prediction of Obesity Risk
Multi-Class Prediction of Obesity Risk is a competition in the 2024 Kaggle Playground Series, using a synthetic data set generated from the Obesity or CVD risk data set (10.34740/kaggle/dsv/7009925). The test set size was $n=13,840$, and the number of participating teams were $m=3,498$ (excluding teams with accuracy lower than chance). The best performing team had $\hat{\theta}^{\,\max}=0.9116$. Figure 7(a) shows the competition observations, $\hat{\theta}_{j}$’s, zoomed in at the $3,081$ best performing teams. The 95% (not multiplicity-corrected) CI for $\hat{\theta}^{\,max}$ is $(0.9067,0.9162)$ and is shown as the red line in Figure 7(a), the lower bound enveloping $920$ of the competition observations. This should raise suspicion that the right tail seen in Figure 7(a) can be the product of multiplicity.
(a) The competition observations
(b) A simulated realisation
(c) Cropped competition observations
(d) A simulated realisation
Figure 7: The competition observations and simulated realisation of classifiers with corresponding probabilities of correct classification. The red line is the 95% CI for $\hat{\theta}^{\,\max}$ in (a).
To demonstrate the impact of ignoring multiplicity, let $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ be equal to the competition observations and calculate the expectation of the maximum value of a realisation. We use the set-up from Section 4.3 with correlation coefficient $\rho=0.6$, and random $Y_{0}$ with $\theta_{0}=\hat{\theta}^{\,max}$. Note that competition observations with low values are excluded in accordance with Eq. 12, but they are not influencing the maximum.
A realisation is shown in Figure 7(b). The upper tail, whose tip represents the maximum observed value, is stretched according to the variances of the $Y_{j}$s.
The estimated expected maximum is $E(\hat{\tilde{\theta}}^{\,max})=0.9159$, far above the observed maximum. The 95% CI is calculated by $B=1,000$ bootstrap samples of size $m$ drawn from the competition observations, and then the simulation as described above is conducted for each bootstrap sample. The values are reported in Table 3. The lower bound exceeds $\hat{\theta}^{\,\max}$, and we can conclude that the maximum observed value is an unlikely candidate for state-of-the-art.
Perhaps a more interesting problem is finding a set $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ that can result in realisations similar to the competition observations in terms of maximum observed value, and let the maximum $\tilde{\theta}_{j}$ be the state-of-the-art candidate.
To obtain a set $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ where the expected maximum is the same as the maximum of the competition observations, $E(\hat{\tilde{\theta}}^{\,\max})=\hat{\theta}^{\,\max}$, we crop the competition observations, as seen in Figure 7(c), and where $E(\hat{\tilde{\theta}}^{\,\max})$ is calculated from simulations as described above. Figure 7(d) shows one realisation of the cropped distribution, where the maximum observed value is close to $\hat{\theta}^{\,max}$. By cropping at $0.9062$ we obtain $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$, and this serves as a new candidate for the state-of-the-art.
It can be argued that the new SOTA candidate is at risk of being a optimistic estimate, since the competition observations are from only one test set. The upper bound of the 95% CI can mitigate such a risk, and the competition observations are cropped to fit that criterion. The CI is calculated by bootstrapping, as described above. The results are shown in Table 3.
Estimating ${\theta}_{SOTA}$
Approach $\hat{\theta}_{SOTA}$ 95% CI $E(\max_{j}\theta_{j})$ teams
no mult. $\hat{\theta}^{\,\max}$ (0.9067,0.9162) - -
Kaggle accuracies $\hat{\theta}^{\,\max}$ (0.9128,0.9191) 0.9159 -
cropping 0.9062 (0.9084,0.9148) $\hat{\theta}^{\,\max}$ 1134
0.9026 (0.9050,$\hat{\theta}^{\,\max}$) 0.9083 1844
Table 3: The 95% CI and the expectation and for the SOTA estimator for different estimation strategies. The bootstrap standard deviations are in the order of magnitude $10^{\text{-}5}$-$10^{\text{-}4}$, and are therefore not displayed.
Table 3 demonstrates that taking the effect of multiplicity into account changes how results are perceived dramatically. A more likely SOTA value is $0.9062$, and even though its reduction from best performance is only about $0.005$, the effect it has on who we consider performing at SOTA-level is dramatic: around one third of the teams have observed performance above the new SOTA. The more conservative approach using the upper bound of the 95% CI to estimate SOTA gives an even more dramatic result, and either way the impact is substantial.
5.1.2 Cassava Leaf Disease
In this example, the highest reported accuracy is unlikely to be the product of multiplicity. For The Cassava Leaf Disease Classification competition, the task was to classify images of cassava leaves. The training set consisted of $21,367$ labelled images of 5 classes, and the test set of approximately $15,000$ images. $3,752$ teams participated (excluding those with accuracy lower than chance).
The three top performances appear as outliers, as seen in Figure 8(a). The 95% CI of $\hat{\theta}^{\,\max}=0.9132$ is $(0.9086,0.9177)$, no other observations are within this interval. By performing simulations without cropping, we get $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$, and although the upper tail is stretched, its tip does not surpass $\hat{\theta}^{\,max}$.
(a) The Kaggle performances
(b) One realisation.
Figure 8: The competition observations and one realisation.
Several explanations for the outlying performances of the three top teams have been put forward, and the interested reader can take a deep-dive in the comment section of the competition. Since this is not the topic of the present work, we will not engage in the discussion here.
5.2 Simulation of AUCs
We discuss here an example, the SIIM-ISIC Melanoma Classification competition from 2020 because it follows many of the common structures of public challenges. The winning team chose many of the verified strategies from Section 2.
The SIIM-ISIC Melanoma Classification competition consisted of more than $40,000$ images of skin lesions from various sites, split into training ($33,126$), validation ($3,295$) and test set ($7,687$). The data set is highly imbalanced, with less than $2\%$ melanomas, which reflects the reality in many cancer detection situations. A full description can be found in Rotemberg et al. (2021). The size of the test set is smaller than the recommendations of, e.g., Roelofs et al. (2019), but reflects the reality of benchmark datasets and public challenges, as seen, e.g., in the overview of Willemink et al. (2020). We want to emphasise that the shortcomings of this competition are not unique, and that is precisely why it serves as a useful example. Participants were allowed to use additional data for training if the data was shared on the competition web site. The prize money was set to $10,000$ USD for the best-performing method, measured as AUC on the hold-out test set, published in the private leaderboard as seen in Figure 9. Participants could get intermediate results from the validation set, published on a public leaderboard.
Figure 9: Histogram of the AUCs from the kaggle leaderboard, test set.
The team ‘All Data Are Ext’ consisting of three Kaggle Grandmasters won the competition, achieving an AUC of $0.9490$ on the test set, and their method and strategy is explained in Ha et al. (2020). Although we show that the AUC is an optimistic estimate for the melanoma detection state-of-the-art, we do not question their rank in the challenge. Quite the contrary; the research in Section 2 supports their strategies. They had a low rank on the public leaderboard (880th out of 3300). Their AUC decreased from the public to the private leaderboard, but it was still within the confidence interval. They used additional data from previous years’ competitions, and they also did data augmentation. They chose ensemble learning as a strategy to avoid overfitting, a strategy implicitly supported by Talagrand (1996). They also trained on multiple classes, a strategy that Feldman et al. (2019) showed will slow down the overfitting.
In a binary classification problem, the calculation of AUC is based on scores of random variables from the positive and the negative class. The score function plus a threshold is the classifier. The AUC is mathematically equivalent to the probability that a random variable from the positive class will be ranked higher, in terms of score, than a random variable from the positive class (Hanley and McNeil, 1982).
Let $score(\cdot)$ be a function that assigns a score to a random variable. Let $S_{+}=score(W_{+})$, where $W_{+}$ comes from the positive class, and $S_{-}=score(W_{-})$, where $W_{-}$ comes from the negative class. Then $\text{AUC}=P(S_{-}<S_{+})=P(S_{-}-S_{+}<0)$, which is equivalent to the cdf at value 00. Hence to simulate classifier scores according to a given true or expected AUC, $a$, we simply need to specify distributions for the variables $S_{-},S_{+}$ such a way that $P(S_{Y}-S_{X}<0)=a$.
This is particularly simple with Normal distributions, since linear combinations of Normal distributions yields another Normal. Let $S_{+}$ and $S_{-}$ have the distributions
$S_{+}$ $\sim\mathcal{N}(\mu_{+},\sigma^{2}_{+})$ (13)
$S_{-}$ $\sim\mathcal{N}(\mu_{-},\sigma^{2}_{-}),$ (14)
where $\mathcal{N}(\mu,\sigma)$ is the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. We then have that
$\text{AUC}=\Phi\left(\frac{0-(\mu_{-}-\mu_{+})}{\sqrt{\sigma_{-}^{2}+\sigma_{+}^{2}}}\right),$ (15)
by linear combination of two Normal distributions, where $\Phi$ is the cdf of the standard Normal distribution. Specifying any three of the four parameters above decides the last one for a chosen AUC = a. Figure 10(a) shows an example where AUC = 0.90.
The distribution of classifier scores, $S$, is a mixture distribution of the $S_{+}$ and the $S_{-}$ distributions, where the weights, $\pi$ and $1-\pi$, correspond to the probability of selecting a positive or a negative random sample, and has pdf
$\pi f_{+}(s)+(1-\pi)f_{-}(s),$ (16)
where $f_{+}(s)$ and $f_{-}(s)$ are the pdfs of the positive and the negative class, respectively. Expected value and variance of $S$ are
$\mu$ $=\pi\mu_{+}+(1-\pi)\mu_{-},\text{ and}$ (17)
$\sigma^{2}$ $=\pi(\sigma_{+}^{2}+\mu_{+}^{2})+(1-\pi)(\sigma_{-}^{2}+\mu_{-}^{2})-\mu^{2}.$ (18)
The uncertainty of AUC can be calculated through bootstrapping, but without access to the predictions for each data point in the test set, the analyst must make some assumptions about the shape of the ROC curve.
Figure 10(a) shows the pdfs of negative and positive scores with an AUC of 0.90. The rug shows an example draw with 3000 images, with the probability of drawing a positive image is 0.017, corresponding to the class balance in the Melanoma challenge. Fig. 10 shows different shapes of ROC curves.
(a) Score pdfs for a classifier with AUC= $0.90$; grey for negatives, black for positives. The rug shows an example draw from these.
(b) ROC curves for a classifier with expected AUC of $0.90$ across 100 realisations of (a).
Figure 10: Example of what an AUC of $0.90$ might look like in practice. The right figure shows $100$ realizations of a ROC curve from a competition with $3000$. Out of the $3000$ images, $52$ are positives, reflecting the proportion of the Kaggle melanoma challenge. The left figure shows the underlying score distributions that generated the curves of Figure 10(b).
However, with high AUC, the shape has less impact, and the analyst’s choices are not suggestive of the overall conclusion.
5.2.1 Simulation results for uncorrelated classifiers
For simplicity we choose $\mu_{-}=0$ and $\sigma_{-}=\sigma_{+}=1$. We then solve for $\mu_{+}$ numerically. Now we can generate scores from the two resulting normal distributions.
To demonstrate the impact that multiplicity can have in a public competition, we start with a simulated example similar to the one in Section 3.4. Consider a classification problem with a test set of size $n=3,000$, number of participating teams, $m=1,000$, where each classifier has a true underlying AUC of $a=0.90$, and the classifiers are independent.
Whereas the distribution of $\hat{\theta}^{\,\max}$ is independent of the class balance as long as the probability of correct prediction is independent of class, this is not true for maximum of $\hat{AUC}$, where the variance increases with class imbalance. We have chosen the same class balance as in the Melanoma challenge, as this reflects real world examples. The 95% CI of a single classifier with observed AUC equal to its true AUC of $0.90$ is (.$(.8558,0.9376)$, and gives a strong indication of what we can expect when $1,000$ teams apply their classifiers on a test set.
Figure 11: Example distribution of simulated observed state of the art with parameters as in Figure 10. This is an analogous figure to Figure 3 above.
Figure 11 shows the simulated distribution of the observed maximum AUC over $10,000$ repetitions of a competition where every classifier has a true AUC of $0.90$. The simulated expected value is $0.9562$, the standard deviation is $0.004459$, and the simulated 95% CI is $(0.9486,0.9662)$.
To give an unbiased estimate of the SOTA performance for the Melanoma challenge, we follow the same procedure as in Section 5.1: The competition observations are cropped in such a manner that their realisations give an expected value of $0.9490$, the observed maximum.
(a) One bootstrap sample from the cropped distribution.
(b) One realisation of the bootstrap sample in (a).
Figure 12: Histogram of cropped competition observations. Here the state of the art is an AUC of 0.8975 which leads to the expected observed maximum AUC in a competition being 0.9490.
Figure 12(a) shows one bootstrap sample from the kaggle observations cropped at $0.8975$. Number of classifiers is set to $m=3,173$, which is as many as there are classifiers in the melanoma challenge with AUC $>0.5$. Figure 12(b) shows a corresponding realisation, where the expected maximum is $0.9490$.
There are $2022$ teams with performance above the new SOTA candidate.
5.2.2 Correlation between classifiers in the AUC simulations
For the moment we have not extended our AUC simulations to include correlations, but it should be possible to adopt a similar setup as in Sections 4.1 and 4.3: each classifier produces scores that correlate to those of a reference classifier. Let $S_{0^{+}}$ and $S_{0^{-}}$ be the scores of the reference classifier, for random variables from the positive and negative class, respectively.
Let $\rho_{+}=corr(S_{0^{+}},S_{j^{+}})$ be the correlation between the reference classifier scores $S_{0^{+}}$ and the scores of classifier $j$, $S_{j^{+}}$ for the positive class. If they both follow Normal distributions as outlined above, then the bivariate distribution is
$\begin{pmatrix}S_{0^{+}}\\ S_{j^{+}}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mu_{0^{+}}\\ \mu_{j^{+}}\end{pmatrix},\begin{pmatrix}\sigma_{0^{+}}^{2}&\rho\sigma_{0^{+}}\sigma_{j^{+}}\\ \rho\sigma_{0^{+}}\sigma_{j^{+}}&\sigma_{j^{+}}^{2}\end{pmatrix}\right),$
and equivalently for the negative class. However, the joint distribution of $(S_{0},S_{j})$ is not uniquely defined by the correlation, and simulations are not straightforward.
6 Discussion
We have addressed the mechanisms of multiplicity and how they affect the bias of the SOTA estimator. When accuracy is used as a performance measure, the exact probability distribution can be provided, and known analyses methods can be engaged. The bias in SOTA is substantial in our example, and possibly hinders new methods to attract interest from the scientific community. We have also demonstrated that dependency might not have a big impact on the bias when the probability of correct prediction is already high. The exception is when the dependency is very strong.
We have presented the simple case with an analytical solution, where the probabilities of correct prediction are known.
We have also presented three case-studies from Kaggle competitions, where we demonstrate how a different SOTA candidate than the observed maximum can be estimated, and how it impacts the perceived results.
This work shows that a multiplicity-adjusted state-of-the-art estimate can be calculated fairly easily through bootstrapping, and can serve as a better reference point when introducing new methods to the scientific community.

### v5

@@
 UiT - The Arctic University of Norway
-1 Introduction: Population vs sample state-of-the-art performance
-Machine learning methods are commonly evaluated and compared by their performance on data sets from public repositories, where methods can be evaluated under identical conditions and across time.
-Theoretical results play an important role in machine learning research, but for the last couple of decades the success of increasingly complex methods has pushed the proof towards the pudding: performance on benchmark data sets.
-The numbers don’t lie: if method A performs better than method B on a specific data set, that is what it did.
-But the truth that the numbers tell is limited.
-We think of a data set, for example a collection of images, as a random sample from an underlying population.
-In science we are interested in the performance on the population, and the observed performance acts as an estimate.
-A good estimator gives a reliable result, but overfitting to the data set, originating from wrong use of cross-validation, data leakage and data reuse, can give overly optimistic estimates.
-An independent test set, often referred to as a hold-out set, is a sample that is not used for development of the method, and gives an unbiased estimator.
-Competitions and challenges with prize money typically withhold the test set so that it becomes impossible for the participants to use it for development.
-Although we know these things we have a tendency to extrapolate the performance metric and interpret it as something about the world rather than something about a data set.
-If we always appended “on MNIST” when we said so-and-so method had accuracy so-and-so nobody would care.
-Who cares about MNIST?
-We hold out data precisely because we want to estimate some sort of long-term behavior on similar data.
-An unbiased estimator solves only half of the problem; the variance tells just as much as the bias regarding the quality of a method.
-For machine learning methods, the sources of variance are plentiful.
-One is the random nature of the test set, the training set and the validation set.
-In addition comes the setting of hyperparameters, which can lead to substantial variation, and minor variance sources such as random seeds and stochastic dropouts.
-The sources of variance that are part of the training are often referred to as influencing the method’s stability or robustness.
-With knowledge about the behaviour of both the bias and the variance of an estimator, the error can be predicted, often in terms of a confidence interval.
-It is certainly more useful to know what the lower limit of the 95% confidence interval is, than the observed performance on a specific test set.
-The highest ranked performance on a problem is referred to as state-of-the-art (SOTA) performance, and is used, among other things, as a reference point for publication of new methods.
-As a specific method’s performance on a test set is just an estimate of its true underlying performance, the reported SOTA performances are estimates, and hence come with bias and variance.
-State-of-the-art performance increasingly relies on benchmark data sets, often in combination with a challenge or a competition.
+1 Estimating state-of-the-art performance
+In this article, we direct attention to the reporting of state-of-the-art classifier performance by presenting the sample maximum estimator distribution function and analysing its bias.
+This analysis is done using the framework of multiplicity.
+Within the fields of research that develop, refine, test and use classifiers, the term state-of-the-art (SOTA) commonly refers to the highest reported performance to date, and is used as a reference point when presenting and comparing models.
+The reported performance of a classifier, originating from applying that classifier to a specific data set, a sample, is in effect an estimate of that classifier’s underlying population performance.
+The highest reported performance among a number of classifiers serves as an estimate of SOTA, although the term ’estimate’ is seldom used explicitly in the SOTA context.
+The highest reported performance corresponds to the sample maximum of population performances, and this estimator is biased when used to estimate SOTA.
+The sample maximum gives in overoptimistic estimates of the maximum population performance, and is problematic as a reference point in research.
+The mechanisms creating this bias comes from multiplicity, and the fact that the we pick the highest sample performance among multiple classifier results.
+In its simplest form, with independent binary classifiers with identical probability of success, the distribution function of the sample maximum estimator exists in closed form with only a few parameters, and can be directly applied to calculate and analyse the bias in a certain situation.
+With growing complexity, such as mutual dependencies and more elaborate performance measures, the distribution function of the estimator is no longer known, but the mechanisms at play and the pitfalls to be aware of are much the same, and limited studies can be done by straightforward simulations.
+Machine learning models are commonly evaluated and compared by their performance on data sets from public repositories, where models can be evaluated under identical conditions and across time.
+Theoretical results play an important role in machine learning research, but for the last couple of decades the success of increasingly complex models has pushed the proof towards the pudding: performance on benchmark data sets.
+The numbers don’t lie: if model A performs better than model B on a specific data set, that is what it did.
+But the truth that these numbers tell is limited.
+We are interested in the population performance, and the sample performance is only an estimate.
+A good estimating procedure gives a reliable result, and there are plenty of pitfalls to be aware of: Overfitting to the data set, originating from wrong use of cross-validation, data leakage and data reuse, can give overly optimistic estimates.
+An independent test set, often referred to as a hold-out set, is a sample that is not used for development of the model, and prevents overfitting.
+Competitions and challenges with prize money typically withhold the test set so that it becomes impossible for the participants to use it for training.
+The highest ranked performance on a problem is referred to as SOTA performance, and is used, among other things, as a reference point for publication of new models.
+As a specific model’s performance on a test set is just an estimate of its true underlying performance, the reported SOTA performances are estimates, and hence come with bias and variance.
 The practice of hold-out test sets in competitions safe-guards against overoptimistic estimates for the individual classifiers.
-Data is often released after the competition has ended, and new methods can be compared to the state-of-the-art performance.
-As we will show in the subsequent sections, using the highest-ranked performance as an estimate for state-of-the-art is a biased estimator, giving overoptimistic results.
-The mechanisms at play are those of multiplicity, a topic that is well-studied in the context of multiple comparisons and multiple testing, but has, as far as the authors are aware of, been nearly absent from the discussion regarding state-of-the-art estimates.
-Biased state-of-the-art has consequences both for practical implementations and for research itself.
+Data is often released after the competition has ended, and new methods can be compared to the SOTA performance.
+As we will show in the subsequent sections, using the highest-ranked performance as an estimate for SOTA is a biased estimator, giving overoptimistic results.
+The mechanisms at play are those of multiplicity, a topic that is well-studied in the context of multiple comparisons and multiple testing, but has, as far as the authors are aware of, been nearly absent from the discussion regarding SOTA estimates.
+Biased SOTA estimation has consequences both for practical implementations and for research itself.
 A health care provider might want to use some machine learning method for tumour detection, but only if the method correctly classifies at least 80%.
-The health care provider closely follows updates on state-of-the-art performance, and are ready to invest time and money in a pilot project once the lower bound surpasses 80%.
-If the state-of-the-art performance is an optimistic estimate, the health care provider has now lost time and money, and perhaps also some faith in research results.
+They follow updates on SOTA performance, and are ready to invest time and money in a pilot project once the lower bound surpasses 80%.
+If the SOTA performance is an optimistic estimate, the health care provider has now lost time and money, and perhaps also some faith in research results.
 For the research community, the situation is perhaps even more dramatic.
-The optimistic state-of-the-art estimate is used as a standard for evaluating new methods.
-Not to say that a new method necessarily needs to beat the current state-of-the-art to catch interest from reviewers and editors, but it certainly helps if its performance is comparable.
+The optimistic SOTA estimate is used as a standard for evaluating new methods.
+Not to say that a new method necessarily needs to beat the current SOTA to catch interest from reviewers and editors, but it certainly helps if its performance is comparable.
 It has been pointed out (see Section 2) that the incremental increase in performance can obstruct the spread of new ideas that have substantial potential, but just haven’t reached their peak yet.
@@
 We demonstrate the impact of multiplicity through simulated examples.
-We show how classifier dependency impacts the variance, but also that the impact is limited when the accuracy is high.
-Finally, we discuss a real-world example; a Kaggle competition from 2020.
+We show how classifier dependency impacts the variance, but also that the impact is limited when the probability of success is high.
+Finally, we discuss three real-world examples that each highlights different aspects multiplicity and SOTA estimation.
 All code is available at https://github.com/3inar/ninety-nine.
 2 Related work
-Multiple comparisons is a well-studied subject within statistics, see, e.g., Tukey (1991) for an interesting discussion on some topics of multiple comparisons, among them the confidence intervals.
+Multiple comparisons is a well-studied subject within statistics, see, e.g., Tukey (1991) for an interesting discussion on some topics of multiple comparisons, among them the confidence intervals (CIs).
 For a direct demonstration of how multiple testing can lead to false discoveries, Bennett et al.
-(2009) gives an example where brain activity is detected in a (dead) salmon “looking” at photos of people.
+(2009) give an example where brain activity is detected in (dead) salmon “looking” at photos of people.
 In the machine learning context, with multiple data sets and classifiers, Demšar (2006) suggests adjusting for multiple testing, according to well-known statistical theory.
@@
 Thompson et al.
-(2020) distinguishing between simultaneous and sequential multiple correction procedures.
+(2020) distinguish between simultaneous and sequential multiple correction procedures.
 In sequential testing, which is the reality of public datasets, the number of tests is constantly updated.
@@
 Blum and Hardt (2015) introduced an algorithm that releases the public leaderboard score only if there has been significant improvement compared to the previous submission.
-This will prevent overfitting by hindering adjustments to random fluctuations in the score, and not true improvements of the method.
+This will prevent overfitting by hindering adjustments to random fluctuations in the score and not true improvements of the method.
 Dwork et al.
-(2017) use the principles of differential privacy to prevent overfitting to the validation set.
+(2017) used the principles of differential privacy to prevent overfitting to the validation set.
 Several publications show that the overfitting is not as bad as expected, and provide possible explanations.
@@
 (2019)) and model similarity (Mania et al.
-(2019)) slows down overfitting.
+(2019)) slow down overfitting.
 Roelofs et al.
@@
 Ma et al.
-(2021) offers a framework and platform for evaluation of NLP methods, where other aspects such as memory use and robustness are evaluated.
+(2021) offered a framework and platform for evaluation of NLP methods, where other aspects such as memory use and robustness are evaluated.
 The robustness/stability of a method, and the variability in performance are crucial aspects that the hold-out test set of public challenges does not offer an immediate solution to.
@@
 (2017) can hinder that, if implemented by the challenge hosts.
-Our contribution is a focus on better estimate for state-of-the-art performance under the current conditions, where the approach of Bouthillier et al.
+Our contribution is a focus on better estimate for SOTA performance under the current conditions, where the approach of Bouthillier et al.
 (2021) does not apply.
-Bousquet and Elisseeff (2002) investigate how sampling randomness influences accuracy, both for regression algorithms and classification algorithms, relying on the work of Talagrand (1996) who stated that “A random variable that depends (in a “smooth” way) on the influence of many independent variables (but not too much of any of them) is essentially constant.”
+Bousquet and Elisseeff (2002) investigated how sampling randomness influences accuracy, both for regression algorithms and classification algorithms, relying on the work of Talagrand (1996) who stated that “A random variable that depends (in a “smooth” way) on the influence of many independent variables (but not too much of any of them) is essentially constant.”
 There is a rich literature on various aspects regarding the use of public challenges as standards for scientific publications, see, e.g., Varoquaux and Cheplygina (2022) for a much broader overview than the one given here.
 Concerns and solutions regarding multiplicity and better performance estimates are active research fields.
-We contribute to this conversation by shining the light on how multiplicity of classifiers’ performance on hold-out test sets create optimistic estimates of state-of-the-art performance.
+We contribute to this conversation by shining the light on how multiplicity of classifiers’ performance on hold-out test sets create optimistic estimates of SOTA performance.
 3 Multiple classifiers and biased state-of-the-art estimation
@@
 $\,\,\,\,X\sim\mathcal{B}(n,1\text{-}\theta)$
-$\hat{\theta}(X)=\frac{n-X}{n}$ $:\text{estimator for }\theta,\text{ unbiased}$
+$\hat{\theta}(X)=\frac{n-X}{n}$ $:\text{estimator for }\theta$
+$\hat{\theta}=\frac{n-x}{n}$ $:\text{estimate of }\theta$
 $P(X=x)$ $:\text{probability of $x$ failures in one experiment}$
 $P_{x}=P(X\leq x)$ $:\text{probability of at most $x$ failures in one experiment}$
+$\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ $:\text{number of failures for classifiers }j=1,\ldots,m$
+$C_{x}$ $:\text{number of classifiers with at most $x$ failures}$
+$\,\,\,\,C_{x}\sim\mathcal{B}(m,P_{x})$
+[…]
+$\,\,\,\,z=0,1,2,\ldots,n$
 […]
 […]
-$\,\,\,\,C_{x}\sim\mathcal{B}(m,P_{x})$
 […]
 […]
-$\hat{\theta}^{\,max}_{SOTA}(\mathbf{X})=\max_{j}\hat{\theta}_{j}$ $:\text{estimator for }\theta_{SOTA},\text{ biased}$
 […]
-3.2 Two coin-flip examples
+[…]
+[…]
+3.2 A coin-flip example
 Multiplicity is, simply put, how the probability of an outcome changes when an experiment is performed multiple times.
-Consider $n=20$ flips of a fair coin where the outcome is the number of heads, referred to as the number of successes in a binomial distribution, $\mathcal{B}(n,\theta)$, where the probability of success is $\theta=0.5$.
-Let the random variable $X$ denote the number of failures, $X\sim\mathcal{B}(n,1\text{-}\theta)$.
-The probability of $15$ heads or more corresponds to observing $x=n-15$ tails or less, that is, the cumulative distribution $P(X\leq x|n,1\text{-}\theta)$.
-This corresponds to the cumulative distribution for the estimator $\hat{\theta}(X)=\frac{n\text{-}X}{n}$, and we have that $P(\hat{\theta}(X)\geq\frac{n\text{-}x}{n})=P(X\leq x)=0.02069$ for $\hat{\theta}\geq 0.75$.
-In other words, if $\theta=0.5$, there is a probability of $0.02069$ that the estimate is $0.75$ or above in a series of $n=20$ trials.
-The probability of exactly $x=5$ can be calculated from the probability mass function (pmf) $P(X=x|n,1\text{-}\theta)=0.01479$, but the interest usually lies with the cumulative distribution.
-We focus on the case where $X\leq x$, but it is straightforward to substitute $P(X\leq x)$ for $P(X=x)$ in the following.
-If we fix $x$ we can define the experiment described above as a Bernoulli trial, where success is defined as $X\leq x$, with corresponding probability $P(X\leq x)$, which we from now on write as $P_{x}$ for ease of notation.
-If we perform the coin-flip experiment $m$ times, and success is defined as $X\leq x$, we have a series of $m$ Bernoulli trials, and the number of successes follow the binomial distribution $\mathcal{B}(m,P_{x})$.
-Let $C_{x}$ be the random variable that denotes the number of successful experiments.
-The probability of, for example, $C_{x}=10$ experiments with at most $x=5$ tails, is calculated from the pmf $P(C_{x}=10|m,P_{x})$.
+Consider an experiment where a fair coin is flipped $n=20$ times and the number of successes is the number of times a classifier correctly predicts the outcome.
+Let the random variable $X$ denote the number of failures, $X\sim\mathcal{B}(n,1\text{-}\theta)$, where $\theta=0.5$.
+An unbiased estimator for $\theta$ is ${\hat{\theta}(X)=(n-X)/n}$.
+The corresponding estimate is referred to as accuracy.
+The probability of, say, $18$ correct predictions or more corresponds to $x=n-18$ failures or less, that is, the cumulative distribution $P(X\leq x|n,1\text{-}\theta)$.
+This corresponds to the cumulative distribution for the estimator $\hat{\theta}(X)=\frac{n\text{-}X}{n}$, and we have that $P(\hat{\theta}(X)\geq\frac{n\text{-}x}{n})=P(X\leq x)=0.00020$ for $\hat{\theta}\geq 0.90$.
+In other words, if $\theta=0.5$, there is a probability of $0.00020$ that the accuracy is $0.90$ or above in a series of $n=20$ trials.
+We can define the experiment described above as a Bernoulli trial, where success is defined as $X\leq x$, with corresponding probability $P(X\leq x)$, which we from now on write as $P_{x}$ for ease of notation.
+If we have $m$ classifiers, we have a series of $m$ Bernoulli trials, and the number of successes follow the binomial distribution $\mathcal{B}(m,P_{x})$.
+Let $C_{x}$ be the random variable that denotes the number of successful experiments, that is, the number of classifiers that correctly predicts at least $n-x$ coin flips.
 Of special interest is the case where at least one experiment succeeds, $P(C_{x}>0)=1-P(C_{x}=0)$.
-In our example, if the coin-flip experiment is performed $m=100$ times, the probability of observing $15$ heads or more at least once is $P(C_{x}>0|m,P_{x})=1-P(C_{x}=0|m,P_{x})$, with $P_{x}=0.02069$ from above.
-In the estimation context, we have that although the probability of $\hat{\theta}\geq 0.75$ is only $0.02069$ for the individual experiments, the probability of at least one $\hat{\theta}\geq 0.75$ is $0.8765$ when the experiment is performed $m=100$ times.
-Consider an experiment where a coin is flipped $n=20$ times, and the number of successes is the number of times a classifier correctly predicts the outcome.
-An unbiased estimator for a classifier’s probability of correct prediction, $\theta$, is ${\hat{\theta}=(n-X)/n}$, where $X$ is the number of failures.
-The corresponding estimate is referred to as accuracy.
-The true state-of-the-art for coin-flip classifiers is $\theta_{SOTA}=0.5$, since it is impossible to train a classifier to predict coin flips.
-If $m=1,000$ classifiers independently give their predictions for $n=20$ flips, we can consider this as a binomial experiment with $1,000$ Bernoulli trials, where the probability of success is the individual classifier’s probability of $n-X\leq n-x$ correct predictions.
-The probability of at least one classifier achieving an accuracy of at least $0.90$ is $0.18231-P(C_{x}=0|m,P_{x})=$ 0.1823, where $P_{x}=P(X\leq 18|n,1\text{-}\theta)=0.00020$.
-This means that using the top-ranked method as an estimate and ignoring multiplicity there is a substantial risk of overestimating the state-of-the-art.
-3.3 The probability distribution of $\hat{\theta}^{\,max}_{SOTA}$
-We define the SOTA probability of correct prediction, $\theta_{SOTA}$, as the highest probability of correct prediction among all classifiers applied to a test set, in accordance with the common set-up for public competition.
-Of interest is $P(\hat{\theta}^{\,max}_{SOTA}|m,n,\theta)$, the probability of $\hat{\theta}^{\,max}_{SOTA}$, being the estimated state of the art from taking the highest ranked classifier, given the number of classifiers, the size of the test set, and the probabilities of correct prediction, $\theta=\{\theta_{1},\ldots,\theta_{m}\}$.
-We will in the following assume independent classifiers, and discuss dependency in Section 4.
-We demonstrate the impact of multiplicity in the simple case of identical probabilities of correct prediction, $\theta_{j}=\theta$ for all classifiers, $j=1,\ldots,m$, and where $\theta_{SOTA}=\theta$.
-We discuss non-identical probabilities in Section 4.2.
-Let $n$ denote the size of the test set, $m$ be the number of classifiers, and $\theta$ be the probability of correct prediction.
-Let $X$ be the random variable that denotes the number of failures (wrong predictions) for a given classifier.
-We have that $X\sim\mathcal{B}(n,1\text{-}\theta)$.
-We denote by $P_{x}$ the cumulative distribution $P(X\leq x)$.
-The random sample, $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ consists of the number of wrong predictions for each classifier $j=1,\ldots,m$ on a test set.
+The true SOTA for coin-flip classifiers is $\theta_{SOTA}=0.5$, since it is impossible to train a classifier to predict coin flips.
+If $m=1,000$ classifiers independently give their predictions for $n=20$ flips, the probability of at least one classifier achieving an accuracy of at least $0.90$ is $0.181-P(C_{x}=0|m,P_{x})=$ 0.18.
+This means that using the top-ranked method as an estimate and ignoring multiplicity there is a substantial risk of grossly overestimating the SOTA.
+The same mechanism of multiplicity applies when $m$ teams submit their classifier’s result on a test set, although in a more complex manner.
+We seek to describe the probability distribution for increasing levels of complexity, and by that offer a tool to avoid overestimation.
+We will in the following use the public competition as a general example, and refer to the teams’ classifiers simply as classifiers, for ease of notation.
+3.3 The probability distribution of $\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})$
+We define the SOTA probability of success, $\theta_{SOTA}$, as the maximum among the $m$ classifiers being evaluated, $\theta_{SOTA}=\max(\Theta)$, $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$.
+Oftentimes, we only have access to sample estimates, $\hat{\Theta}=\{\hat{\theta}_{1},\ldots,\hat{\theta}_{m}\}$.
+Let $\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})=\max(\{\hat{\theta}_{1}(X),\hat{\theta}_{2}(X),\ldots,\hat{\theta}_{m}(X)\})$ be the naive, sample maximum estimator for $\theta_{SOTA}$.
+In the coin flip classifier example above, we have that $\theta_{SOTA}=0.5$, but $P(\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})\geq 0.90)=0.18$, and it is clear that we risk an overoptimistic estimate.
+Of interest is $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})|m,n,\Theta)$, which we will denote by $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ for ease of notation.
+Similarly, we use $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for $\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})$ when there is no risk of confusion.
+We here give $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ in the simple case of independent classifiers with identical probabilities of success: $\theta_{j}=\theta$ for all classifiers, $j=1,\ldots,m$.
+We discuss non-identical probabilities and dependency in Section 4.
+Let $X$ be the random variable that denotes the number of failures (wrong predictions) for a given classifier, with $X\sim\mathcal{B}(n,1\text{-}\theta)$.
+The random sample, $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ consists of the number of wrong predictions for each classifier $j=1,\ldots,m$ on a sample.
 Let $C_{x}$ be the random variable that denotes the number of classifiers with at most $x$ failures.
-We have that $C_{x}\sim\mathcal{B}(m,P_{x})$, because each classifier will independently have a probability $P_{x}$ of having at most $x$ wrong predictions.
-We are interested in the probability of at least one classifier having at most $x$ wrong predictions, because it corresponds to the probability function of the $\hat{\theta}^{\,max}_{SOTA}$ estimator, $P(\hat{\theta}^{\,max}_{SOTA}>\frac{n-x}{n})$, in other words, at least one $\hat{\theta}_{j}>\frac{n-x}{n}$.
-We can derive this from $C_{x}$:
-$P(C_{x}>0|m,n,\theta).$ (1)
-Note that this corresponds to the cumulative distribution function (cdf) $F(z)$ for the random variable $Z$, denoting the most wrong predictions among all classifier.
-We have cdf $F(z)$, the probability that the most wrong predictions among all classifiers is less than or equal to a number $z$, can easily be derived from the binomial distribution
+We have that $C_{x}\sim\mathcal{B}(m,P_{x})$.
+We are interested in the probability of at least one classifier having at most $x$ wrong predictions, because it corresponds to the probability function of the $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ estimator, $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}}\geq\frac{n-x}{n})$.
+We can derive this from $P(C_{x}>0|m,n,\theta)$.
+This corresponds to the cdf for the random variable $Z$ denoting the most wrong predictions among all classifiers:
 $F(z)$ $=P(C_{z}>0|m,n,\theta)=1-P(C_{z}=0|m,P_{z})$
 $=1-(1-P_{z})^{m}.$ (1)
-Let $f(z)$ denote the probability mass function of $Z$. We then have
+Let $f(z)=P(Z=z)$ denote the pmf of $Z$, the probability of at least one classifier having exactly $z$ wrong predictions, and no classifiers having fewer than $z$ wrong predictions. We then have
 $f(z)$ $=F(z)-F(z-1)=1-(1-P_{z})^{m}-\{1-(1-P_{z-1})^{m}\}$
@@
 $=-\sum_{k=1}^{m}(-1)^{k}\binom{m}{k}(1-\sum_{\ell=0}^{z-1}P(Z=\ell))^{m-k}P(Z=z)^{k}.$ (2)
-The last line is derived by binomial expansion.
-The practical relevance of the binomial expansion is limited by the need to approximate binomial coefficient for large $m$.
-From Eq.
-3.3, the expected value and variance can easily be calculated.
-An example of $F(z)$ can be seen in Fig.
-2, and the corresponding $f(z)$ in Fig.
-3.
-The pmf $f(z)=P(Z=z)$ denotes the probability of at least one classifier having exactly $z$ wrong predictions, and no classifiers having fewer than $z$ wrong predictions.
+The variable $Z$ corresponds to $\hat{\theta}_{SOTA}^{\,max}$ in the sense that the smallest number of failures among all classifiers corresponds to the sample maximum probability of success.
+Equations 3.3 and 3.3 can be used to analyse $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for a given situation, as we do in the next section, where the impact that the three parameters have on the bias is also presented.
 3.4 A simulated public competition example
-The multiplicity mechanisms described above apply to public competitions where several teams submit their predictions on a test set.
-Winning teams are those with highest accuracy, $\max_{j}\hat{\theta}_{j}$, where $\hat{\theta}_{j}=(n-x_{j})/n$ is the accuracy of a classifier $j$ on the test set.
-This is often referred to as the state of the art, and hence, we have that $\hat{\theta}^{\,max}_{SOTA}(\mathbf{x})=\max_{j}\hat{\theta}_{j}$.
-As exemplified in Section 3.2, multiplicity biases the estimation of $\theta_{SOTA}$, and Section 3.3 gives us the tools to analyze its impact.
-In the following, we will use a significance level of $\alpha=0.05$ and two-sided confidence intervals.
-Consider a classification problem with a test set of size $n=3,000$, and a classifier with probability of correct prediction $\theta=0.90$.
-The pmf of a single classifier is shown as a histogram in Fig.
+The multiplicity mechanism that we described in the coin-flip example applies to public competitions where multiple teams submit the predictions made by their classifiers on a test set.
+The classifier with highest performance is declared winner, and is often referred to as SOTA later on.
+As exemplified in Section 3.2, multiplicity biases the naive sample maximum $\theta_{SOTA}$-estimator, and Section 3.3 gives us the tools to analyze its impact.
+In the following, we will use a significance level of $\alpha=0.05$ and two-sided CIs.
+Consider a classification problem with a test set of size $n=3,000$, and a classifier with probability of success $\theta=0.90$.
+The pmf of a single $\hat{\theta}(X)$ is shown as the blue curve in Fig.
+1, where the dotted line indicates $E(\hat{\theta})=\theta$.
+The 95% CI for $\hat{\theta}=0.90$ is $(0.8887,0.9105)$, and we denote it by $(\theta_{\alpha/2},\theta_{1\text{-}\alpha/2})$, displayed as a blue line on the horizontal axis.
+If there are $m=1,000$ independent classifiers, each with probability of success $\theta=0.90$, the probability of at least one classifier achieving at least $\hat{\theta}=\theta_{1\text{-}\alpha/2}$ is close to 1.
+This comes as no surprise given the definition of a CI.
+It means that we can be pretty certain that the sample performance of the top ranked classifier is above the upper limit of the CI of the expected performance.
+Consider a new classifier with probability of success $\theta_{m+1}=\theta_{1\text{-}\alpha/2}$.
+The corresponding pmf is displayed as the red curve in Fig.
+1, with the dotted line indicating $E(\hat{\theta}_{m+1})=\theta_{1\text{-}\alpha/2}$.
+By common scientific standards, the expected performance is considered significantly better than the expected performance of any of the $1,000$ classifiers with $\theta=0.90$, and is of interest to the research community.
+However, the multiplicity effect can conceal this.
+In the competition context, this is not necessarily a problem, since each classifier has the same probability of achieving the highest sample performance given that they have the same $\theta$.
+In a scientific context, using $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as an estimate for $\theta_{SOTA}$ without multiplicity corrections is problematic.
+The upper limit of the $(1\text{-}\alpha)$ CI of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ is $0.9213$, displayed on the horizontal axis in Fig.
+1, denoted as $\theta^{m}_{1\text{-}\alpha/2}$.
+We have $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})=0.9173$, calculated from Eq.
+3.3, also displayed on the horizontal axis in Fig.
 1.
-The 95% confidence interval for $\theta$ if we have observed $\hat{\theta}=0.90$ is $(0.8887,0.9105)$, and we denote it by $(\theta_{\alpha/2},\theta_{1\text{-}\alpha/2})$.
-If there are $m=1,000$ classifiers, each with probability of correct prediction $\theta=0.90$, the probability of at least one team achieving at least $\theta_{1\text{-}\alpha/2}$ is close to 1.
-This comes as no surprise given the definition of a confidence interval.
-It means that we can be almost certain that the accuracy of the top ranked classifier, which is the default SOTA, is above the upper limit of the confidence interval of the true performance.
-On the other hand, it is straightforward to show that if each individual classifier has a probability of $P_{x}=0.00002532$ for at most $x$ failures, then, with a probability of $\alpha/2$, at least one team will achieve at least $\hat{\theta}=0.9213$.
-We denote this by $\theta^{m}_{1\text{-}\alpha/2}$, the multiplicity-adjusted upper limit of the $(1\text{-}\alpha)$ confidence interval.
-Similarly, the expected value of $\hat{\theta}^{\,max}_{SOTA}$ can easily be calculated from Eq.
-3.3, and we have $E(\hat{\theta}^{\,max}_{SOTA})=0.9173$.
-Both are optimistic estimates, knowing that $\theta_{SOTA}$ is $0.90$.
-Within the competition context, this is not necessarily a problem, because each team has the same probability of achieving the highest accuracy given that they have the same $\theta$.
-In a scientific context, using $\max_{j}\hat{\theta}_{j}$ as an estimate for $\theta_{SOTA}$ without multiplicity corrections is problematic.
-Consider a new classifier with probability of correct prediction $\theta_{m+1}=\theta_{1\text{-}\alpha/2}$ from above, displayed as the red curve in Fig.
+The probability of the new classifier beating the upper limit of the sample maximum CI is very low: $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})=0.02$, illustrated by the darker green area in Fig.
 1.
-By common scientific standards, the expected performance, $E(\hat{\theta}_{m+1})=\theta_{m+1}$ is considered significantly better than the expected performance of any of the $1,000$ classifiers with $\theta=0.90$, and is of interest to the research community.
-But the probability of this classifier beating the upper limit of the multiplicity adjusted confidence interval is very low: $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})=0.0190$, and can be seen as the red area in Fig.
+Similarly, $P(\hat{\theta}_{m+1}\geq E(\operatorname{\hat{\theta}^{\,max}_{SOTA}}))=0.10$ seen as the entire green area in Fig.
 1.
-Similarly, $P(\hat{\theta}_{m+1}\geq E(\hat{\theta}^{\,max}_{SOTA}))=0.0996$, seen as the blue plus red area in Fig.
-1.
-This means that by using the maximum accuracy as SOTA, we can expect that 9 out of 10 classifiers that are in fact better will still be naively considered worse.
-The probability mass functions of single $\hat{\theta}$.
-Figure 1: The grey histogram shows the pmf of a classifier with probability of correct prediction $\theta$.
-The red curve is the pmf of a classifier with probability of correct prediction $\theta_{1\text{-}\alpha/2}$, the upper limit of the $95\%$ CI for $\hat{\theta}(\mathbf{x})=\theta$.
-$E(\hat{\theta}^{\,max}_{SOTA})$ denotes the expected value of the highest ranked among $m=1,000$ classifiers.
-$\theta^{m}_{1\text{-}\alpha/2}$ is the multiplicity corrected upper limit of the $95\%$ CI.
-The shaded areas corresponds to $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})$ (red) and $P\big{(}\hat{\theta}_{m+1}\geq E(\hat{\theta}^{\,max}_{SOTA})\big{)}$ (blue+red).
-The cumulative distribution function, $F(z)$ from Eq.
-3.3, is displayed in Fig.
-2, and describes the cumulative probability of at least one classifier having at most $z$ failures.
-Or, in other words, the cumulative probability of maximum accuracy.
-The number of failures are displayed on the top horizontal axis, and the corresponding accuracy on the bottom horizontal axis.
+This means that by using the sample maximum as an estimator for SOTA, we can expect that 9 out of 10 classifiers that are significantly better will still be naively considered worse.
+The pmfs of two single $\hat{\theta}(X)$ compared to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ statistics
+Figure 1: Against a mass of identical classifiers, a single significantly better classifier has little chance of beating the state of the art: The blue and the red curves show the pmfs of $\hat{\theta}(X)$ for classifiers with probability of success $\theta$ and $\theta_{1\text{-}\alpha/2}$, respectively.
+$\theta_{1\text{-}\alpha/2}$ is the upper limit of the CI of $\hat{\theta}(\mathbf{x})=\theta$.
+The CI is displayed as the blue horizontal line.
+$E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ and $\theta^{m}_{1\text{-}\alpha/2}$ denote the expected value and the upper limit of the CI of the sample maximum estimator for $m=1,000$ classifiers.
+The shaded areas corresponds to $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})$ and $P\big{(}\hat{\theta}_{m+1}\geq E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})\big{)}$, and reflects the slim chances that the single better classifier has of beating the multiple worse classifiers.
+For the entire distribution of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, see Fig. 2.
+Cdf and pmf for at least one classifier having at most $z$ failures, corresponding to the sample maximum estimator
+(a) Cdf $F(z)$, see Eq. 3.3
+(b) Pmf $f(z)$, see Eq. 3.3
+Figure 2: The distribution and the probability function of the sample maximum estimator for $n=3,000$, $\theta=0.90$, and $m=1,000$.
+The horizontal axes denotes the number of failures, $z$, at the top and the corresponding sample maximum estimates, $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, at the bottom.
 Note that the horizontal axis is reversed compared to Fig.
 1.
-Cumulative distribution function, $F(z|m,n,\theta)$, for at least one classifier having at most $z$ failures.
-Figure 2: The figure shows the cumulative distribution function for at least one classifier having at most $z$ failures.
-The horizontal axis denotes the number of failures, $z$, on top and the accuracy at the bottom.
-The corresponding pmf, $f(z)$ from Eq. 3.3, is displayed as a histogram in Fig. 3.
-Probability mass function, $f(z|m,n,\theta)$, for the smallest number of failures, $z$, among all classifiers.
-Figure 3: The figure shows the corresponding pmf for Fig. 2. The horizontal axis denotes the number of failures, $z$, on top and the corresponding accuracy at the bottom.
-From Eq.
-3.3 we get the expected value and variance.
-For $m=1,000$, $n=3,000$ and $\theta=0.9$ this corresponds to $E(\hat{\theta}^{\,max}_{SOTA})=0.9173$ with a standard deviation of $0.001817$.
-We can see how different values for $m$, $n$ and $\theta$ affects the $\hat{\theta}^{\,max}_{SOTA}$ distribution in Table 1.
-Expected values and standard deviations for $\hat{\theta}^{\,max}_{SOTA}$
-m n $\theta$ $E(\hat{\theta}^{\,max}_{SOTA})$ $\sigma_{\hat{\theta}^{\,max}_{SOTA}}$
-1,000 3,000 0.85 0.8707 0.002197
+The distribution of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ will vary with $m$, $n$ and $\theta$, and Table 1 shows the expectation and standard deviation for a small collection of examples.
+To complement the table, Fig.
+3 a)-c) show the bias of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as a function of $m$, $n$ and $\theta$, respectively.
+Expected values and standard deviations for $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$
+m n $\theta$ $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ $\sigma_{\operatorname{\hat{\theta}^{\,max}_{SOTA}}}$
 1,000 3,000 0.90 0.9173 0.001817
-1,000 3,000 0.95 0.9624 0.001277
 100 3,000 0.90 0.9135 0.002250
-500 3,000 0.90 0.9163 0.001923
+5,000 3,000 0.90 0.9196 0.001623
 1,000 1,000 0.90 0.9294 0.003007
 1,000 10,000 0.90 0.9096 0.001022
-Table 1: The expectation and standard deviation for the SOTA estimator with varying test set size, number of classifiers and probability of correct prediction.
+1,000 3,000 0.85 0.8707 0.002197
+1,000 3,000 0.95 0.9624 0.001277
+Table 1: The expected value and standard deviation for the SOTA estimator with varying test set size, number of classifiers and probability of correct success.
 The bold line is the example used in Fig.
+2 and corresponds to the intersection between the gray vertical lines and the solid lines in Fig.
 3.
-4 Non-i.i.d.
-The probability distribution in Section 3.4 is based on independent classifiers with identical $\theta$s.
-This assumption is rarely true in practice.
-A classifier’s ability to perform better than random guesses is due to correlation between features and labels in a training set.
-The teams will pick up on many of the same features and correlations, regardless of which method they choose.
-In addition, many teams will use different versions of related methods.
-We investigate dependency and non-identical classifiers, and the impact on the bias.
-Table 2 gives an overview, details are in the following sections.
-Upper bounds, expected values and standard deviations for $\hat{\theta}^{\,max}_{SOTA}$
-section independent identical upper 95% CI expected value standard deviation
-3.4 YES YES 0.9213 0.9173 0.001817
-4.1 NO YES 0.9207 0.9140 0.003481
-4.1 NO YES 0.9173 0.9140 0.001484
-4.2 YES NO 0.9177 0.9130 0.002129
-4.3 NO NO 0.9173 0.9101 0.003649
-Table 2: The upper bound of the 95% CI, the expectation and standard deviation for the SOTA estimator for different combinations of independency and identicality.
-The true state-of-the-art is $\theta=0.90$.
-4.1 Dependent, identical classifiers
-Let $Y_{j}$ be the Bernoulli variable with value $1$ if classifier $j$ predicts a data point correctly.
-We have $P(Y_{j}=1)=\theta_{j}$, and $\theta_{j}=\theta$ for classifiers with equal probability of success.
-Mania et al.
-(2019) calculated the similarity, defined as the fraction of two models giving the same output, which is equivalent to
-$P\big{(}\{Y_{j},Y_{k}\}=\{0,0\}\big{)}+P(\{Y_{j},Y_{k}\}=\{1,1\})$ (4)
-$=(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}$
-for independent models. For dependent models with correlation $\rho_{jk}$, this becomes
-$(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}+2\rho_{jk}\sqrt{\theta_{j}\theta_{k}(1-\theta_{j})(1-\theta_{k})}.$ (5)
-By setting $\theta_{j}=\theta$ and $\rho_{jk}=\rho$, $\rho$ can easily be estimated from the estimations of $\theta$, $P\big{(}\{Y_{j},Y_{k}\}=\{0,0\}\big{)}$ and $P(\{Y_{j},Y_{k}\}=\{1,1\})$.
-In Mania et al.
-(2019), they used the average accuracy for top-performing ImageNet models, $\hat{\theta}=0.756$.
-For details, see Mania et al.
-(2019).
-Note that these are multi-class problems, so a lower accuracy than for binary classification problems is to be expected.
-From their calculated similarity of $0.85$, the estimated correlation is $0.36$.
+For each curve, the two other parameters are kept constant, and each figure displays five curves, corresponding to Table 1.
+The bias of the sample maximum estimator
+Figure 3: The bias of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as a function of each of its three parameters; number of classifiers, test set size and probability of success.
+The intersection between the black solid lines and the gray vertical lines correspond to the bold line in Table 2.
+The dotted and dashed lines correspond to variation of parameter values found in Table 2, colour-coded accordingly.
+In Fig.
+3(a), we see that the impact of number of classifiers is dramatic, and multiplicity-considerations must be treated with full attention starting at $m=2$.
+We can see this from the cdf in Eq.
+3.3, where $m$ is the exponent.
+In Fig.
+3(b) is that the test set size also has large impact on the bias.
+With small test sets, the variance of each individual $\hat{\theta}_{j}(X)$ is high, increasing the probability of sample estimates far from $\theta$.
+As seen in Fig.
+3(c), the bias decreases as $\theta$ approaches 1.
+The variance of the individual $\hat{\theta}_{j}(X)$s is at its highest for $\theta=0.5$ and decreases monotonically, and it is this variance that is reflected in the bias, similar to the effect of the test set size.
+The change in absolute numbers may seem modest, but the practical implications are large, as demonstrated in Section 5, where we analyse the results from three public competitions.
+4 Dependency and non-identical $\theta$s
+The probability distribution in Section 3.4 is that of independent classifiers with identical $\theta$s, an assumption that is rarely true in practice.
+We expand our investigation to non-identical $\theta$s and dependency.
+Table 2 and Figure 5 give an overview of some exemplifying modelss, details are in the following sections.
+Characteristics of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for dependent and non-identical classifiers
+Section Independent Identical Upper 95% CI $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ $\sigma_{\operatorname{\hat{\theta}^{\,max}_{SOTA}}}$
+3.4 YES YES 0.9213 0.9173 0.0018
+4.1 YES NO 0.9177 0.9130 0.0021
+4.2 NO YES 0.9207 0.9140 0.0035
+4.3 NO NO 0.9173 0.9101 0.0036
+Table 2: The upper bound of the 95% CI, the expected value and standard deviation for the sample maximum estimator for different combinations of independency and identicality.
+The true SOTA is $\theta=0.90$, the test set size is $n=3,000$, the number of classifiers is $m=1,000$.
+The correlation coefficient and the $\Theta$-vector are found in their respective subsections.
+Bias for non-identical and dependent classifiers.
+(a) Bias as a function of $\min(\mathbf{\Theta})$, with constant $\max(\mathbf{\Theta})=0.90$
+(b) Bias as a function of correlation coefficient.
+Figure 4: Bias of the sample maximum estimator as a function of $\min(\Theta)$ for non-identical classifiers and of $\rho_{0}$ for conditional independent classifiers.
+(a) The right tip of the curve is identical classifiers, and following the curve from right to left the $\mathbf{\Theta}$-vector stretches out, and the bias decreases.
+The intersection with the solid gray vertical line corresponds to the cdf and pmf in Fig.
+5, and produces the CI, the expectation and the standard deviation in line two in Table 2.
+The intersections between the gray vertical lines correspond to the left tips of the black lines in Fig.
+4(b), where the classifiers are independent.
+(b) The red line are identical classifiers, as discussed in Section 4.2.
+The left tip of the red curve corresponds to the i.i.d.
+classifiers in Section 3.4, and the intersection with the grey vertical line corresponds to line three in Table 2.
+The black lines show bias as a function of correlation coefficient demonstrated for different $\mathbf{\Theta}$-vectors.
+The intersection between the solid black curve and the grey vertical line corresponds to line four in Table 2.
+The scale of the $y$-axis is kept the same as in Fig.
+3 for easier comparison of which parameters have the most impact.
+Overall, Table 2 demonstrates that dependency and non-identical $\theta$s reduces the bias of the $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ estimator, and assumptions regarding dependency and identicality are crucial to include in a model.
+Figure 4 shows how the bias changes as a function of $\Theta$ and $\rho_{0}$ for these specific models.
+We use the set-up from Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, $\theta_{SOTA}=0.90$, and $100,000$ repetitions when simulations are needed.
+We refer to classifiers with identical $\theta$s as identical classifiers.
+4.1 Non-identical, independent classifiers
+We are interested in generalising Eq.
+3.3 to the non-identical case, and we start by noting that we now have $P_{xj}=P(X\leq x|\theta_{j})$, and we seek the sum of those so that we can find the proper expression for
+$F(z|\mathbf{\Theta})=1-P(C_{z}=0|m,n,\mathbf{\Theta}).$
+The Poisson Binomial distribution, see, e.g., Wang1993, describes the sum of $m$ non-identical independent Bernoulli trials with success probabilities $p_{1},p_{2},\ldots,p_{m}$.
+The pmf for $K$ number of successes is
+$P(K=k)=\sum\limits_{{A\in G_{k}}}\prod\limits_{{i\in A}}p_{i}\prod\limits_{{j\in A^{c}}}(1-p_{j}),$ (3)
+where $G_{k}$ is all subsets of size $k$ that can be selected from {1, 2, …, m}.
+We then get the cdf for non-identical classifiers
+$F(z|\mathbf{\Theta})$ $=1-\sum\limits_{{A\in G_{0}}}\prod\limits_{{i\in A}}P_{zi}\prod\limits_{{j\in A^{c}}}(1-P_{zj})=1-\prod\limits_{j=1}^{m}\big{(}1-P_{zj}\big{)}.$ (4)
+The pmf is
+$f(z|\mathbf{\Theta})$ $=\prod\limits_{j=1}^{m}\big{(}1-P_{z\text{-}1j}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-P_{zj}\big{)}$ (5)
+$=\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z\text{-1}}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}.$
+Analytically, it is more complex than the case of identical classifiers, but it is easy to simulate.
+We use the same parameters as in Section 3.4, replacing $\theta=0.90$ with the vector $\mathbf{\Theta}$, where $\min(\mathbf{\Theta})=0.875,\,\max(\mathbf{\Theta})=0.90$ and equally spaced elements.
+The cdf and pmf are displayed in Fig.
+5(a) and Fig.
+5(b).
+Cdf and pmf for non-identical classifiers.
+(a) Cdf $F(z)$
+(b) Pmf $f(z)$
+Figure 5: The distribution and the probability function of the sample maximum estimator for non-identical $\theta$s, where $\min(\mathbf{\Theta})=0.875,\,\max(\mathbf{\Theta})=0.90$ and equally spaced elements.
+The upper bound of the 95% CI is $0.9177$.
+The $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ is $0.9129$, with standard deviation of $0.0021$.
+Comparing this to the results in Section 3.4, we see that the bias is smaller and the variance larger.
+This is because the classifiers at the lower end does not contribute to the sample maximum, and the reduced number of contributing classifiers gives a higher variance.
+The exact reduction/increase in bias and variance depends on $\mathbf{\Theta}$, and can at its most extreme cases completely eradicate the multiplicity effect, as we shall see in Section 5.1.2.
+Figure 4(a) shows how $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ changes with $\min(\mathbf{\Theta})$ for equally spaced elements with $\max(\mathbf{\Theta})=0.90$.
+What we see here is that with non-identical classifiers the bias decreases rapidly at first as $\min(\mathbf{\Theta})$ moves away from $\max(\mathbf{\Theta})$, but the rate at which this happens also decreases rapidly.
+This means that taking into account that classifier performances are heterogeneous is important, and small variations from identicality can have great impact on SOTA estimation.
+In Section 5, we demonstrate for other $\mathbf{\Theta}$s, based on observations.
+4.2 Identical, dependent classifiers
+Even though the different teams work independently, their resulting classifiers are not necessarily statistically independent.
+A classifier’s ability to perform better than random guesses is due to the correlations between features and labels in the training set.
+In their training phase, different classifiers will pick up on many of the same features and correlations, regardless of which classifier they train.
+In addition, many teams will use different versions of related models, often pre-trained on the same data set.
+This scenario can be modelled by conditional independency.
+Let $Y_{0}$ be a Bernoulli variable with $P(Y_{0}=1)=\theta_{0}$.
+Let $Y_{j}$ be the Bernoulli variable with value $1$ if classifier $j$ predicts a data point correctly with $P(Y_{j}=1)=\theta_{j}$, and let $j=1,\ldots,m$ be the $m$ classifiers contributing to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$.
 Boland et al.
-(1989) describes dependency in the majority systems context, with $\rho_{0j}=\text{corr}(Y_{0},Y_{j}),\,j=1,\ldots m$, and the $Y_{j}$s are independent given $Y_{0}$.
-This gives $\rho_{jk}=\rho_{0j}^{2}$.
-The observations from $Y_{0}$ does not contribute directly to $\hat{\theta}^{\,max}_{SOTA}$, but can be seen as, e.g., the influence that a common training set has on the classifiers.
+(1989) described dependency in the majority systems context, where $\rho_{0j}=\text{corr}(Y_{0},Y_{j}),\,j=1,\ldots m$, and the $Y_{j}$s are independent given $Y_{0}$.
+The observations from $Y_{0}$ does not contribute directly to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, but can be seen as, e.g., the influence that a common training set has on the classifiers.
 The binomial distribution for dependent Bernoulli trials is not an easy problem, see, e.g., Ladd (1975) and Hisakado et al.
 (2006).
-We here demonstrate by simulation how dependency reduces the bias.
-We use the estimated correlation from Mania et al.
-(2019), which corresponds to $\rho_{0j}=0.6$ in the multiplicity set-up with an external $Y_{0}$.
-We use the set-up from Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, and probability of correct prediction $\theta=0.90$.
-We simulate two scenarios; one where $Y_{0}$ is random, and one where the $y_{0}$’s are fixed with exactly $(n-x)/n=0.90$ correct predictions.
-The fixed scenario can correspond to fixed training conditions, e.g., models pre-trained on ImageNet.
-The random scenario picks up the randomness in, e.g., a training set.
-Although all teams have the same training set, the training set is a random sample.
-The upper bound of the $95\%$ confidence interval for the dependent classifiers is at $0.9173$ for the fixed $y_{0}$ and $0.9207$ for the random $Y_{0}$, but they have the same expected value.
-Minimum number of failures in $m=1,000$ classifiers.
-(a) $\rho_{0j}=0.6$, random $Y$
-(b) $\rho_{0j}=0.6$, fixed $y$
-Figure 4: The minimum number of failures in $100,000$ simulations, with probability of success $\theta=0.9$.
-Note that, with $\theta=0.90$, the similarity as described by Mania et al.
+The Bernoulli trials can be simulated, see Boland et al.
+(1989) for conditional distributions.
+We give an example below for the case where $\rho_{0j}=\rho_{0}$ and $\theta_{j}=\theta$ for all classifiers.
+We set $\rho_{0}$ based on the work of Mania et al.
+(2019).
+Mania et al. (2019) calculated model similarity, defined as the probability of two classifiers giving the same output, which is equivalent to
+$(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}+2\rho_{jk}\sqrt{\theta_{j}\theta_{k}(1-\theta_{j})(1-\theta_{k})}$ (6)
+for dependent models with correlation $\rho_{jk}$.
+The average accuracy for top-performing ImageNet models, $\hat{\theta}=0.756$, and the calculated similarity of $0.85$ gives $\rho=0.36$ if we assume the same correlation for all classifiers.
+We have that $\rho_{jk}=\rho_{0j}^{2}$, and hence our proposed correlation $\rho_{0}$ is set to $0.6$.
+Note that, for $\theta=0.90$, the similarity as described by Mania et al.
 (2019) is $0.82$ for two independent classifiers.
-When the probability of correct prediction is high, there is little room for variation between classifiers, and the inherent similarity is high.
-4.2 Non-identical, independent classifiers
-The Poisson Binomial distribution, see, e.g., Wang (1993), describes the sum of $n$ non-identical independent Bernoulli trials with success probabilities $p_{1},p_{2},\ldots,p_{n}$.
-The pmf for $K$ number of successes is
-$P(K=k)=\sum\limits_{{A\in G_{k}}}\prod\limits_{{i\in A}}p_{i}\prod\limits_{{j\in A^{c}}}(1-p_{j}),$ (6)
-where $G_{k}$ is all subsets that can be selected from {1, 2, …, n}.
-In the multiple classifier context, we are interested in generalising Eq. 3.3 to the non-identical case.
-$F(z)=P(C_{z}>0|m,n,\mathbf{\theta})=1-P(C_{z}=0|m,n,\mathbf{\theta}),$ (7)
-where $P(C_{z}=0)$ corresponds to 00 classifiers with non-identical probabilities of success, $P(Z\leq z|\theta_{j})$. The cdf is
-$F(z)$ $=1-\sum\limits_{{A\in G_{0}}}\prod\limits_{{i\in A}}P(Z\leq z|\theta_{i})\prod\limits_{{j\in A^{c}}}(1-P(Z\leq z|\theta_{j}))$
-$=1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}.$ (8)
-The pmf is
-$f(z)$ $=F(z)-F(z-1)$
-$=1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}-\{1-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z\text{-}1|\theta_{j})\big{)}\}$
-$=\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z\text{-}1|\theta_{j})\big{)}-\prod\limits_{j=1}^{m}\big{(}1-P(Z\leq z|\theta_{j})\big{)}$ (9)
-$=\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z\text{-1}}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}$
-Analytically, it is more complex than the case of identical classifiers, but it is easy to simulate.
-We use a similar set-up as in Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, and probabilities of correct prediction $\theta=\{\theta_{1},\ldots,\theta_{m}\}$, where $\min\theta=0.875,\,\max\theta=0.90$, and they are equally spaced.
-The simulated non-identical upper bound of the 95% confidence interval is $0.9177$, with $100,000$ repetitions.
-The $E(\hat{\theta}^{\,max}_{SOTA})$ is $0.9129$, with standard deviation of $0.0021$.
-The cdf and pmf are displayed in Fig.
-5 and Fig.
-6.
-Cumulative distribution function for non-identical classifiers.
-Figure 5:
-Probability mass function for non-identical classifiers.
-Figure 6:
+When the probability of correct prediction is high, there is little room for variation between classifier outcomes.
+Let $P(Y_{j}=1)=\theta=0.90$ for all classifiers, and $P(Y_{0}=1)=\theta=0.90$.
+From our simulations, we get $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})=0.9140$, a decrease in bias compared to the independent and identical classifiers.
+However, the variance has increased due to the random nature of $Y_{0}$, so the upper 95% CI is essentially the same.
+See, Table 2 for exact numbers.
+To give an idea of how the randomness of $Y_{0}$ influences the variance, we fix the $y_{0}$’s with exactly $(n-x)/n=0.90$ correct predictions.
+The expected value is the same as with random $Y_{0}$s, but the standard deviation has now decreased from $0.0035$ to $0.0015$.
+In the real world, the dependencies have both random and fixed sources.
+A random source would, e.g., be the training set and the test set, random samples from an underlying distribution.
+A fixed source can be, e.g., fixed training conditions, such as pre-trained models.
+For $\rho_{0j}=0$, we have independent classifiers, the same scenario as in Section 3.4, and for $\rho_{0j}=1$, we have $m$ classifiers producing the same output, and multiplicity is not a concern.
+For the whole range, see the red curve in Figure 4(b).
+What we see here is that the correlation coefficient value in a model has a great impact on the estimated bias if the correlation is substantial.
+This can explain that in mature fields, where most competing teams will use variants of the same model, the multiplicity will not impact the overestimation of SOTA as much as in young fields where the competing teams have more diverse models.
+This aligns with what is often observed in science: overoptimistic results in the beginning, and more sober as the field evolves.
 4.3 Non-identical, dependent classifiers
@@
 An analytical solution is not known Hisakado et al.
-(2006), but simulations are easy.
+(2006), but simulations are straightforward.
 The results of Boland et al.
-(1989) are easily extended, and we have that
+(1989) can be extended, and we have that
 $P(Y_{j}|Y_{0}=1)$ $=\frac{\rho\sqrt{\theta_{j}(1-\theta_{j})\theta_{0}(1-\theta_{0})}+\theta_{j}\theta_{0}}{\theta_{0}}$ (7)
 $P(Y_{j}|Y_{0}=0)$ $=\frac{-\rho\sqrt{\theta_{j}(1-\theta_{j})\theta_{0}(1-\theta_{0})}+\theta_{j}(1-\theta_{0})}{(1-\theta_{0})}.$ (8)
-Note that, since $P\leq 1$, it follows from 11 that
+Note that, since $P\leq 1$, it follows from Eq. 8 that
 $\theta_{j}\geq\frac{\rho^{2}\frac{\theta_{0}}{1-\theta_{0}}}{1+\rho^{2}\frac{\theta_{0}}{1-\theta_{0}}},$ (9)
-and similarly for $0\leq P$ and 10.
-We get the a the simulated dependent upper bound of the 0.95 confidence interval of $0.91733$, with $100,000$ repetitions.
+and similarly for $0\leq P$ and Eq. 7.
+We get an upper bound of the $95\%$ CI of $0.91733$ and an expected value of $0.9101$, which are lower than both the dependent and the non-identical model separately.
+The standard deviation is $0.0036$, which is higher than the two other models.
+This sums of up much of Table 2, where more complex, and therefore more realistic, models reduces the bias but increases the variances.
+Figure 4(b) shows the impact of correlation for non-identical classifiers.
+What we see is that the bias decreases slowly for low correlations, but that the rate increases as $\rho_{0}$ increases.
+Importantly, Eq.
+9 restricts the discrepancy between $\min(\mathbf{\Theta})$ and $\max(\mathbf{\Theta})$, and the graphs are right-truncated.
+From Fig.
+3(a) we observe that the impact from number of classifiers flattens off, instead of restricting $\rho_{0}$ in our model, we can exclude classifiers by performance without large effect on the estimated bias.
+That is, restricting $\min(\mathbf{\Theta})$ instead of $\rho_{0}$.
+What we also observe in Fig.
+4(a) is that change in $\min(\mathbf{\Theta})$ has less impact on the lower end of the scale, so this also supports a decision of excluding teams with low performance without impacting the bias-estimation too much.
 5 Real world examples
-As we have shown above with simulated data, the observed maximum is an overoptimistic estimate for SOTA.
-These results are based on known state-of-the-art, but faced with real world data, the problem to be solved is of opposite nature: The observed $\hat{\theta}^{\,max}$ is known, but the true SOTA must be estimated.
+As we have shown above with simulated data, the sample maximum is an overoptimistic estimate for SOTA.
+These results are based on known SOTA, but faced with real world data the problem to be solved is of opposite nature: The sample $\max(\hat{\Theta})$ is known, but the true SOTA must be estimated.
 In the public competition or public data set context there are many factors that come in to play, making the bias calculations everything but straightforward.
@@
 To estimate multiplicity effect on the AUC scores is not as straightforward as with accuracy, but the mechanisms are the same and AUC is a more common ranking criterion than accuracy.
-In a public competition with $m$ participating teams, let $\hat{\theta}_{j}$, $j=1,\ldots,m$ be the estimated probability of correct prediction for each classifier based on the performance on an independent test set of size $n$.
-We will refer to these as the competition observations.
-$\hat{\theta}^{\,max}$ is the observed performance of the winning team.
+In a public competition with $m$ participating teams, let $\hat{\Theta}=\{\hat{\theta}_{1},\hat{\theta}_{2},\ldots,\hat{\theta}_{m}\}$, be the estimated probability of success for each classifier based on the performance on an independent test sample of size $n$.
+The the winning team is identified by the sample maximum, $\max(\hat{\Theta})$.
 Among the different scenarios presented in Section 4, the non-identical dependent is the most realistic model.
-Whereas in Table 2 we simulated $E(\hat{\theta}^{\,max})$ from known $\theta_{1},\ldots,\theta_{m}$, the task is now to estimate $\theta_{SOTA}$ based on $\hat{\theta}_{1},\ldots,\hat{\theta}_{m}$.
-A ’natural’ approach is to find $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ that gives $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$.
-That is, find a set of $\theta$’s so that the expected value of the observed maximum on a test set of size $n$ is equal to the competition observed maximum.
-We use a straightforward set-up, where the simulated $\hat{\tilde{\theta}}_{j}$ is the mean value of $n$ realisations of $Y_{j}$, where $P(Y_{j})=\tilde{\theta}_{j}$.
+Whereas in Table 2 we report $E(\hat{\theta}^{\,max}_{SOTA})$ from known $\Theta$, the task is now to estimate $\theta_{SOTA}$ based on $\hat{\Theta}$.
+A straightforward approach is to find $\Theta^{\prime}=\{\theta^{\prime}_{1},\theta^{\prime}_{2},\ldots,\theta^{\prime}_{m}\}$ that gives $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$.
+That is, find a set of $\theta$’s so that the expected value of its sample maximum on a test set of size $n$ is equal to the competition sample maximum.
+The SOTA estimate is then $\max(\Theta^{\prime})$.
+We use a set-up where the simulated $\hat{\theta}^{\prime}_{j}$s are the mean values of $n$ realisations of $Y_{j}$, where $P(Y_{j})={\theta}^{\prime}_{j}$.
 The simulation is repeated $10,000$ times, and for each repetition the observed maximum value is recorded.
-$E(\hat{\tilde{\theta}}^{\,\max})$ is the mean value over the repetitions.
-In the studies in Section 4, the $\theta_{1},\ldots,\theta_{m}$ were chosen to be equally spaced, but this does not align with the competition observations, as seen in Figures 7(a), 8(a) and 9.
-We base our suggested $\tilde{\theta}_{j}$’s on cropping the competition observations, as seen, e.g., in Figure 7(c), but this is far from the only approach.
-Shrinking, which we also investigated, might give a more realistic set of $\tilde{\theta}_{j}$’s, but it also requires an extra parameter.
+$E(\max(\hat{\Theta}^{\prime}))$ is the mean value over the repetitions.
+In the studies in Section 4, the $\theta_{1},\ldots,\theta_{m}$ were chosen to be equally spaced, but this does not align with the real world sample estimates, as seen in Figures 6(a), 7(a) and 8.
+We base our suggested ${\theta}^{\prime}_{j}$’s on cropping the sample estimates, as seen, e.g., in Figure 6(c), but this is far from the only approach.
+Shrinking, which we also investigated, might give a more realistic set of ${\theta}^{\prime}_{j}$’s, but it also requires an extra parameter.
 For the examples presented here shrinking gave similar results to cropping, and are therefore not included.
-An alternative to the expected value of the observed maximum is the upper limit of the 95% CI.
-The latter is more in line with common interpretation of the uncertainty of observed results for research, but which to choose will depend on the aim of the analysis.
+An alternative to the expected value of the sample maximum is the upper limit of the 95% CI.
+The latter is more in line with common interpretation of the uncertainty of observed results in research, but which to choose will depend on the aim of the analysis.
 We demonstrate the impact in the Obesity Risk example, reported in Table 3.
-Not every $\hat{\theta}^{\,max}$ is a result of multiplicity, and we demonstrate this in the Cassava Leaf example.
+Not every $\max(\hat{\Theta})$ is a result of multiplicity, and we demonstrate this in the Cassava Leaf example.
 5.1 Estimating $\theta_{SOTA}$
 5.1.1 Multi-Class Prediction of Obesity Risk
-Multi-Class Prediction of Obesity Risk is a competition in the 2024 Kaggle Playground Series, using a synthetic data set generated from the Obesity or CVD risk data set (10.34740/kaggle/dsv/7009925).
-The test set size was $n=13,840$, and the number of participating teams were $m=3,498$ (excluding teams with accuracy lower than chance).
-The best performing team had $\hat{\theta}^{\,\max}=0.9116$.
-Figure 7(a) shows the competition observations, $\hat{\theta}_{j}$’s, zoomed in at the $3,081$ best performing teams.
-The 95% (not multiplicity-corrected) CI for $\hat{\theta}^{\,max}$ is $(0.9067,0.9162)$ and is shown as the red line in Figure 7(a), the lower bound enveloping $920$ of the competition observations.
+Multi-Class Prediction of Obesity Risk is a competition in the 2024 Kaggle Playground Series, using a synthetic data set generated from the Obesity or CVD risk data set (Palechor and de la Hoz Manotas, 2023).
+The test sample size was $n=13,840$, and the number of participating teams were $m=3,498$ (excluding teams with accuracy lower than chance).
+The maximum sample estimate was $\max(\hat{\Theta})=0.9116$.
+Figure 6(a) shows the sample estimates, $\hat{\theta}_{j}$’s, zoomed in at the $3,081$ classifiers with sample estimates above $0.88$.
+The 95% CI for $\hat{\theta}=0.9116$ is $(0.9067,0.9162)$ and is shown as the red line in Figure 6(a), the lower bound enveloping $920$ of the sample estimates.
 This should raise suspicion that the right tail seen in Figure 6(a) can be the product of multiplicity.
-(a) The competition observations
-(b) A simulated realisation
-(c) Cropped competition observations
-(d) A simulated realisation
+(a) Sample estimates
+(b) A simulated realisation from a)
+(c) Cropped sample estimates
+(d) A simulated realisation from c)
 Figure 6: The competition observations and simulated realisation of classifiers with corresponding probabilities of correct classification.
 The red line is the 95% CI for $\hat{\theta}^{\,\max}$ in (a).
-To demonstrate the impact of ignoring multiplicity, let $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ be equal to the competition observations and calculate the expectation of the maximum value of a realisation.
-We use the set-up from Section 4.3 with correlation coefficient $\rho=0.6$, and random $Y_{0}$ with $\theta_{0}=\hat{\theta}^{\,max}$.
-Note that competition observations with low values are excluded in accordance with Eq.
-12, but they are not influencing the maximum.
-A realisation is shown in Figure 7(b). The upper tail, whose tip represents the maximum observed value, is stretched according to the variances of the $Y_{j}$s.
-The estimated expected maximum is $E(\hat{\tilde{\theta}}^{\,max})=0.9159$, far above the observed maximum.
-The 95% CI is calculated by $B=1,000$ bootstrap samples of size $m$ drawn from the competition observations, and then the simulation as described above is conducted for each bootstrap sample.
-The values are reported in Table 3.
-The lower bound exceeds $\hat{\theta}^{\,\max}$, and we can conclude that the maximum observed value is an unlikely candidate for state-of-the-art.
-Perhaps a more interesting problem is finding a set $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ that can result in realisations similar to the competition observations in terms of maximum observed value, and let the maximum $\tilde{\theta}_{j}$ be the state-of-the-art candidate.
-To obtain a set $\tilde{\theta}_{1},\ldots,\tilde{\theta}_{m}$ where the expected maximum is the same as the maximum of the competition observations, $E(\hat{\tilde{\theta}}^{\,\max})=\hat{\theta}^{\,\max}$, we crop the competition observations, as seen in Figure 7(c), and where $E(\hat{\tilde{\theta}}^{\,\max})$ is calculated from simulations as described above.
-Figure 7(d) shows one realisation of the cropped distribution, where the maximum observed value is close to $\hat{\theta}^{\,max}$.
-By cropping at $0.9062$ we obtain $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$, and this serves as a new candidate for the state-of-the-art.
-It can be argued that the new SOTA candidate is at risk of being a optimistic estimate, since the competition observations are from only one test set.
-The upper bound of the 95% CI can mitigate such a risk, and the competition observations are cropped to fit that criterion.
+To demonstrate the impact of ignoring multiplicity, let ${\theta}^{\prime}_{1},\ldots,{\theta}^{\prime}_{m}$ be equal to the sample estimates and calculate the expected value of the sample maximum, $E(\max(\hat{\Theta}^{\prime}))$.
+We use the set-up from Section 4.3 with correlation coefficient $\rho_{0}=0.6$, and $P(Y_{0})=\theta_{0}=\max({\Theta}^{\prime})$.
+Note that sample estimates with low values are excluded in accordance with Eq.
+9, but their influence on the expected maximum is negligible.
+A realisation is shown in Figure 6(b).
+The upper tail, whose tip represents the maximum sample estimate, is stretched according to the variances of the $Y_{j}$s.
+The estimated expected sample maximum is $E(\max(\hat{{\Theta}}^{\prime}))=0.9159$, far above the sample maximum in the competition.
+The 95% CI is reported in Table 3.
+The lower bound exceeds $\max(\hat{\Theta})$, and we can conclude that the sample maximum is an unlikely candidate for SOTA.
+The CI is calculated by $B=1,000$ bootstrap samples of size $m$ drawn from the sample estimates, and then the simulation as described above is conducted for each bootstrap sample.
+Instead, we can find a set $\Theta^{\prime}$ that results in realisations, $\hat{\Theta}^{\prime}$, similar to the competition sample estimates in terms of sample maximum, and let $\max(\Theta^{\prime})$ be the SOTA candidate.
+To obtain a set $\Theta^{\prime}$ where $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$, we crop the competition sample estimates, as seen in Figure 6(c), and where $E(\max(\hat{\Theta}^{\prime}))$ is calculated from simulations as described above.
+Figure 6(d) shows one realisation of the cropped distribution, where $\max(\hat{\Theta}^{\prime})$ is close to $\max(\hat{\Theta})$.
+By cropping at $0.9062$ we obtain $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$, and this serves as a new candidate for the SOTA.
+It can be argued that the new SOTA candidate is at risk of being a optimistic estimate, since the competition observations are from only one test sample.
+The upper bound of the 95% CI can mitigate such a risk, and the competition sample estimates are cropped to fit that criterion.
 The CI is calculated by bootstrapping, as described above.
@@
 Estimating ${\theta}_{SOTA}$
-Approach $\hat{\theta}_{SOTA}$ 95% CI $E(\max_{j}\theta_{j})$ teams
-no mult. $\hat{\theta}^{\,\max}$ (0.9067,0.9162) - -
-Kaggle accuracies $\hat{\theta}^{\,\max}$ (0.9128,0.9191) 0.9159 -
-cropping 0.9062 (0.9084,0.9148) $\hat{\theta}^{\,\max}$ 1134
-0.9026 (0.9050,$\hat{\theta}^{\,\max}$) 0.9083 1844
-Table 3: The 95% CI and the expectation and for the SOTA estimator for different estimation strategies.
+Approach $\max(\Theta^{\prime})$ 95% CI $E(\max(\hat{\Theta}^{\prime}))$ teams
+no mult. $\max(\hat{\Theta})$ (0.9067, 0.9162) - -
+$\Theta^{\prime}=\hat{\Theta}$ $\max(\hat{\Theta})$ (0.9128, 0.9191) 0.9159 -
+cropping 0.9062 (0.9084, 0.9148) $\max(\hat{\Theta})$ 1134
+0.9026 (0.9050, $\max(\hat{\Theta})$) 0.9083 1844
+Table 3: The 95% CI and the expected value for sample maximum, $\max(\hat{\Theta}^{\prime})$, for different SOTA estimation strategies, where SOTA is defined as $\max(\Theta^{\prime})$.
+The competition sample maximum is $\max(\hat{\Theta})=0.9116$.
 The bootstrap standard deviations are in the order of magnitude $10^{\text{-}5}$-$10^{\text{-}4}$, and are therefore not displayed.
+The rightmost column shows number of classifiers with sample performance above the estimated SOTA.
 Table 3 demonstrates that taking the effect of multiplicity into account changes how results are perceived dramatically.
-A more likely SOTA value is $0.9062$, and even though its reduction from best performance is only about $0.005$, the effect it has on who we consider performing at SOTA-level is dramatic: around one third of the teams have observed performance above the new SOTA.
+The new SOTA value is $0.9062$, and even though its reduction from best performance is only about $0.005$, the effect it has on who we consider performing at SOTA-level is dramatic: around one third of the classifiers have sample performance above the new SOTA.
 The more conservative approach using the upper bound of the 95% CI to estimate SOTA gives an even more dramatic result, and either way the impact is substantial.
 5.1.2 Cassava Leaf Disease
-In this example, the highest reported accuracy is unlikely to be the product of multiplicity.
+In this example, the maximum sample estimate is unlikely to be the product of multiplicity.
 For The Cassava Leaf Disease Classification competition, the task was to classify images of cassava leaves.
-The training set consisted of $21,367$ labelled images of 5 classes, and the test set of approximately $15,000$ images.
-$3,752$ teams participated (excluding those with accuracy lower than chance).
-The three top performances appear as outliers, as seen in Figure 8(a).
-The 95% CI of $\hat{\theta}^{\,\max}=0.9132$ is $(0.9086,0.9177)$, no other observations are within this interval.
-By performing simulations without cropping, we get $E(\hat{\tilde{\theta}}^{\,max})=\hat{\theta}^{\,\max}$, and although the upper tail is stretched, its tip does not surpass $\hat{\theta}^{\,max}$.
+The training set consisted of $21,367$ labelled images of five classes, and the test set of approximately $15,000$ images.
+$3,752$ teams participated (excluding those with sample performance lower than chance).
+The three top performances appear as outliers (0.9132, 0.9043 and 0.9028) as seen in Figure 7(a).
+Several explanations for the outlying performances of the three top teams have been put forward, and the interested reader can take a deep-dive in the comment section of the competition.
+Since this is not the topic of the present work, we will not engage in the discussion here.
+The 95% CI of $\hat{\theta}=0.9132$ is $(0.9086,0.9177)$, no other sample performance is within this interval.
+By performing simulations without cropping, we get $E(\max(\hat{\theta}^{\prime}))=\max(\hat{\Theta})$, and although the upper tail is stretched, its tip does not surpass $\max(\hat{\Theta})$.
 (a) The Kaggle performances
 (b) One realisation.
-Figure 8: The competition observations and one realisation.
-Several explanations for the outlying performances of the three top teams have been put forward, and the interested reader can take a deep-dive in the comment section of the competition.
-Since this is not the topic of the present work, we will not engage in the discussion here.
-5.2 Simulation of AUCs
-We discuss here an example, the SIIM-ISIC Melanoma Classification competition from 2020 because it follows many of the common structures of public challenges.
-The winning team chose many of the verified strategies from Section 2.
+Figure 7: Cassava Leaf sample performances and one realisation.
+This analysis demonstrates that the suggested approach for SOTA estimation is robust in the sense that when multiplicity is not a likely explanation to the sample maximum value, the SOTA value is not altered by this approach.
+5.2 Estimating $AUC_{SOTA}$
+5.2.1 Melanoma Classification
+We will use the Melanoma Classification competition from 2020 as an illustrative example of $AUC_{SOTA}$ estimation because it follows many of the common structures of public challenges.
+The results are presented in Section 5.2.3.
 The SIIM-ISIC Melanoma Classification competition consisted of more than $40,000$ images of skin lesions from various sites, split into training ($33,126$), validation ($3,295$) and test set ($7,687$).
@@
 The size of the test set is smaller than the recommendations of, e.g., Roelofs et al.
-(2019), but reflects the reality of benchmark datasets and public challenges, as seen, e.g., in the overview of Willemink et al.
+(2019), but reflects the reality of benchmark datasets and public challenges, as seen in, e.g., the overview of Willemink et al.
 (2020).
@@
 Participants were allowed to use additional data for training if the data was shared on the competition web site.
-The prize money was set to $10,000$ USD for the best-performing method, measured as AUC on the hold-out test set, published in the private leaderboard as seen in Figure 9.
+The prize money was set to $10,000$ USD for the best-performing method, measured as AUC on the hold-out test set.
+Figure 8 shows the results.
 Participants could get intermediate results from the validation set, published on a public leaderboard.
-Figure 9: Histogram of the AUCs from the kaggle leaderboard, test set.
-The team ‘All Data Are Ext’ consisting of three Kaggle Grandmasters won the competition, achieving an AUC of $0.9490$ on the test set, and their method and strategy is explained in Ha et al.
+Figure 8: AUCs for the Melanoma Classification competition, calculated on a hold-out test set, cropped from below for better visualisation.
+The team ‘All Data Are Ext’, consisting of three Kaggle Grandmasters, won the competition, achieving an AUC of $0.9490$ on the test set, and their method and strategy is explained in Ha et al.
 (2020).
-Although we show that the AUC is an optimistic estimate for the melanoma detection state-of-the-art, we do not question their rank in the challenge.
-Quite the contrary; the research in Section 2 supports their strategies.
-They had a low rank on the public leaderboard (880th out of 3300).
-Their AUC decreased from the public to the private leaderboard, but it was still within the confidence interval.
+Although we show that the AUC is an optimistic estimate for the melanoma detection SOTA, we do not question their rank in the challenge.
+Quite the contrary; the winning team chose many of the verified strategies from Section 2.
 They used additional data from previous years’ competitions, and they also did data augmentation.
@@
 (2019) showed will slow down the overfitting.
+In the following sections, we introduce a strategy for estimating $AUC_{SOTA}$, using simulated examples and the Melanoma Challenge for illustration.
+5.2.2 Simulation of AUC
 In a binary classification problem, the calculation of AUC is based on scores of random variables from the positive and the negative class.
-The score function plus a threshold is the classifier.
 The AUC is mathematically equivalent to the probability that a random variable from the positive class will be ranked higher, in terms of score, than a random variable from the positive class (Hanley and McNeil, 1982).
 Let $score(\cdot)$ be a function that assigns a score to a random variable.
-Let $S_{+}=score(W_{+})$, where $W_{+}$ comes from the positive class, and $S_{-}=score(W_{-})$, where $W_{-}$ comes from the negative class.
+Let $S_{+}=score(W_{+})$, where $W_{+}$ is a random variable from the positive class, and $S_{-}=score(W_{-})$, where $W_{-}$ is a random variable from the negative class.
 Then $\text{AUC}=P(S_{-}<S_{+})=P(S_{-}-S_{+}<0)$, which is equivalent to the cdf at value 00.
-Hence to simulate classifier scores according to a given true or expected AUC, $a$, we simply need to specify distributions for the variables $S_{-},S_{+}$ such a way that $P(S_{Y}-S_{X}<0)=a$.
-This is particularly simple with Normal distributions, since linear combinations of Normal distributions yields another Normal. Let $S_{+}$ and $S_{-}$ have the distributions
+To simulate classifier scores according to a given AUC $=a$, we need to specify distributions for the variables $S_{-},S_{+}$ in such a way that $P(S_{-}-S_{+}<0)=a$.
+We use Normal distributions for illustration.
+Let $S_{+}$ and $S_{-}$ have the distributions
 $S_{+}$ $\sim\mathcal{N}(\mu_{+},\sigma^{2}_{+})$ (10)
 $S_{-}$ $\sim\mathcal{N}(\mu_{-},\sigma^{2}_{-}),$ (11)
-where $\mathcal{N}(\mu,\sigma)$ is the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. We then have that
+where $\mathcal{N}(\mu,\sigma^{2})$ is the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. We then have that
 $\text{AUC}=\Phi\left(\frac{0-(\mu_{-}-\mu_{+})}{\sqrt{\sigma_{-}^{2}+\sigma_{+}^{2}}}\right),$ (12)
-by linear combination of two Normal distributions, where $\Phi$ is the cdf of the standard Normal distribution.
+where $\Phi$ is the cdf of the standard Normal distribution.
 Specifying any three of the four parameters above decides the last one for a chosen AUC $=a$.
@@
 $\sigma^{2}$ $=\pi(\sigma_{+}^{2}+\mu_{+}^{2})+(1-\pi)(\sigma_{-}^{2}+\mu_{-}^{2})-\mu^{2}.$ (15)
-The uncertainty of AUC can be calculated through bootstrapping, but without access to the predictions for each data point in the test set, the analyst must make some assumptions about the shape of the ROC curve.
-Figure 10(a) shows the pdfs of negative and positive scores with an AUC of 0.90.
-The rug shows an example draw with 3000 images, with the probability of drawing a positive image is 0.017, corresponding to the class balance in the Melanoma challenge.
-Fig.
-10 shows different shapes of ROC curves.
-(a) Score pdfs for a classifier with AUC= $0.90$; grey for negatives, black for positives. The rug shows an example draw from these.
-(b) ROC curves for a classifier with expected AUC of $0.90$ across 100 realisations of (a).
+Figure 9(a) shows an example of Normally distributed negative and positive scores and an AUC of 0.90.
+The rug shows an example draw with $n=3000$ and $\pi=0.017$, corresponding to the class balance in the Melanoma challenge.
+(a) Score pdfs for a classifier with AUC= $0.90$; grey for negatives, black for positives. The rug shows an example draw from these with class balance $\pi=0.017$.
+(b) ROC curves for a classifier with AUC $=0.90$ across 100 realisations of (a).
 Figure 9: Example of what an AUC of $0.90$ might look like in practice.
-The right figure shows $100$ realizations of a ROC curve from a competition with $3000$.
-Out of the $3000$ images, $52$ are positives, reflecting the proportion of the Kaggle melanoma challenge.
-The left figure shows the underlying score distributions that generated the curves of Figure 10(b).
+The right figure shows $100$ realizations of a ROC curves.
+The left figure shows the underlying score distributions that generated the curves in (b).
+The choice of distributions has an impact on the final SOTA estimation, as it affects the variance of the ROC curves.
 However, with high AUC, the shape has less impact, and the analyst’s choices are not suggestive of the overall conclusion.
 5.2.3 Simulation results for uncorrelated classifiers
-For simplicity we choose $\mu_{-}=0$ and $\sigma_{-}=\sigma_{+}=1$. We then solve for $\mu_{+}$ numerically. Now we can generate scores from the two resulting normal distributions.
 To demonstrate the impact that multiplicity can have in a public competition, we start with a simulated example similar to the one in Section 3.4.
 Consider a classification problem with a test set of size $n=3,000$, number of participating teams, $m=1,000$, where each classifier has a true underlying AUC of $a=0.90$, and the classifiers are independent.
-Whereas the distribution of $\hat{\theta}^{\,\max}$ is independent of the class balance as long as the probability of correct prediction is independent of class, this is not true for maximum of $\hat{AUC}$, where the variance increases with class imbalance.
-We have chosen the same class balance as in the Melanoma challenge, as this reflects real world examples.
-The 95% CI of a single classifier with observed AUC equal to its true AUC of $0.90$ is (.$(.8558,0.9376)$, and gives a strong indication of what we can expect when $1,000$ teams apply their classifiers on a test set.
-Figure 11: Example distribution of simulated observed state of the art with parameters as in Figure 10. This is an analogous figure to Figure 3 above.
-Figure 11 shows the simulated distribution of the observed maximum AUC over $10,000$ repetitions of a competition where every classifier has a true AUC of $0.90$.
+Whereas the distribution of the sample maximum for probability of success is independent of class balance as long as the probability of success is independent of class, this is not true for the sample maximum of AUC, where the variance increases with class imbalance.
+We have chosen the same class balance as in the Melanoma challenge, $\pi=0.017$, as this reflects real world examples.
+For simplicity we choose $\mu_{-}=0$ and $\sigma_{-}=\sigma_{+}=1$.
+We then solve for $\mu_{+}$ numerically.
+Now we can generate scores from the resulting Normal distribution.
+The 95% CI of a single classifier with $\hat{AUC}=0.90$ is (.$(.8558,0.9376)$, and gives a strong indication of what we can expect when $1,000$ teams apply their classifiers on a test set.
+Figure 10: Simulated distribution of the sample maximum AUC with parameters as in Figure 9. This is an analogous figure to Figure 2(b) in Section 3.4.
+Figure 10 shows the simulated distribution of the sample maximum AUC over $10,000$ repetitions of a competition where every classifier has a true AUC of $0.90$.
 The simulated expected value is $0.9562$, the standard deviation is $0.004459$, and the simulated 95% CI is $(0.9486,0.9662)$.
-To give an unbiased estimate of the SOTA performance for the Melanoma challenge, we follow the same procedure as in Section 5.1: The competition observations are cropped in such a manner that their realisations give an expected value of $0.9490$, the observed maximum.
-(a) One bootstrap sample from the cropped distribution.
-(b) One realisation of the bootstrap sample in (a).
-Figure 12: Histogram of cropped competition observations. Here the state of the art is an AUC of 0.8975 which leads to the expected observed maximum AUC in a competition being 0.9490.
+Comparing this to the results in Section 3.4, we see that the bias is much larger for the sample maximum AUC.
+This is partly due to the large variance created by the class imbalance.
+Although the sample size is $n=3,000$, only $51$ of those belong to the positive class, creating large variance in the estimator.
+Again, it is clear that the observed maximum is an unlikely SOTA candidate, and to give a better estimate of the SOTA for the Melanoma challenge, we follow the same procedure as in Section 5.1: The competition sample estimates are cropped in such a manner that their realisations give an expected value of $0.9490$, the competition sample maximum.
+(a) Truncated AUC distribution based on the Melanoma challenge scores. The part of the distribution removed by truncation is shaded in dark grey.
+(b) A realization from the truncated distribution of (a).
+Figure 11: Histogram of cropped competition AUC sample estimates. Here the state of the art is an AUC of 0.8975 which leads to the expected sample maximum AUC of 0.9490.
 Figure 11(a) shows one bootstrap sample from the kaggle observations cropped at $0.8975$.
 Number of classifiers is set to $m=3,173$, which is as many as there are classifiers in the melanoma challenge with AUC $>0.5$.
-Figure 12(b) shows a corresponding realisation, where the expected maximum is $0.9490$.
-There are $2022$ teams with performance above the new SOTA candidate.
+Figure 11(b) shows a corresponding realisation, where the maximum is close to $0.9490$.
+There are $2,022$ teams with performance above the new SOTA candidate.
+The discrepancy between the sample maximum and the new SOTA candidate is big, and have big implications.
+From an AUC close to $95\%$ to just under $90\%$ is a substantial drop in performance in a cancer detection setting.
+In addition, two-thirds of the classifiers have performances above the new SOTA estimate.
+As was shown in Section 4, illustrated in Figure 4(b), the bias is reduced for correlated classifiers, so the analysis with independence assumption is an overestimation of the bias.
+In this challenge, we can assume multiple sources of dependency: Access to the same training set, models pre-trained on the same data, similar models, etc.
+However, the wide CI for a single classifier tells the story of an estimator with large variance, and hence, the bias of the sample maximum will be substantial.
 5.2.4 Correlation between classifiers in the AUC simulations
@@
 $\begin{pmatrix}S_{0^{+}}\\ S_{j^{+}}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mu_{0^{+}}\\ \mu_{j^{+}}\end{pmatrix},\begin{pmatrix}\sigma_{0^{+}}^{2}&\rho\sigma_{0^{+}}\sigma_{j^{+}}\\ \rho\sigma_{0^{+}}\sigma_{j^{+}}&\sigma_{j^{+}}^{2}\end{pmatrix}\right),$
-and equivalently for the negative class. However, the joint distribution of $(S_{0},S_{j})$ is not uniquely defined by the correlation, and simulations are not straightforward.
+and equivalently for the negative class.
+However, the joint distribution of $(S_{0},S_{j})$ is not uniquely defined by the correlation, and additional assumptions must therefore be made.
+We leave this as future work.
 6 Discussion
 We have addressed the mechanisms of multiplicity and how they affect the bias of the SOTA estimator.
-When accuracy is used as a performance measure, the exact probability distribution can be provided, and known analyses methods can be engaged.
-The bias in SOTA is substantial in our example, and possibly hinders new methods to attract interest from the scientific community.
-We have also demonstrated that dependency might not have a big impact on the bias when the probability of correct prediction is already high.
-The exception is when the dependency is very strong.
+When probability of success is used as a performance measure, the exact distribution of the sample maximum estimator can be provided in the simpler cases, and known analyses methods can be engaged.
+The bias in SOTA is substantial in our examples, and possibly hinders new methods to attract interest from the scientific community.
 We have presented the simple case with an analytical solution, where the probabilities of correct prediction are known.
 We have also presented three case-studies from Kaggle competitions, where we demonstrate how a different SOTA candidate than the observed maximum can be estimated, and how it impacts the perceived results.
-This work shows that a multiplicity-adjusted state-of-the-art estimate can be calculated fairly easily through bootstrapping, and can serve as a better reference point when introducing new methods to the scientific community.
+This work shows that for some cases, an unbiased SOTA estimate can be calculated fairly easily through bootstrapping, and can serve as a better reference point when introducing new methods to the scientific community.
+In other cases, the SOTA estimation is more complicated.

### v6

@@
 Accounting for multiplicity in machine learning benchmark performance
-Accounting for multiplicity in machine learning benchmark performance
-Kajsa Møllersen1 and Einar Holsbø2
-(1Department of Community Medicine
-2Department of Computer Science
-UiT - The Arctic University of Norway
+Abstract
+State-of-the-art (SOTA) performance refers to the highest performance achieved by some model on a test sample, preferably under controlled conditions such as public data (reproducibility) or public challenges (independent sample).
+Thousands of classifiers are applied, and the highest performance becomes the new reference point for a particular problem.
+In effect, this set-up is an estimate of the expected best performance among all classifiers applied to a random sample; a sample maximum estimate.
+In this paper, we argue that SOTA should instead be estimated by the expected performance of the best classifier, which can be done without knowing which classifier it is.
+Our contribution is the formal distinction between the two, and an investigation into the practical consequences of using the former to estimate the latter.
+This is done by presenting sample maximum estimator distributions for non-identical and dependent classifiers.
+We illustrate the impact on real world examples from public challenges.
 1 Estimating state-of-the-art performance
-In this article, we direct attention to the reporting of state-of-the-art classifier performance by presenting the sample maximum estimator distribution function and analysing its bias.
-This analysis is done using the framework of multiplicity.
-Within the fields of research that develop, refine, test and use classifiers, the term state-of-the-art (SOTA) commonly refers to the highest reported performance to date, and is used as a reference point when presenting and comparing models.
-The reported performance of a classifier, originating from applying that classifier to a specific data set, a sample, is in effect an estimate of that classifier’s underlying population performance.
-The highest reported performance among a number of classifiers serves as an estimate of SOTA, although the term ’estimate’ is seldom used explicitly in the SOTA context.
-The highest reported performance corresponds to the sample maximum of population performances, and this estimator is biased when used to estimate SOTA.
-The sample maximum gives in overoptimistic estimates of the maximum population performance, and is problematic as a reference point in research.
-The mechanisms creating this bias comes from multiplicity, and the fact that the we pick the highest sample performance among multiple classifier results.
-In its simplest form, with independent binary classifiers with identical probability of success, the distribution function of the sample maximum estimator exists in closed form with only a few parameters, and can be directly applied to calculate and analyse the bias in a certain situation.
+Machine learning (ML) models are commonly evaluated and compared by their performance on data sets from public repositories, where models can be evaluated under identical conditions and across time.
+Theoretical results play an important role in ML research, but for the last couple of decades the success of increasingly complex models has pushed the evidence towards performance on benchmark data sets.
+In this paper, we direct attention to the reporting of state-of-the-art (SOTA) performance.
+We restrict our attention to classifier performance, but the overall mechanisms apply to other models.
+The term performance will in general refer to quantification of the classifier’s ability to correctly classify the instances in a sample, and we will specify when necessary.
+Other performance measures, such as computation time, will not be addressed here.
+Because of the complexity of both models and their training and test samples, the performances are estimated by applying the models to data samples.
+Within the fields of research that develop, refine, test and use classifiers, the term SOTA commonly refers to the highest reported performance to date, and is used as a reference point when presenting and comparing models.
+The highest performance serves as an estimate of SOTA, although the term “estimate” is seldom used explicitly.
+There is an important distinction between best performance among a number of classifiers, and performance of best classifier, and this paper makes an effort to bring attention to this distinction.
+1.1 SOTA performance
+SOTA refers to the best model, and hence SOTA performance is the performance of that model.
+Let $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$ be the probability of success for $m$ classifiers, and let $\hat{\theta}_{j}$, $j=1,\ldots,m$, be their corresponding unbiased estimates.
+In a public challenge, $\Theta$ would be the classifier performances of the $m$ participants.
+Taking into account that the participants of a challenge is somewhat arbitrary, and that their classifiers are adjusted to a training sample, we model the classifier ensemble as random sample where $\theta_{j}\sim P(\theta)$, and define SOTA performance as follows
+Definition 1 (SOTA classifier performance) Let $\theta_{j}$ be the random variable that denotes the performance of classifier $j$.
+In an ensemble of $m$ classifiers with performances $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$, the SOTA classifier performance is the expected best performance
+$\theta_{SOTA}=\text{E}\max\Theta.$ (1)
+For non-random classifier ensembles, the definition holds since the expectation of a constant is just the constant itself, $\theta_{SOTA}=\max\Theta$.
+This definition answers the question what can we expect the best classifier’s performance to be?, when applied to a new data sample.
+Note that picking the highest estimated performance derived from an ensemble of classifiers applied to a test sample does not answer this question.
+1.2 Sample maximum performance
+ML models are commonly ranked by their performance on benchmark datasets, published as leaderboards, often in public challenges where the test sample is withheld to avoid overfitting.
+The best performance to date, or among the participants in a challenge, is referred to as SOTA, see, for example paperswithcode.com.
+The published performances are estimates, $\hat{\theta}_{j}$’s, based on the application of models on a specific benchmark or test sample.
+Although seldom formalised, these are often unbiased estimates, and we will treat them as such.
+The test samples act as random samples, all the while the point of reporting these performances is the assumption that they can be generalised to similar data sets.
+Hence, the $\hat{\theta}_{j}$’s are a random sample, and the best performance is the sample maximum estimate, which we define for clarity.
+Definition 2 (Sample maximum estimator) Let $\theta_{j}$ be the random variable that denotes the performance of classifier $j$ and let $\hat{\theta}_{j}(X)$ be the corresponding unbiased estimator.
+For an ensemble of $m$ classifiers with performances $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$, the sample maximum performance estimator is
+$\operatorname{\hat{\theta}_{\max}(\mathbf{X})}=\text{E}\max(\hat{\Theta}(\mathbf{X})),$ (2)
+where $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ quantifies successes for classifiers $j=1,\ldots,m$.
+The sample maximum answers the question What is the best expected performance of an ensemble of classifiers?
+When $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ is used as an estimator for $\theta_{SOTA}$, it is biased.
+The two terms look similar at a glance, the only thing distinguishing them are $\hat{\Theta}(\mathbf{X})$ vs $\Theta$.
+Taking into account that $\hat{\theta}(X)$ is an unbiased estimator for $\theta$, the distinction can easily be overlooked.
+The difference reflects a multiplicity effect.
+1.3 Multiplicity effect
+Multiplicity is, simply put, how the probability of an outcome changes when an experiment is performed multiple times.
+As we will show in the subsequent sections, using the maximum performance of an ensemble as an estimate for SOTA is a biased estimator giving overoptimistic results.
+The bias is a multiplicity effect; multiple models being applied to the same sample.
+The optimistic estimate is often used as a reference point for evaluating new models.
+Not to say that a new model necessarily needs to beat the current SOTA to catch interest from reviewers and editors, but it certainly helps if its performance is comparable.
+Many aspects of public benchmark datasets and public challenges have been thoroughly discussed, and we present some of the topics connected to multiplicity in Section 2.
+As far as the authors are aware of, the multiplicity effect has been nearly absent from the discussion regarding SOTA estimates.
+We present the sample maximum estimator distribution function and its bias for some classifier scenarios.
+In its simplest form, independent classifiers with identical probability of success, the distribution function of the sample maximum estimator exists in closed form.
 With growing complexity, such as mutual dependencies and more elaborate performance measures, the distribution function of the estimator is no longer known, but the mechanisms at play and the pitfalls to be aware of are much the same, and limited studies can be done by straightforward simulations.
-Machine learning models are commonly evaluated and compared by their performance on data sets from public repositories, where models can be evaluated under identical conditions and across time.
-Theoretical results play an important role in machine learning research, but for the last couple of decades the success of increasingly complex models has pushed the proof towards the pudding: performance on benchmark data sets.
-The numbers don’t lie: if model A performs better than model B on a specific data set, that is what it did.
-But the truth that these numbers tell is limited.
-We are interested in the population performance, and the sample performance is only an estimate.
-A good estimating procedure gives a reliable result, and there are plenty of pitfalls to be aware of: Overfitting to the data set, originating from wrong use of cross-validation, data leakage and data reuse, can give overly optimistic estimates.
-An independent test set, often referred to as a hold-out set, is a sample that is not used for development of the model, and prevents overfitting.
-Competitions and challenges with prize money typically withhold the test set so that it becomes impossible for the participants to use it for training.
-The highest ranked performance on a problem is referred to as SOTA performance, and is used, among other things, as a reference point for publication of new models.
-As a specific model’s performance on a test set is just an estimate of its true underlying performance, the reported SOTA performances are estimates, and hence come with bias and variance.
-The practice of hold-out test sets in competitions safe-guards against overoptimistic estimates for the individual classifiers.
-Data is often released after the competition has ended, and new methods can be compared to the SOTA performance.
-As we will show in the subsequent sections, using the highest-ranked performance as an estimate for SOTA is a biased estimator, giving overoptimistic results.
-The mechanisms at play are those of multiplicity, a topic that is well-studied in the context of multiple comparisons and multiple testing, but has, as far as the authors are aware of, been nearly absent from the discussion regarding SOTA estimates.
-Biased SOTA estimation has consequences both for practical implementations and for research itself.
-A health care provider might want to use some machine learning method for tumour detection, but only if the method correctly classifies at least 80%.
-They follow updates on SOTA performance, and are ready to invest time and money in a pilot project once the lower bound surpasses 80%.
-If the SOTA performance is an optimistic estimate, the health care provider has now lost time and money, and perhaps also some faith in research results.
-For the research community, the situation is perhaps even more dramatic.
-The optimistic SOTA estimate is used as a standard for evaluating new methods.
-Not to say that a new method necessarily needs to beat the current SOTA to catch interest from reviewers and editors, but it certainly helps if its performance is comparable.
-It has been pointed out (see Section 2) that the incremental increase in performance can obstruct the spread of new ideas that have substantial potential, but just haven’t reached their peak yet.
-In this article, we provide a probability distribution for the case of multiple classifiers.
-We demonstrate the impact of multiplicity through simulated examples.
-We show how classifier dependency impacts the variance, but also that the impact is limited when the probability of success is high.
-Finally, we discuss three real-world examples that each highlights different aspects multiplicity and SOTA estimation.
+We demonstrate the impact of multiplicity through simulated examples, and show how classifier dependency impacts bias and variance.
+Finally, we discuss three real-world examples that each highlights different aspects of multiplicity and SOTA estimation.
+From these real-world examples, we suggest a SOTA estimate, in line with Definition 1.
+This is a new perspective for multiplicity effects in the ML context, a step aside from overfitting and re-use of test sets.
 All code is available at https://github.com/3inar/ninety-nine.
 2 Related work
-Multiple comparisons is a well-studied subject within statistics, see, e.g., Tukey (1991) for an interesting discussion on some topics of multiple comparisons, among them the confidence intervals (CIs).
+Multiple comparisons is a well-studied subject within statistics, see, for example, Tukey (1991) for an interesting discussion on some topics, among them confidence intervals (CIs).
 For a direct demonstration of how multiple testing can lead to false discoveries, Bennett et al.
-(2009) give an example where brain activity is detected in (dead) salmon “looking” at photos of people.
-In the machine learning context, with multiple data sets and classifiers, Demšar (2006) suggests adjusting for multiple testing, according to well-known statistical theory.
-Our article focuses on the estimate itself and its uncertainty, and not comparison of methods.
-The re-use of benchmark datasets has been addressed in several publications.
-Salzberg (1997) point to the problem that arises because these datasets are static, and therefore run the risk of being exhausted by overfitting.
-Exhausted public data sets still serve important functions, e.g., a new idea can show to perform satisfactory, even when it is not superior.
+(2009) gave an example where brain activity is detected in (dead) salmon “looking” at photos of people.
+In the ML context, with multiple data sets and classifiers, Demšar (2006) suggested adjusting for multiple testing, according to well-known statistical theory.
+This paper focuses on the estimate itself and its uncertainty, and not comparison of methods.
+The use and re-use of benchmark datasets has been addressed in several publications.
+Longjohn et al.
+(2025) drew attention to the role of data repositories with leaderboards and their lack of metric uncertainty.
+Salzberg (1997) pointed to the problem that arises because these datasets are static, and therefore run the risk of being exhausted by overfitting.
+Exhausted public data sets still serve important functions, for example, a new idea can show to perform satisfactory, even when it is not superior.
 Thompson et al.
-(2020) distinguish between simultaneous and sequential multiple correction procedures.
+(2020) distinguished between simultaneous and sequential multiple correction procedures.
 In sequential testing, which is the reality of public datasets, the number of tests is constantly updated.
@@
 Teney et al.
-(2020) performed a successful attack on the VQA-CP benchmark dataset, that is, they are able to attain high performance with a non-generalisable classifier.
+(2020) performed a successful attack on the VQA-CP benchmark dataset (introduced by Agrawal et al., 2018); they were able to attain high performance with a non-generalisable classifier.
 They point towards the risk of ML methods capturing idiosyncrasies of a dataset.
@@
 motivating their concern of methods being tailored to a dataset rather than a problem.
-The lack of statistically significant differences among top performing methods is demonstrated by, e.g., Everingham et al.
+The lack of statistically significant differences among top performing methods is demonstrated by, for example, Everingham et al.
 (2015) and Fernández-Delgado et al.
 (2014).
-The same mechanisms of overfitting are present in challenges through multiple submissions to the validation set.
+The same mechanisms of overfitting are present in public challenges through multiple submissions on the validation sample.
 Blum and Hardt (2015) introduced an algorithm that releases the public leaderboard score only if there has been significant improvement compared to the previous submission.
@@
 (2019) constructed new, independent test sets of CIFAR-10 and ImageNet.
-They found that the accuracy drop is dramatic, around 10%.
-However, the order of the best performing methods is preserved, which can be explained by the fact that multiple classes (Feldman et al.
-(2019)) and model similarity (Mania et al.
-(2019)) slow down overfitting.
+They found a dramatic drop in accuracy, around 10%.
+However, the order of the best performing methods is preserved, which can be explained by the fact that multiple classes (Feldman et al., 2019) and model similarity (Mania et al., 2019) slow down overfitting.
 Roelofs et al.
 (2019) compared the results from the private leaderboard (one submission per team) and the public leaderboard (several submissions per team) in Kaggle challenges.
-Several of their analyses indicate overfitting even though their main conclusion is that there is little evidence of overfitting.
-Overfitting and its consequences are avoided in most challenges by withholding the test set so that a method can only be evaluated once and not adapted to the test set performance.
-Hold-out datasets also prevents wrong use of cross-validation, see, e.g., Fernández-Delgado et al.
+Several of their analyses indicate overfitting even though their main conclusion was that there is little evidence of overfitting.
+Overfitting and its consequences are avoided in most challenges by withholding the test sample so that a method can only be evaluated once and not adapted to the result.
+Hold-out datasets also prevents wrong use of cross-validation, see, for example, Fernández-Delgado et al.
 (2014) and the seminal paper of Dietterich (1998).
-Another contribution from public challenges is the requirement of reproducibility to collect the prize money, which is a well-known concern in science, see, e.g., Gundersen and Kjensmo (2018).
+Another contribution from public challenges is the requirement of reproducibility to collect the prize money, which is a well-known concern in science, see, for example, Gundersen and Kjensmo (2018).
 Public challenges typically rank the participants by a (single) performance measure, and can fall victims of Goodhart’s law, paraphrased as “When a measure becomes a target, it ceases to be a good measure.” by Teney et al.
 (2020).
+Møllersen et al.
+(2017) demonstrated how five different performance measures completely rearranges the leaderboard for the top participants in the ISBI 2016 challenge Skin Lesion Analysis towards Melanoma Detection (Gutman et al., 2016).
 Ma et al.
 (2021) offered a framework and platform for evaluation of NLP methods, where other aspects such as memory use and robustness are evaluated.
-The robustness/stability of a method, and the variability in performance are crucial aspects that the hold-out test set of public challenges does not offer an immediate solution to.
+The robustness of a method, and the variability in performance are crucial aspects that the hold-out test set of public challenges do not offer an immediate solution to.
 Bouthillier et al.
@@
 Bousquet and Elisseeff (2002) investigated how sampling randomness influences accuracy, both for regression algorithms and classification algorithms, relying on the work of Talagrand (1996) who stated that “A random variable that depends (in a “smooth” way) on the influence of many independent variables (but not too much of any of them) is essentially constant.”
-There is a rich literature on various aspects regarding the use of public challenges as standards for scientific publications, see, e.g., Varoquaux and Cheplygina (2022) for a much broader overview than the one given here.
-Concerns and solutions regarding multiplicity and better performance estimates are active research fields.
+There is a rich literature on various aspects regarding the use of public challenges as standards for scientific publications, see, for example, Varoquaux and Cheplygina (2022) for a much broader overview than the one given here.
 We contribute to this conversation by shining the light on how multiplicity of classifiers’ performance on hold-out test sets create optimistic estimates of SOTA performance.
 3 Multiple classifiers and biased state-of-the-art estimation
-3.1 Notation
+In the following sections, we take a closer look at the sample maximum estimator by presenting its distribution and investigating its bias and variance when used for SOTA estimation.
+We wrap the presentation in a public challenge context to refer back to previous work presented in Section 2.
+We use standard deviation instead of variance, because standard deviation is below $0.005$ for our numerical examples.
+We use the following default parameter values: $n=3,000$, $\theta=0.90$, and $m=1,000$, $\theta_{SOTA}=0.9$.
+3.1 Notation and parameter values
 […]
@@
 […]
-$\,\,\,\,X\sim\mathcal{B}(n,1\text{-}\theta)$
-$\hat{\theta}(X)=\frac{n-X}{n}$ $:\text{estimator for }\theta$
-$\hat{\theta}=\frac{n-x}{n}$ $:\text{estimate of }\theta$
-$P(X=x)$ $:\text{probability of $x$ failures in one experiment}$
-$P_{x}=P(X\leq x)$ $:\text{probability of at most $x$ failures in one experiment}$
-$\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ $:\text{number of failures for classifiers }j=1,\ldots,m$
-$C_{x}$ $:\text{number of classifiers with at most $x$ failures}$
+$\,\,\,\,X\sim\mathcal{B}(n,\theta)$
+$\hat{\theta}(X)=\frac{X}{n}$ $:\text{estimator for }\theta$
+$P(X=x)$ $:\text{probability of $x$ successes}$
+$P_{x}=P(X\leq x)$ $:\text{probability of at most $x$ successes}$
+$\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ $:\text{number of successes for classifiers }j=1,\ldots,m$
+$C_{x}$ $:\text{number of classifiers with at most $x$ successes}$
 $\,\,\,\,C_{x}\sim\mathcal{B}(m,P_{x})$
-[…]
-$\,\,\,\,z=0,1,2,\ldots,n$
+$\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$ $:\text{probability of success for classifiers }j=1,\ldots,m$
+$\,\,\,\,\theta\sim P(\theta)$
+$\Theta^{\prime}=\{\theta^{\prime}_{1},\theta^{\prime}_{2},\ldots,\theta^{\prime}_{m}\}$ $:\text{empirical Bayes estimate for }P(\theta)$
+$\hat{\Theta}(\mathbf{X})=\{\hat{\theta}_{1}(X_{1}),\hat{\theta}_{2}(X_{2}),\ldots,\hat{\theta}_{m}(X_{3})\}$ $:\text{estimator for }\Theta$
 […]
@@
 […]
+Some of the expectations, standard deviations and distributions need to be estimated by simulations.
+In the hierarchical set-up, we use $B=1,000$ draws from $P(\theta)$ in Eq.
+1b, and $rep=100,000$ when drawing random samples of $\mathbf{X}$ for each $\theta_{j}$.
+We denote the estimates without the “hat”, leaning on the Law of Large Numbers and the access to arbitrary large $B$ and $rep$.
+We refer to classifiers with identical probability of success as identical classifiers. We use significance level $\alpha=0.05$ and two-sided CIs.
+3.2 The invention of a coin-flip classifier
+We here give a motivating example, related to multiple testing, to illustrate the effect of multiplicity.
+For any classifier that predicts the outcome of the flip of a fair coin, the probability of success is $0.5$, and hence the SOTA performance is $\theta_{SOTA}=0.5$.
+If we had a public challenge where the aim was to make a coin-flip classifier, what would the SOTA performance be?
+Consider an experiment where a classifier predicts the outcome of a fair coin being flipped $n=20$ times.
+Let the random variable $X$ denote the number of successes, $X\sim\mathcal{B}(n,\theta)$, and let ${\hat{\theta}(X)=X/n}$.
+We are interested in the probability of observing a large $\hat{\theta}$, say, $\geq 0.90$, and wrongfully getting the impression that a coin-flip classifier has been invented.
+For a single classifier, this is calculated from $1-P(X\leq x|n,\theta)$, which in our example gives us a probability of $0.00020$, reasonably low for making the false claim of coin-flip classifier invention.
+For the multiplicity setting with $m$ independent classifiers, define the “success” of a classifier as $\{X\leq x\}$.
+Each classifier is then a Bernoulli trial.
+Let $C_{x}$ be the random variable that denotes the number of classifiers with success.
+We have that $C_{x}\sim\mathcal{B}(m,P_{x})$, where $P_{x}=P(X\leq x|n,\theta)$.
+Of special interest in this example is one or more classifiers performing at $0.90$ or above, defined as “failure” in this set-up.
+This can be written as $1-P(C_{x}=m)=1-P_{x}^{m}$.
+$P_{x}^{m}$ is equivalent to the probability of the sample maximum, or the $m$th order statistic, for random samples.
+With $m=1,0000$ classifiers, the probability of at least one classifier performing at $0.90$ or above is $0.18$, a substantial risk of grossly overestimating $\theta_{SOTA}$.
+The same mechanism as described here is present when highest model performance is reported from public challenge leaderboards or public data sets.
+Even without the usual suspects of data re-use, overfitting, wrong use of cross-validation, multiple testing or data leakage, multiplicity, when not properly accounted for, can give the impression of the invention of a coin-flip classifier.
+3.3 The probability distribution of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$
+Let $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$ be the probability of success for $m$ classifiers, and let $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ be the set of random variables that denotes the number of successes (correct predictions) for each classifier on a sample of size $n$.
+Let the statistic $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}=\max(\{\hat{\theta}_{1}(X_{1}),\hat{\theta}_{2}(X_{2}),\ldots,\hat{\theta}_{m}(X_{m})\})$ be the naive sample maximum estimator for $\theta_{SOTA}$.
+We are interested in the probability distribution of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$.
+In the coin-flip example, the $\theta_{j}$’s were equal and fixed, but we can generalize to $\theta_{j}\sim P(\theta)$, where the $m$ classifiers from where $\mathbf{X}$ originates are themselves a random sample.
+We can model the probability mass function (pmf) as a hierarchical distribution as follows:
+$\operatorname{\hat{\theta}_{\max}(\mathbf{X})}|\Theta$ $\sim P(\operatorname{\hat{\theta}_{\max}(\mathbf{X})}|\Theta)$ (1a)
+$\theta_{j}$ $\sim P(\theta).$ (1b)
+The sampling procedure is illustrated in Figure 1.
+Sampling procedure for the iid model
 […]
+Figure 1: The $\theta_{j}$’s are sampled from $P(\theta)$. The $x_{j}$’s are then sampled from their respective distributions, and $\hat{\theta}_{\max}$ follows directly.
+The sample maximum is equivalent to at least one classifier having more than $x$ correct predictions, and can be written as $1-P(\operatorname{\hat{\theta}_{\max}(\mathbf{X})}\leq x/n)$, which in turn corresponds to $1-F(x|\Theta)$.
+In the case where $\Theta$ is fixed the hierarchical distribution reduces to Eq.
+1a, and we denote it simply by $F(x)$.
+We will continue this section with fixed $\Theta$, and introduce random $\Theta$ in Section 4.
+We first present identical and independent classifiers, for which the sample maximum estimator is just a special case of the order statistic.
+We have the following cumulative distribution function (cdf):
+$F(x)$ $=P_{x}^{m}=P(X\leq x)^{m}=\big{(}\sum_{\ell=0}^{x}P(X=\ell)\big{)}^{m},$ (2)
+and the corresponding pmf:
+$f(x)$ $=P_{x}^{m}-P_{x-1}^{m}$
+$=\big{(}\sum_{\ell=0}^{x}P(X=\ell)\big{)}^{m}-\big{(}\sum_{\ell=0}^{x-1}P(X=\ell)\big{)}^{m}$ (3)
+$=\sum_{k=0}^{m-1}\binom{m}{k}\big{(}\sum_{\ell=0}^{x-1}P(X=\ell)\big{)}^{k}P(X=\ell)^{m-k}$
+where $f(x=0)=F(x=0)$. See Casella and Berger (2002, p. 228) for details. Figure 2 shows an example with default parameter values.
+Cdf and pdf for identical, independent classifiers
+(a) Cdf $F(x)$, see Eq. 2
+(b) Pmf $f(x)$, see Eq. 3.3
+Figure 2: The cdf and pmf of the sample maximum estimator for $n=3,000$, $\theta=0.90$, and $m=1,000$.
+The horizontal axes denotes the number of successes, $x$, at the bottom and the corresponding sample maximum estimates, $\operatorname{\hat{\theta}_{\max}}$, at the top.
+3.4 A simulated public challenge
+Consider a public challenge consisting of a classification problem on a test sample of size $n=3,000$, and $m=1,000$ participants.
+Let the classifiers be independent with identical probability of success, $\theta=0.90$, and hence $\theta_{SOTA}=0.90$.
+The classifier with the highest performance, $\hat{\theta}$, is declared winner, commonly referred to as SOTA.
+Figure 3 illustrates the slim chances a single better classifier has of beating an ensemble of worse classifiers.
+The pmfs of two single $\hat{\theta}(X)$ compared to the $\operatorname{\hat{\theta}_{\max}}$ statistics
 […]
+In Figure 3, the blue curve shows the pmf of a single $\hat{\theta}(X)$.
+The CI of $\hat{\theta}=E\hat{\theta}=\theta$, denoted by $(\theta_{\alpha/2},\theta_{1\text{-}\alpha/2})$, is the blue line on the horizontal axis.
+If we take multiplicity into account, the expected performance, in terms of sample maximum for the $m=1,000$ participating teams, is $E\operatorname{\hat{\theta}_{\max}}=0.9173$, indicated on the horizontal axis together with the upper limit of its CI, denoted as $\theta^{m}_{1\text{-}\alpha/2}$.
+Consider a new classifier with probability of success $\theta_{m+1}=\theta_{1\text{-}\alpha/2}$.
+The corresponding pmf is displayed as the red curve in Figure 3, with the dotted line indicating $\text{E}\hat{\theta}_{m+1}=\theta_{1\text{-}\alpha/2}$.
+By common scientific standards, a classifier performance of $\hat{\theta}=\theta_{m+1}$ is considered significantly better than a performance of $\hat{\theta}=\theta$, but the multiplicity effect can conceal this.
+The probability of the new and better classifier beating the ensemble of worse classifiers is low: $P(\hat{\theta}_{m+1}\geq\text{E}\operatorname{\hat{\theta}_{\max}})=0.10$, displayed as the green areas under the red curve.
+This means that by using the sample maximum as an estimator for SOTA, we can expect that 9 out of 10 classifiers that are significantly better will still be naively considered worse.
+The distribution of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ for these default parameters is displayed in Figure 2.
+To get an insight in how the parameter values impacts the distribution, Figure 4 shows the bias and standard deviation of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ as a function of $m$, $n$ and $\theta$, respectively.
+The bias and standard deviation of the sample maximum estimator
+Figure 4: Bias and standard deviation of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ as a function of each of its three parameters.
+For the black lines, two parameters are set to the default values $n=3,000$, $\theta=0.9$ and $m=1,000$; the third is on the x-axis.
+The other lines represent offsets, with values displayed in the legends.
+Intersections with the the gray vertical lines correspond to entries in Table 3 in Appendix A.
+In Figure 4(a), we see that the impact of number of classifiers is dramatic, and multiplicity considerations must be treated with full attention starting at $m=2$.
+Figure 4(c) demonstrates that the test set size also has large impact on the bias, which can be traced back to the standard error, $\sqrt{\theta(1-\theta)/n}$.
+The standard error also explains the decrease in bias as $\theta$ approaches 1, as seen in Figure 4(e): The standard deviation of the individual $\hat{\theta}_{j}(X)$’s is at its highest for $\theta=0.5$ and decreases monotonically.
+Figures 4(b), 4(d) and 4(f) show the corresponding standard deviations. For numerical values, see Table 3 in Appendix A.
+4 Non-identical $\theta$’s and classifier dependency
+The probability distribution in Section 3.4 is that of independent classifiers with identical $\theta$’s, an assumption that is rarely true in practice.
+We expand our investigation to non-identical $\theta$’s and classifier dependency, non-iid for short.
+We model this as the hierarchical distribution presented in Section 3.3:
+$\operatorname{\hat{\theta}_{\max}(\mathbf{X})}|\Theta$ $\sim P(\operatorname{\hat{\theta}_{\max}(\mathbf{X})}|\Theta,m,n,\rho)$ (1a)
+$\theta_{j}$ $\sim P(\theta),$ (1b)
+where $m,n,\rho$ are fixed, and $\rho$ is the correlation coefficient, see Section 4.2.
+In this model, the classifiers are a random sample from the distribution $P(\theta)$.
+$\theta_{SOTA}$ is no longer a fixed number, but defined as the expected maximum, $\theta_{SOTA}=E\max(\Theta)$.
+For the demonstrations in the following subsections we will use $\theta\sim\mathcal{U}(a,b)$, the continuous uniform distribution, for which the sample maximum statistic, $T$, has probability density function (pdf)
+$f(t)=\frac{m(t-a)^{m-1}}{(b-a)^{m}},$
+and expected value
+$\theta_{SOTA}=E\max(\Theta)=ET=\frac{mb+a}{m+1}.$
+We use the set-up from Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$ and $\theta_{SOTA}=0.90$.
+We have that $E\max\hat{\Theta}(\mathbf{X})=E(E(\max\hat{\Theta}(\mathbf{X})|\Theta))$, and use Monte Carlo integration for the calculations.
+Table 1 and Figure 5 give an overview of some exemplifying models, details are in the following sections.
+Overview of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ for dependent and non-identical classifiers
+Section Independent Identical Upper 95% CI $\text{E}\operatorname{\hat{\theta}_{\max}}$ $\sigma_{\operatorname{\hat{\theta}_{\max}}}$
+3.4 YES YES 0.9213 0.9173 0.0018
+4.1 YES NO 0.9177 0.9129 0.0021
+4.2 NO YES 0.9207 0.9140 0.0035
+4.3 NO NO 0.9173 0.9101 0.0036
+Table 1: The upper bound of the CI, the expected value and standard deviation for $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ for different combinations of independency and identicality.
+$\theta_{SOTA}=0.90$, $n=3,000$, $m=1,000$.
+The correlation coefficient and $P(\theta)$ are found in their respective subsections.
+Bias and standard deviation for non-iid classifiers
+(a) Bias and standard deviation as a function of $b-a$ in $\mathcal{U}(a,b)$
+(b) Bias as a function of correlation coefficient.
+(c) Standard deviation as a function of correlation coefficient.
+Figure 5: Bias and standard deviation for $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ as a function of $d=b-a$ in $\mathcal{U}(a,b)$ for non-identical classifiers and of $\rho_{0}$ for conditional independent classifiers.
+The scale of the $y$-axes are kept the same as in Fig.
+4 for easier comparison.
+(a) The left tip of the curve is identical classifiers.
+The support of $\mathcal{U}(a,b)$ stretches out along the x-axis, and the bias decreases.
+The intersection with the solid vertical line corresponds to the cdf and pmf in Fig.
+6, and row 2 in Table 1.
+The intersections with the broken lines correspond to the left tips of the black lines in Fig.
+5(b) where the classifiers are independent.
+(b) The red line is identical classifiers, discussed in Section 4.2.
+The left tip of the red curve corresponds to the iid classifiers in Section 3.4, and the intersection with the vertical line corresponds to row 3 in Table 1.
+The black lines show bias as a function of correlation coefficient demonstrated for different $d$’s.
+The intersection between the solid black curve and the grey vertical line corresponds to row 4 in Table 1.
+(c) The corresponding standard deviation.
+Overall, Table 1 demonstrates that dependency and non-identical $\theta$s reduce the bias of the $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ estimator, and assumptions regarding dependency and identicality are crucial to include in a model.
+Figure 5 shows how the bias and standard deviation change as functions of $\Theta$ and $\rho_{0}$ for these models.
+4.1 Non-identical, independent classifiers
+We are interested in generalising Eq.
+2 to the non-identical case.
+We start by noting that we now have $P_{xj}=P(X\leq x|\theta_{j})$, and we seek the sum of those so that we can find the proper expression for $F(x|\Theta)=P(C_{x}=m|\Theta)$.
+The Poisson Binomial distribution describes the sum of $m$ non-identical independent Bernoulli trials with success probabilities $p_{1},p_{2},\ldots,p_{m}$ (Wang, 1993).
+The pmf for $K$ number of successes is
+$P(K=k)=\sum\limits_{{A\in G_{k}}}\prod\limits_{{j\in A}}p_{j}\prod\limits_{{i\in A^{c}}}(1-p_{i}),$
+where $G_{k}$ is all subsets of size $k$ that can be selected from {1, 2, …, m}. $G_{m}$ contains only one subset $A=\{1,2,\ldots,m\}$, and its complement $A^{C}$ is empty.
+We get the following cdf for non-identical classifiers:
+$F(x|\Theta)$ $=\sum\limits_{{A\in G_{m}}}\prod\limits_{{j\in A}}P_{xj}\prod\limits_{{i\in A^{c}}}(1-P_{xi})=\prod\limits_{j=1}^{m}P_{xj}.$ (4)
+The corresponding pmf is
+$f(x|\Theta)$ $=\prod\limits_{j=1}^{m}P_{xj}-\prod\limits_{j=1}^{m}P_{x\text{-}1j}$ (5)
+$=\prod\limits_{j=1}^{m}\sum\limits_{k=0}^{x}\binom{n}{k}\theta_{j}^{k}(1-\theta_{j})^{n\text{-}k}-\prod\limits_{j=1}^{m}\sum\limits_{k=0}^{x\text{-1}}\binom{n}{k}\theta_{j}^{k}(1-\theta_{j})^{n\text{-}k}.$
+Analytically, it is more complex than the case of identical classifiers, but it is easy to simulate. An example of cdf and pmf are displayed in Figure 6.
+Cdf and pmf for non-identical classifiers.
+(a) Cdf $F(z)$, see Eq. 4
+(b) Pmf $f(z)$, see Eq. 5
+Figure 6: The distribution and the probability function of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ for non-identical $\theta$’s, with $\theta_{j}\sim\mathcal{U}(a,b)$ where $d=b-a=0.025$.
+See Fig.
+2 for the corresponding distributions for identical classifiers.
+From row 2 in Table 1 we see that the bias is smaller, but the standard deviation is larger compared to the iid case.
+In our case, we can see this as the classifiers at the lower end not contributing to the sample maximum, and the reduced number of contributing classifiers gives a higher variance.
+The exact reduction/increase in bias and variance depends on $P(\theta)$, and can at its most extreme completely eradicate the multiplicity effect, as we shall see in Section 5.1.2.
+Figure 5(a) shows how bias and standard deviation changes with $d$ for fixed $\theta_{SOTA}=0.90$.
+The bias decreases rapidly at first, but the rate goes down fast.
+This means that taking into account that classifier performances are heterogeneous is important, and even small variations from identicality can have great impact on SOTA estimation.
+In Section 5, we demonstrate for other $\Theta$s, based on observations.
+4.2 Identical, dependent classifiers
+Even though the different teams work independently, their resulting classifiers are not necessarily statistically independent.
+A classifier’s ability to perform better than random guesses is due to dependencies between features and labels in the training sample.
+In their training phase, different classifiers will pick up on many of the same features and dependencies, regardless of which classifier they train.
+In addition, many teams will use different versions of related models, often pre-trained on the same data set.
+This scenario can be modelled by conditional independency.
+Let $Y_{0}$ be a Bernoulli variable with $P(Y_{0}=1)=\theta_{0}$.
+Let $Y_{j}$ be the Bernoulli variable with value $1$ if classifier $j$ predicts a data point correctly with $P(Y_{j}=1)=\theta_{j}$, and let $j=1,\ldots,m$ be the $m$ classifiers contributing to $\operatorname{\hat{\theta}_{\max}}$.
+Boland et al.
+(1989) described dependency in the majority systems context, where $\rho_{0j}=\text{corr}(Y_{0},Y_{j}),\,j=1,\ldots m$, and the $Y_{j}$s are independent given $Y_{0}$.
+$Y_{0}$ does not contribute directly to $\operatorname{\hat{\theta}_{\max}}$, but can be seen as, for example, the influence that a common training sample has on the classifiers.
+The binomial distribution for dependent Bernoulli trials is not an easy problem, see, for example, Ladd (1975) and Hisakado et al. (2006). We can model the pmf as a hierarchical distribution:
+$\operatorname{\hat{\theta}_{\max}(\mathbf{X})}|\Theta,Y_{0}$ $\sim P(\operatorname{\hat{\theta}_{\max}(\mathbf{X})}|\Theta,Y_{0})$
+$\theta\sim P(\theta),$ $\,\,\,Y_{0}\sim P(Y_{0}|\theta_{0}).$
+The Bernoulli trials can be simulated, see Boland et al.
+(1989) for conditional distributions.
+The sampling procedure is shown in Figure 7, illustrating the dependency step between the $\theta_{j}$’s and the $x_{j}$’s in Figure 1.
+Sampling procedure for the non-iid model
 […]
+Figure 7: The parameter $\theta_{0}$ defines the distribution from which the $Y_{0}$’s are sampled.
+The $\theta_{j}$’s are sampled from $P(\theta)$ as in Fig.
+1.
+Random samples of $Y_{j}$s are taken from their respective distributions, and $x_{j}$ follows directly, which in turn gives $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ as in Fig.
+1.
+$P(Y_{j}|\mathbf{y}_{0},\theta_{j})$ is described in Eq.
+6 and 7.
+We give an example for the case where $\rho_{0j}=\rho_{0}$ and $\theta_{j}=\theta$ for all classifiers, setting $\rho_{0}$ based on the work of Mania et al. (2019).
+Mania et al. (2019) calculated model similarity, defined as the probability of two classifiers giving the same output, which is equivalent to
+$(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}+2\rho_{jk}\sqrt{\theta_{j}\theta_{k}(1-\theta_{j})(1-\theta_{k})}$
+for models with correlation $\rho_{jk}$.
+The average accuracy for top-performing ImageNet models, $\hat{\theta}=0.756$, and the calculated similarity of $0.85$ gives $\rho=0.36$ under the assumption of equal correlation for all classifiers.
+We have that $\rho_{jk}=\rho_{0j}^{2}$, and hence our proposed correlation $\rho_{0}$ is set to $0.6$.
+Note that, for $\theta=0.90$, the similarity as described by Mania et al.
+(2019) is $0.82$ for two independent classifiers.
+When the probability of correct prediction is high, there is little room for variation between classifier outcomes.
+Let $\theta=0.90$ for all classifiers, and $\theta_{0}=0.90$.
+Comparing rows 1 and 3 in Table 1, we see a decrease in bias compared to the iid classifiers.
+However, the standard deviation has increased due to the random nature of $Y_{0}$, so the upper CI is essentially the same.
+To give an idea of how the randomness of $Y_{0}$ influences the standard deviation, we fix the $y_{0}$’s with exactly $\theta_{0}=0.90$ correct predictions.
+The expected value is the same as with random $Y_{0}$s, but the standard deviation has now decreased from $0.0035$ to $0.0015$.
+For real datasets, the dependencies have both random and fixed sources.
+The training and the test samples are random sources, a fixed source could be pre-trained models.
+For $\rho_{0j}=0$, we have independent classifiers, the same scenario as in Section 3.4, and for $\rho_{0j}=1$, we have $m$ classifiers producing the same output, and multiplicity is not a concern.
+For the whole range, see the red curve in Figure 5(b).
+We see here that the correlation coefficient value in a model has a great impact on the estimated bias if the correlation is substantial.
+This can explain that in mature fields, where most competing teams will use variants of the same model, the multiplicity will not impact the overestimation of SOTA as much as in young fields where the competing teams have more diverse models.
+This aligns with what is often observed: overoptimistic results in the beginning, and more sober as the field evolves.
+4.3 Non-identical, dependent classifiers
+The most complex, but also the most realistic case, is when classifiers are non-identical and dependent.
+An analytical solution is not known (Hisakado et al., 2006), but simulations are straightforward.
+The results of Boland et al.
+(1989) can be extended, and we have that
+$P(Y_{j}|Y_{0}=1)$ $=\frac{\rho\sqrt{\theta_{j}(1-\theta_{j})\theta_{0}(1-\theta_{0})}+\theta_{j}\theta_{0}}{\theta_{0}}$ (6)
+and
+$P(Y_{j}|Y_{0}=0)$ $=\frac{-\rho\sqrt{\theta_{j}(1-\theta_{j})\theta_{0}(1-\theta_{0})}+\theta_{j}(1-\theta_{0})}{(1-\theta_{0})}.$ (7)
+Note that, since $P\leq 1$, it follows from Eq. 7 that
+$\theta_{j}\geq\frac{\rho^{2}\frac{\theta_{0}}{1-\theta_{0}}}{1+\rho^{2}\frac{\theta_{0}}{1-\theta_{0}}},$ (8)
+and similarly for $0\leq P$ and Eq. 6.
+The results in Table 1 show that this non-identical, dependent model has the lowest bias, but the highest standard deviation.
+This sums of up much of Table 1, where more complex, and therefore more realistic, models reduces the bias but increases the variances.
+Figure 5(b) shows the impact of correlation for non-identical classifiers.
+The bias decreases slowly for low correlations, but the rate increases as $\rho_{0}$ increases.
+Note that Eq.
+8 restricts $d=b-a$, and the graphs are right-truncated.
+From Figure 4(a) we see that the impact from number of classifiers flattens off, and instead of restricting $\rho_{0}$ in our model, we can exclude classifiers by performance without large effect on the estimated bias.
+That is, restricting $d$ instead of $\rho_{0}$.
+What we also observe in Figure 5(a) is that change in $d$ has less impact on the higher end of the scale, so this also supports a decision of excluding teams with low performance without impacting the bias estimation too much.
+5 Real world examples
+As we have shown above, the sample maximum is an overoptimistic estimate for SOTA.
+These results are based on known SOTA, but faced with real world data the problem is of opposite nature: The sample maximum is known, but SOTA must be estimated.
+In the public challenge or public data set context there are many factors that come in to play, making bias calculations everything but straightforward.
+We here demonstrate different aspects of real world analysis through SOTA estimation of three different cases: Obesity Risk (5.1.1), Cassava Leaf (5.1.2), and Melanoma (5.2.3).
+Obesity Risk and Cassava Leaf have accuracy, defined as $\hat{\theta}$, as their ranking criterion, whereas Melanoma has the more common AUC, the area under the receiver operating characteristic (ROC) curve, as ranking criterion.
+Estimating multiplicity effect on the AUC’s is not as straightforward as with accuracy, but the mechanisms are the same.
+5.1 Estimating $\theta_{SOTA}$
+In a public challenge with $m$ participating teams, let $\hat{\Theta}=\{\hat{\theta}_{1},\hat{\theta}_{2},\ldots,\hat{\theta}_{m}\}$ be the estimated probability of success for classifiers $j=1,\ldots,m$ based on the performance on an independent test sample of size $n$.
+The winning team is identified by the sample maximum, $\max(\hat{\Theta})$.
+Among the different scenarios presented in Section 4, the non-identical dependent is the most realistic model.
+Whereas in Table 1 we report $\text{E}\operatorname{\hat{\theta}_{\max}}$ from known $P(\theta)$, the task is now to estimate $\theta_{SOTA}$ based on $\hat{\Theta}$.
+A straightforward approach is to propose a distribution $\theta^{\prime}_{j}\sim P(\theta^{\prime})$ whose expected sample maximum equals the observed sample maximum, $E\max(\hat{\Theta}^{\prime}(\mathbf{X}))=\max(\hat{\Theta})$.
+Deciding on $P(\theta^{\prime})$ is a crucial part, and we use an empirical Bayes approach, where we estimate $P(\theta^{\prime})$ from $\hat{\Theta}$.
+We base our suggested $P(\theta^{\prime})$ on shrinking the sample estimates, as seen in Figures 8(c) and 11(a), so that $E\max(\hat{\Theta}^{\prime}(\mathbf{X}))=\max(\hat{\Theta})$.
+The shrink parameter is found numerically.
+Cropping, an alternative approach, gave similar results to shrinking, and the results are presented in the Appendix.
+An alternative to the expected value is the upper limit of the 95% CI.
+The latter is more in line with common practices in research, and which to choose will depend on the aim of the analysis.
+We demonstrate the impact in the Obesity Risk example, reported in Table 2.
+Not every $\max(\hat{\Theta})$ is a result of multiplicity, and we demonstrate this in the Cassava Leaf example.
+5.1.1 Multi-Class Prediction of Obesity Risk
+Multi-Class Prediction of Obesity Risk is a challenge in the 2024 Kaggle Playground Series, using a synthetic data set generated from the Obesity or CVD risk data set (Palechor and de la Hoz Manotas, 2023).
+The test sample size was $n=13,840$, and the number of teams were $m=3,558$, excluding teams with accuracy lower than chance.
+Figure 8(a) shows the sample estimates, $\hat{\theta}_{j}$’s, zoomed in at the classifiers with sample estimates above $0.88$ for better visualisation.
+The maximum sample estimate is $\operatorname{\hat{\theta}_{\max}}=0.9116$, shown as the vertical dotted magenta line in all subfigures of Figure 8 together with the $\hat{\theta}_{j}$’s in gray, for reference.
+The 95% CI for the single $\operatorname{\hat{\theta}_{\max}}$ is shown as the magenta horizontal line in Figure 8(a), enveloping $921$ teams.
+This should raise suspicion that the right tail seen in Figure 8(a) can be the product of multiplicity.
+Obesity risk: $\Theta^{\prime}$ and corresponding realisations
+(a) $\hat{\theta}_{\max}$ and its CI. Row 1, Table 2
+(b) A realisation from a), $\text{E}\max(\hat{\Theta}^{\prime})$ and its CI. Row 2, Table 2.
+(c) Shrunk data, $\Theta^{\prime}$, and the new $\theta_{SOTA}$ (green). Row 3, Table 2.
+(d) A realisation from c), $\text{E}\max(\hat{\Theta}^{\prime})=\hat{\theta}_{\max}$ and its CI (red). Row 3, Table 2.
+Figure 8: The upper row shows the sample estimates and a corresponding realisation.
+The lower row shows a shrunk version and a corresponding realisation.
+For exact numbers, see Table 2.
+$\hat{\Theta}$ (gray) and $\hat{\theta}_{\max}$ (magenta) for reference.
+To demonstrate the impact of ignoring the multiplicity effect, we set the proposed distribution ${\theta}^{\prime}_{1},\ldots,{\theta}^{\prime}_{m}$ equal to the sample estimates, $\hat{\theta}_{j}$’s, and calculate the expected value of the sample maximum, $\text{E}\max(\hat{\Theta}^{\prime})$.
+We use the set-up from Section 4.3 with correlation coefficient $\rho_{0}=0.6$, and $P(Y_{0})=\theta_{0}=\max({\Theta}^{\prime})$.
+Note that sample estimates with low values are excluded in accordance with Eq.
+8, but their influence on the expected maximum is negligible.
+A realisation is shown in Figure 8(b).
+The upper tail, whose tip represents the maximum sample estimate, is stretched according to the variances of the $Y_{j}$s.
+The expected sample maximum is $\text{E}\max(\hat{{\Theta}}^{\prime})=0.9159$, far above the sample maximum in the challenge.
+The 95% CI is reported in Table 2.
+The lower bound exceeds $\max(\hat{\Theta})$, and we can conclude that the sample maximum is an unlikely candidate for SOTA.
+Instead, we can find a set $\Theta^{\prime}$ that results in realisations, $\hat{\Theta}^{\prime}$, similar to the challenge sample estimates in terms of sample maximum, and let $\max(\Theta^{\prime})$ be the SOTA candidate.
+To obtain a set $\Theta^{\prime}$ where $\text{E}\max(\hat{\Theta}^{\prime})=\max(\hat{\Theta})$, we shrink the challenge sample estimates
+$\theta^{\prime}_{j}=w\cdot\hat{\theta}_{j}+(1-w)\cdot\frac{1}{|classes|},$ (9)
+where $|classes|$ is the number of classes, and $0\leq w\leq 1$.
+The shrinking point, $1/|classes|$, is the random classifier.
+The weight, $w$, balances between the observations and the random element.
+Figure 8(c) shows the proposed $\Theta^{\prime}$.
+Figure 8(d) shows one realisation of the shrunk distribution, where $\max(\hat{\Theta}^{\prime})$ coincides with $\max(\hat{\Theta})$, and the corresponding CI.
+The empirical Bayes distribution, with $\text{E}\max(\hat{\Theta}^{\prime})=\max(\hat{\Theta})$, has $\max({\Theta}^{\prime})=0.9070$ and this serves as a new candidate for the SOTA.
+It can be argued that the new SOTA candidate is at risk of being an optimistic estimate, since the challenge observations are from only one test sample.
+The upper bound of the 95% CI can mitigate such a risk, and the challenge sample estimates are shrunk to fit that criterion.
+The results are shown in row 4 in Table 2.
+Estimating ${\theta}_{SOTA}$
+Approach $\max(\Theta^{\prime})$ 95% CI $\text{E}\max(\hat{\Theta}^{\prime})$ teams
+no mult. $\operatorname{\hat{\theta}_{\max}}$ (0.9067, 0.9162) - -
+$\Theta^{\prime}=\hat{\Theta}$ $\operatorname{\hat{\theta}_{\max}}$ (0.9128, 0.9191) 0.9159 -
+shrinking 0.9070 (0.9084, 0.9148) $\operatorname{\hat{\theta}_{\max}}$ 841
+0.9036 (0.9050, $\operatorname{\hat{\theta}_{\max}}$) 0.9082 1726
+Table 2: The 95% CI and the expected value for sample maximum, $\max(\hat{\Theta}^{\prime})$, for different SOTA estimation strategies, where SOTA is defined as $\max(\Theta^{\prime})$.
+The challenge sample maximum is $\operatorname{\hat{\theta}_{\max}}=0.9116$.
+The rightmost column shows number of classifiers with sample performance above the estimated SOTA.
+Table 2 demonstrates that taking the effect of multiplicity into account changes how results are perceived.
+The new SOTA value is $0.9070$, and even though its reduction from best performance is less than $0.005$, the effect it has on who we consider performing at SOTA-level is dramatic: around a quarter of the classifiers have sample performance above the new SOTA.
+The more conservative approach using the upper bound of the 95% CI to estimate SOTA gives an even more dramatic result, and either way the impact is substantial.
+5.1.2 Cassava Leaf Disease
+In this example, the maximum sample estimate is unlikely to be the product of multiplicity.
+For The Cassava Leaf Disease Classification challenge (Mwebaze et al., 2020), the task was to classify images of cassava leaves.
+The training sample consisted of $21,367$ images of five classes, and the test sample of approximately $15,000$ images.
+$3,752$ teams participated, excluding those with sample performance lower than chance.
+The three top performances appear as outliers: 0.9132, 0.9043 and 0.9028 as seen in Figure 9(a).
+Several explanations for the outlying performances have been put forward, and the interested reader can take a deep-dive in the comment section of the challenge.
+Since this is not the topic of the present work, we will not engage in the discussion here.
+The 95% CI of $\operatorname{\hat{\theta}_{\max}}$ is $(0.9086,0.9177)$, no other performance is within this interval.
+Already here we can decline a multiplicity-induced bias as a likely model.
+We can take one more step, and set $\Theta^{\prime}=\hat{\Theta}$.
+We get $\text{E}\max(\hat{\Theta}^{\prime})<\operatorname{\hat{\theta}_{\max}}$, and although the upper tail is stretched, its tip does not surpass $\operatorname{\hat{\theta}_{\max}}$, as illustrated in Figure 9(b).
+Cassava Leaf: $\Theta^{\prime}$ and corresponding realisations
+(a) Sample estimates and single CI for $\hat{\theta}_{\max}$
+(b) A simulated realization from a), the expected sample maximum and its CI in blue, $\hat{\theta}_{\max}$ in magenta for reference
+Figure 9: Cassava Leaf sample estimates as observed, and a corresponding simulated realisation.
+It is easy to see that the proposed shrinkage model in Eq.
+9 does not have a solution with the restriction that $\text{E}\max(\hat{\Theta}^{\prime})=\operatorname{\hat{\theta}_{\max}}$.
+This demonstrates how the suggested approach is inapplicable in an example where it is clearly not suited.
+5.2 Estimating $AUC_{SOTA}$
+In the following sections, we introduce a strategy for estimating $AUC_{SOTA}$, using simulated examples and the Melanoma Challenge for illustration.
+5.2.1 Simulation of AUC
+In a binary classification problem, the calculation of AUC is based on scores of random variables from the positive and the negative class.
+The AUC is mathematically equivalent to the probability that a random variable from the positive class will be ranked higher, in terms of score, than a random variable from the negative class (Hanley and McNeil, 1982).
+Let $score(\cdot)$ be a function that assigns a score to a random variable.
+Let $S_{+}=score(W_{+})$, where $W_{+}$ is a random variable from the positive class, and $S_{-}=score(W_{-})$, where $W_{-}$ is a random variable from the negative class.
+The distribution of classifier scores, $S$, is a mixture distribution of the $S_{+}$ and the $S_{-}$ distributions, where the weights, $\pi$ and $1-\pi$, correspond to the probability of drawing a positive or a negative random variable, and has pdf
+$\pi f_{+}(s)+(1-\pi)f_{-}(s),$
+where $f_{+}(s)$ and $f_{-}(s)$ are the pdfs of the positive and the negative class, respectively.
+Then $\text{AUC}=P(S_{-}<S_{+})=P(S_{-}-S_{+}<0)$, which is equivalent to the cdf at value 00.
+To simulate classifier scores according to a given AUC $=a$, we need to specify distributions for the variables $S_{-},S_{+}$ in such a way that $P(S_{-}-S_{+}<0)=a$.
+We use Normal distributions for illustration.
+Let $S_{+}$ and $S_{-}$ have the distributions
+$S_{+}$ $\sim\mathcal{N}(\mu_{+},\sigma^{2}_{+})$
+$S_{-}$ $\sim\mathcal{N}(\mu_{-},\sigma^{2}_{-}),$
+where $\mathcal{N}(\mu,\sigma^{2})$ is the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. We then have that
+$\text{AUC}=\Phi\left(\frac{0-(\mu_{-}-\mu_{+})}{\sqrt{\sigma_{-}^{2}+\sigma_{+}^{2}}}\right),$ (10)
+where $\Phi$ is the cdf of the standard Normal distribution.
+Figure 10(a) shows an example of two Normal distributions with means and variances chosen so that AUC = $0.90$.
+The rugs show a sample of size $n=3,000$ from the corresponding mixture distribution with $\pi=0.017$.
+The observations can be visualised as an ROC curve, and Fig 10(b) shows curves for $m=1,000$ samples.
+We consider the observed AUC for each sample, $\hat{\text{AUC}}_{j}$, $j=1,\ldots,m$, to be the sample estimates, analogous to $\hat{\Theta}=\{\hat{\theta}_{1},\ldots,\hat{\theta}_{m}\}$ in Section 3.3.
+The sample maximum AUC is $\max(\{\hat{\text{AUC}}_{1},\hat{\text{AUC}}_{2},\ldots,\hat{\text{AUC}}_{m}\})$, and is displayed as the black line in Fig 10(b).
+Figure 10(c) shows a simulated distribution of the sample maximum where the sample maximum from Fig 10(b) is close to its expected value.
+From score distributions to sample maximum AUC
+(a) Score distributions for a classifier: grey for negatives, black for positives. The rugs (bottom, top) show an example draw from each.
+(b) ROC curves for classifiers as in (a). The curves that fall within the inner quartile of all AUCs are dark grey, the ones outside are light grey. The black line is the sample maximum AUC of .958.
+(c) A simulated distribution of the sample maximum AUC. This is analogous figure to Figure 2(b) in Section 3.4.
 […]
-3.2 A coin-flip example
-Multiplicity is, simply put, how the probability of an outcome changes when an experiment is performed multiple times.
-Consider an experiment where a fair coin is flipped $n=20$ times and the number of successes is the number of times a classifier correctly predicts the outcome.
-Let the random variable $X$ denote the number of failures, $X\sim\mathcal{B}(n,1\text{-}\theta)$, where $\theta=0.5$.
-An unbiased estimator for $\theta$ is ${\hat{\theta}(X)=(n-X)/n}$.
-The corresponding estimate is referred to as accuracy.
-The probability of, say, $18$ correct predictions or more corresponds to $x=n-18$ failures or less, that is, the cumulative distribution $P(X\leq x|n,1\text{-}\theta)$.
-This corresponds to the cumulative distribution for the estimator $\hat{\theta}(X)=\frac{n\text{-}X}{n}$, and we have that $P(\hat{\theta}(X)\geq\frac{n\text{-}x}{n})=P(X\leq x)=0.00020$ for $\hat{\theta}\geq 0.90$.
-In other words, if $\theta=0.5$, there is a probability of $0.00020$ that the accuracy is $0.90$ or above in a series of $n=20$ trials.
-We can define the experiment described above as a Bernoulli trial, where success is defined as $X\leq x$, with corresponding probability $P(X\leq x)$, which we from now on write as $P_{x}$ for ease of notation.
-If we have $m$ classifiers, we have a series of $m$ Bernoulli trials, and the number of successes follow the binomial distribution $\mathcal{B}(m,P_{x})$.
-Let $C_{x}$ be the random variable that denotes the number of successful experiments, that is, the number of classifiers that correctly predicts at least $n-x$ coin flips.
-Of special interest is the case where at least one experiment succeeds, $P(C_{x}>0)=1-P(C_{x}=0)$.
-The true SOTA for coin-flip classifiers is $\theta_{SOTA}=0.5$, since it is impossible to train a classifier to predict coin flips.
-If $m=1,000$ classifiers independently give their predictions for $n=20$ flips, the probability of at least one classifier achieving an accuracy of at least $0.90$ is $0.181-P(C_{x}=0|m,P_{x})=$ 0.18.
-This means that using the top-ranked method as an estimate and ignoring multiplicity there is a substantial risk of grossly overestimating the SOTA.
-The same mechanism of multiplicity applies when $m$ teams submit their classifier’s result on a test set, although in a more complex manner.
-We seek to describe the probability distribution for increasing levels of complexity, and by that offer a tool to avoid overestimation.
-We will in the following use the public competition as a general example, and refer to the teams’ classifiers simply as classifiers, for ease of notation.
-3.3 The probability distribution of $\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})$
-We define the SOTA probability of success, $\theta_{SOTA}$, as the maximum among the $m$ classifiers being evaluated, $\theta_{SOTA}=\max(\Theta)$, $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$.
-Oftentimes, we only have access to sample estimates, $\hat{\Theta}=\{\hat{\theta}_{1},\ldots,\hat{\theta}_{m}\}$.
-Let $\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})=\max(\{\hat{\theta}_{1}(X),\hat{\theta}_{2}(X),\ldots,\hat{\theta}_{m}(X)\})$ be the naive, sample maximum estimator for $\theta_{SOTA}$.
-In the coin flip classifier example above, we have that $\theta_{SOTA}=0.5$, but $P(\hat{\theta}^{\,\max}_{SOTA}(\mathbf{X})\geq 0.90)=0.18$, and it is clear that we risk an overoptimistic estimate.
-Of interest is $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})|m,n,\Theta)$, which we will denote by $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ for ease of notation.
-Similarly, we use $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for $\operatorname{\hat{\theta}^{\,max}_{SOTA}}(\mathbf{X})$ when there is no risk of confusion.
-We here give $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ in the simple case of independent classifiers with identical probabilities of success: $\theta_{j}=\theta$ for all classifiers, $j=1,\ldots,m$.
-We discuss non-identical probabilities and dependency in Section 4.
-Let $X$ be the random variable that denotes the number of failures (wrong predictions) for a given classifier, with $X\sim\mathcal{B}(n,1\text{-}\theta)$.
-The random sample, $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ consists of the number of wrong predictions for each classifier $j=1,\ldots,m$ on a sample.
-Let $C_{x}$ be the random variable that denotes the number of classifiers with at most $x$ failures.
-We have that $C_{x}\sim\mathcal{B}(m,P_{x})$.
-We are interested in the probability of at least one classifier having at most $x$ wrong predictions, because it corresponds to the probability function of the $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ estimator, $P(\operatorname{\hat{\theta}^{\,max}_{SOTA}}\geq\frac{n-x}{n})$.
-We can derive this from $P(C_{x}>0|m,n,\theta)$.
-This corresponds to the cdf for the random variable $Z$ denoting the most wrong predictions among all classifiers:
-$F(z)$ $=P(C_{z}>0|m,n,\theta)=1-P(C_{z}=0|m,P_{z})$
-$=1-(1-P_{z})^{m}.$ (1)
-Let $f(z)=P(Z=z)$ denote the pmf of $Z$, the probability of at least one classifier having exactly $z$ wrong predictions, and no classifiers having fewer than $z$ wrong predictions. We then have
-$f(z)$ $=F(z)-F(z-1)=1-(1-P_{z})^{m}-\{1-(1-P_{z-1})^{m}\}$
-$=(1-P_{z-1})^{m}-(1-P_{z})^{m}=$
-$=(1-\sum_{\ell=0}^{z-1}P(Z=\ell))^{m}-(1-\sum_{\ell=0}^{z-1}P(Z=\ell)-P(Z=z))^{m}$
-$=-\sum_{k=1}^{m}(-1)^{k}\binom{m}{k}(1-\sum_{\ell=0}^{z-1}P(Z=\ell))^{m-k}P(Z=z)^{k}.$ (2)
-The variable $Z$ corresponds to $\hat{\theta}_{SOTA}^{\,max}$ in the sense that the smallest number of failures among all classifiers corresponds to the sample maximum probability of success.
-Equations 3.3 and 3.3 can be used to analyse $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for a given situation, as we do in the next section, where the impact that the three parameters have on the bias is also presented.
-3.4 A simulated public competition example
-The multiplicity mechanism that we described in the coin-flip example applies to public competitions where multiple teams submit the predictions made by their classifiers on a test set.
-The classifier with highest performance is declared winner, and is often referred to as SOTA later on.
-As exemplified in Section 3.2, multiplicity biases the naive sample maximum $\theta_{SOTA}$-estimator, and Section 3.3 gives us the tools to analyze its impact.
-In the following, we will use a significance level of $\alpha=0.05$ and two-sided CIs.
-Consider a classification problem with a test set of size $n=3,000$, and a classifier with probability of success $\theta=0.90$.
-The pmf of a single $\hat{\theta}(X)$ is shown as the blue curve in Fig.
-1, where the dotted line indicates $E(\hat{\theta})=\theta$.
-The 95% CI for $\hat{\theta}=0.90$ is $(0.8887,0.9105)$, and we denote it by $(\theta_{\alpha/2},\theta_{1\text{-}\alpha/2})$, displayed as a blue line on the horizontal axis.
-If there are $m=1,000$ independent classifiers, each with probability of success $\theta=0.90$, the probability of at least one classifier achieving at least $\hat{\theta}=\theta_{1\text{-}\alpha/2}$ is close to 1.
-This comes as no surprise given the definition of a CI.
-It means that we can be pretty certain that the sample performance of the top ranked classifier is above the upper limit of the CI of the expected performance.
-Consider a new classifier with probability of success $\theta_{m+1}=\theta_{1\text{-}\alpha/2}$.
-The corresponding pmf is displayed as the red curve in Fig.
-1, with the dotted line indicating $E(\hat{\theta}_{m+1})=\theta_{1\text{-}\alpha/2}$.
-By common scientific standards, the expected performance is considered significantly better than the expected performance of any of the $1,000$ classifiers with $\theta=0.90$, and is of interest to the research community.
-However, the multiplicity effect can conceal this.
-In the competition context, this is not necessarily a problem, since each classifier has the same probability of achieving the highest sample performance given that they have the same $\theta$.
-In a scientific context, using $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as an estimate for $\theta_{SOTA}$ without multiplicity corrections is problematic.
-The upper limit of the $(1\text{-}\alpha)$ CI of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ is $0.9213$, displayed on the horizontal axis in Fig.
-1, denoted as $\theta^{m}_{1\text{-}\alpha/2}$.
-We have $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})=0.9173$, calculated from Eq.
-3.3, also displayed on the horizontal axis in Fig.
-1.
-The probability of the new classifier beating the upper limit of the sample maximum CI is very low: $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})=0.02$, illustrated by the darker green area in Fig.
-1.
-Similarly, $P(\hat{\theta}_{m+1}\geq E(\operatorname{\hat{\theta}^{\,max}_{SOTA}}))=0.10$ seen as the entire green area in Fig.
-1.
-This means that by using the sample maximum as an estimator for SOTA, we can expect that 9 out of 10 classifiers that are significantly better will still be naively considered worse.
-The pmfs of two single $\hat{\theta}(X)$ compared to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ statistics
-Figure 1: Against a mass of identical classifiers, a single significantly better classifier has little chance of beating the state of the art: The blue and the red curves show the pmfs of $\hat{\theta}(X)$ for classifiers with probability of success $\theta$ and $\theta_{1\text{-}\alpha/2}$, respectively.
-$\theta_{1\text{-}\alpha/2}$ is the upper limit of the CI of $\hat{\theta}(\mathbf{x})=\theta$.
-The CI is displayed as the blue horizontal line.
-$E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ and $\theta^{m}_{1\text{-}\alpha/2}$ denote the expected value and the upper limit of the CI of the sample maximum estimator for $m=1,000$ classifiers.
-The shaded areas corresponds to $P(\hat{\theta}_{m+1}\geq\theta^{m}_{1\text{-}\alpha/2})$ and $P\big{(}\hat{\theta}_{m+1}\geq E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})\big{)}$, and reflects the slim chances that the single better classifier has of beating the multiple worse classifiers.
-For the entire distribution of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, see Fig. 2.
-Cdf and pmf for at least one classifier having at most $z$ failures, corresponding to the sample maximum estimator
-(a) Cdf $F(z)$, see Eq. 3.3
-(b) Pmf $f(z)$, see Eq. 3.3
-Figure 2: The distribution and the probability function of the sample maximum estimator for $n=3,000$, $\theta=0.90$, and $m=1,000$.
-The horizontal axes denotes the number of failures, $z$, at the top and the corresponding sample maximum estimates, $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, at the bottom.
-Note that the horizontal axis is reversed compared to Fig.
-1.
-The distribution of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ will vary with $m$, $n$ and $\theta$, and Table 1 shows the expectation and standard deviation for a small collection of examples.
-To complement the table, Fig.
-3 a)-c) show the bias of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as a function of $m$, $n$ and $\theta$, respectively.
-Expected values and standard deviations for $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$
-m n $\theta$ $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ $\sigma_{\operatorname{\hat{\theta}^{\,max}_{SOTA}}}$
+5.2.2 Simulation results for uncorrelated classifiers
+To demonstrate the impact that multiplicity can have, we start with a simulated example similar to the one in Section 3.4.
+Consider a classification problem with $n=3,000$, $m=1,000$, where each classifier has an AUC of $a=0.90$, and the classifiers are independent.
+Whereas the distribution of the sample maximum for probability of success is independent of class balance as long as the probability of success is independent of class, this is not true for the sample maximum of AUC, where the variance increases with class imbalance.
+We have chosen the same class balance as in the Melanoma challenge, $\pi=0.017$, as this reflects real world examples.
+For simplicity we choose $\mu_{-}=0$ and $\sigma_{-}=\sigma_{+}=1$.
+As we see from Eq.
+10 the AUC is completely decided by the ratio between the difference in means and the pooled standard deviation.
+Changing the scale of one implies the same change of scale in the other.
+Perhaps the strongest assumption is that of equal variances, but we see unrealistically variable AUCs from letting the difference between the two be too large, so keeping them equal was the simplest choice.
+We then solve for $\mu_{+}$ numerically.
+Now we can generate scores from the resulting Normal distributions.
+Figure 10(c) shows the simulated distribution of the sample maximum AUC over $10,000$ repetitions of a challenge where every classifier has AUC of $0.90$.
+The simulated expected value is $0.9562$, the standard deviation is $0.004459$, and the simulated 95% CI is $(0.9486,0.9662)$.
+Comparing this to the results in the upper row of Table 3 , we see that the bias is much larger for the sample maximum AUC.
+This is partly due to the large variance created by the class imbalance.
+Although the sample size is $n=3,000$, only $51$ of those belong to the positive class, creating large variance in the estimator.
+5.2.3 Melanoma Classification
+We will use the Melanoma Classification challenge from 2020 as an illustrative example of $AUC_{SOTA}$ estimation because it follows many of the common structures of public challenges.
+The SIIM-ISIC Melanoma Classification (Zawacki et al., 2020) challenge consisted of more than $40,000$ images of skin lesions from various sites, split into training ($33,126$), validation ($3,295$) and test set ($7,687$).
+The data set is highly imbalanced, with less than $2\%$ melanomas, which reflects the reality in many cancer detection situations.
+A full description can be found in Rotemberg et al.
+(2021).
+The size of the test set is smaller than the recommendations of, for example, Roelofs et al.
+(2019), but reflects the reality of benchmark datasets and public challenges, as seen in, for example, the overview of Willemink et al.
+(2020).
+We want to emphasise that the shortcomings of this challenge are not unique, and that is precisely why it serves as a useful example.
+The prize money was set to $10,000$ USD for the best-performing method, measured as AUC on the test set.
+Figure 11(a) shows the results.
+The team ‘All Data Are Ext’, consisting of three Kaggle Grandmasters, won the challenge, achieving an AUC of $0.9490$, and their method and strategy is explained in Ha et al.
+(2020).
+Although we show that the AUC is an optimistic estimate for the melanoma detection SOTA, we do not question their rank in the challenge.
+Quite the contrary; the winning team chose many of the verified strategies from Section 2.
+They used additional data from previous years’ challenges, and they also did data augmentation.
+They chose ensemble learning as a strategy to avoid overfitting, a strategy implicitly supported by Talagrand (1996).
+They also trained on multiple classes, a strategy that Feldman et al.
+(2019) showed will slow down the overfitting.
+The observed maximum is an unlikely SOTA candidate, and to give a better estimate of the SOTA for the Melanoma challenge, we follow the same procedure as in Section 5.1: The challenge observations are shrunk in such a manner that simulating from the shrunk set, $AUC^{\prime}$, gives an expected sample maximum AUC of $0.9490$, the challenge sample maximum.
+Use of shrinkage distribution to suggest a SOTA
+(a) Shrinkage AUC distribution (black) based on the Melanoma challenge results (grey).
+The vertical red line shows the highest Melanoma challenge AUC of .949.949.949.949 and the horizontal red line shows a $95\%$ CI around this.
+(b) A sample from the shrinkage distribution of (a).
+The vertical line is the expected sample maximum, engineered to coincide with the observed Melanoma challenge maximum of $0.949$.
+The horizontal line shows the $95\%$ CI of the sample maximum.
+Figure 11: Histograms for AUC sample estimates shrunk toward $0.5$.
+In the shrinkage distribution the state of the art is an AUC of $0.8992$ which leads to the expected sample maximum AUC of 0.9490.
+AUCs below .6.6.6.6 not shown to make the more interesting upper region more readable.
+Figure 11(a) shows the AUCs of the Melanoma challenge test sample shrunk toward $0.5$, with maximum $0.8992$.
+Figure 11(b) shows a corresponding sample, where the maximum is close to $0.9490$.
+There are 2013 teams with performance above the new SOTA candidate.
+The discrepancy between the sample maximum and the new SOTA candidate is big and has big implications.
+From an AUC close to $95\%$ to barely $90\%$ is a substantial drop in performance.
+Almost two thirds of the classifiers have performances above the new SOTA estimate.
+As was shown in Section 4, illustrated in Figure 5(b), the bias is reduced for correlated classifiers, so the analysis assuming independence overestimates the bias.
+In this challenge, we can assume multiple sources of dependency: Access to the same training set, models pre-trained on the same data, similar models, etc.
+However, the wide CI for a single classifier tells the story of an estimator with large variance, and hence the bias of the sample maximum will nevertheless be substantial.
+5.3 Future work
+There are several interesting aspects that has not been addressed in this work, that the authors briefly touched upon while preparing the manuscript.
+5.3.1 Correlation between classifiers in the AUC simulations
+For the moment, we have not extended our AUC simulations to include correlations, but it should be possible to adopt a similar setup as in Sections 4.2 and 4.3: each classifier produces scores that correlate to those of a reference classifier.
+Let $S_{0^{+}}$ and $S_{0^{-}}$ be the scores of the reference classifier, for random variables from the positive and negative class, respectively.
+Let $\rho_{+}=corr(S_{0^{+}},S_{j^{+}})$ be the correlation between the reference classifier scores $S_{0^{+}}$ and the scores of classifier $j$, $S_{j^{+}}$ for the positive class.
+If they both follow Normal distributions as outlined above, then the bivariate distribution is
+$\begin{pmatrix}S_{0^{+}}\\ S_{j^{+}}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mu_{0^{+}}\\ \mu_{j^{+}}\end{pmatrix},\begin{pmatrix}\sigma_{0^{+}}^{2}&\rho\sigma_{0^{+}}\sigma_{j^{+}}\\ \rho\sigma_{0^{+}}\sigma_{j^{+}}&\sigma_{j^{+}}^{2}\end{pmatrix}\right),$
+and equivalently for the negative class. However, the joint distribution of $(S_{0},S_{j})$ is not uniquely defined by the correlation, and additional assumptions must therefore be made.
+I
+5.3.2 Approximate Bayesian Posterior for SOTA
+In Section 5 we consider observed data, $\mathbf{x}$, as coming from a hierarchical model like that of Figure 1 and try to find a $P(\theta)$ with expected maximum coinciding with $\max(\mathbf{x})$.
+If we frame this as a Bayesian argument we could say that we are tying to find a prior that aligns well with our observed data and our model for the data.
+To find a prior we consider a very simple family of discrete distributions based on the empirical distribution of our data.
+In the usual Bayesian arguments the data model is expressed as a likelihood, $f(\mathbf{x}\mid\theta)$.
+We rely instead on simulations for the more complex models above.
+We could instead consider a Bayesian analysis where we specify a prior for $\theta$ and try to find a posterior distribution for $\max(\theta)$.
+Approximate Bayesian Computation (Rubin, 1984) is quite close to what we do above in that instead of a likelihood the analysis uses a simulation of the data generating process.
+Specifically we could specify some weakly informative prior, $P(\theta)$, simulate data, $\mathbf{x}^{\prime}$, based on a draw, $\bar{\theta}\sim P(\theta)$, and check if $|\max(\mathbf{x}^{\prime})-\max(\mathbf{x})|<\epsilon$ for some acceptable $\epsilon>0$.
+If yes, we consider $\max(\bar{\theta})$ a draw from the posterior distribution, $P(\max(\theta)\mid\mathbf{x})$, otherwise we discard it.
+Repeating this would give us a posterior distribution for $\max(\theta)$ without us having to specify some ad-hoc family of priors, requiring instead a single prior distribution and the specification of $\epsilon$.
+6 Discussion
+Some of the insights presented here are obvious to anyone with a basic understanding of probability, like the introductory coin-flip example in Section 3.2.
+The mechanisms we have described are similar to those we find in regression to the mean, a well-known phenomenon in various fields such as livestock breeding and sport results.
+We believe that this has been neglected mainly because of the lack of formalisation: What is meant by SOTA performance, and what question does it answer?
+The main contribution has been to phrase the question, both the one that is commonly answered by reporting the maximum performance, and the question that can give an unbiased estimate of the SOTA performance.
+Although our models span from the special case of iid classifiers to heterogeneous and correlated ones, the analysis is limited by the assumptions one has to make regarding distributions and correlation coefficients.
+An important lesson for practical applications is that simpler models tend to overestimate the bias, and hence underestimate the SOTA.
+Looking at the application to real data in Section 5, we see that the implications of shifting the focus from best sample performance to best classifier performance has huge implications.
+Not necessarily so much in the new SOTA performance estimate, like the $0.005$ shift in Section 5.1.1, but definitely in which classifiers we consider performing at SOTA level - in this example, $25\%$ of the participating teams.
+In the Melanoma example in Section 5.2.3, the decrease in estimated SOTA performance is more striking, and we must keep in mind the wisdom from before, that simpler models tend to overestimate the bias.
+However, any model will uncover bias, and any bias will have quite dramatic results in terms of number of teams performing at SOTA level.
+What does this mean for how we view the results from public challenges, such as Kaggle?
+First, we would like to emphasize that we appreciate the positive contributions that public challenges make, and that this is meant as a modification of how their results are viewed in the aftermath.
+The beauty of the approach presented in this work - that we do not even have to know which classifier is the best to estimate the best performance - becomes the curse for the public challenge: We do not know which team is the best, only which was good combined with luck.
+The approach described here offers a method to quantify that luck, or uncertainty as we would name it in statistics.
+Like anyone with a basic understanding of probability can draw the conclusions in the coin-flip example, anyone with a basic understanding of statistics will look at the distribution of observed performances, as in Figure 8(a), and suspect that the right tail is a product of sample variance.
+This work offers a way of connecting the dots from coin-flip to public challenge, and reveals some of the plentiful hurdles along that route.
+In some fields of science, where hundreds of thousands of tests are performed simultaneously, the effects of multiplicity simply cannot be ignored.
+But as we see from Figures 4(a) and 4(b), this is not a thousand-classifier problem, it is a two-classifier problem.
+As soon as you have two classifiers, the observed maximum is appreciably biased.
+It is obvious from Figure 8(d) that the chosen approach to find a good empirical distribution is not optimal.
+A sample from the empirical distribution in Figure 8(c) should ideally result in a sample similar to the observed sample estimates, that is, an overlap between the black and the gray histograms in Figure 8(d).
+Whereas this is not that difficult to achieve, it requires more parameters and more tweaking, and does not result in any generalizable model.
+When deep-diving into a specific problem, it can be worthwhile spending some extra time refining those models, and presenting different results depending on model choice.
+There are several approaches that can contribute to the problem of multiplicity, as presented in Section 2.
+Common for the solutions presented there are that they require additional data or more elaborate set-ups for public challenges.
+The approach presented here offers a tool that can be applied to already published results, and does not require convincing anyone else to change their preferred approach to finding the best solution.
+Appendix A Numerical results
+Expected values and standard deviations for $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$
+[…]
 1,000 3,000 0.90 0.9173 0.001817
@@
 1,000 3,000 0.95 0.9624 0.001277
-Table 1: The expected value and standard deviation for the SOTA estimator with varying test set size, number of classifiers and probability of correct success.
-The bold line is the example used in Fig.
-2 and corresponds to the intersection between the gray vertical lines and the solid lines in Fig.
-3.
-For each curve, the two other parameters are kept constant, and each figure displays five curves, corresponding to Table 1.
-The bias of the sample maximum estimator
-Figure 3: The bias of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ as a function of each of its three parameters; number of classifiers, test set size and probability of success.
-The intersection between the black solid lines and the gray vertical lines correspond to the bold line in Table 2.
-The dotted and dashed lines correspond to variation of parameter values found in Table 2, colour-coded accordingly.
-In Fig.
-3(a), we see that the impact of number of classifiers is dramatic, and multiplicity-considerations must be treated with full attention starting at $m=2$.
-We can see this from the cdf in Eq.
-3.3, where $m$ is the exponent.
-In Fig.
-3(b) is that the test set size also has large impact on the bias.
-With small test sets, the variance of each individual $\hat{\theta}_{j}(X)$ is high, increasing the probability of sample estimates far from $\theta$.
-As seen in Fig.
-3(c), the bias decreases as $\theta$ approaches 1.
-The variance of the individual $\hat{\theta}_{j}(X)$s is at its highest for $\theta=0.5$ and decreases monotonically, and it is this variance that is reflected in the bias, similar to the effect of the test set size.
-The change in absolute numbers may seem modest, but the practical implications are large, as demonstrated in Section 5, where we analyse the results from three public competitions.
-4 Dependency and non-identical $\theta$s
-The probability distribution in Section 3.4 is that of independent classifiers with identical $\theta$s, an assumption that is rarely true in practice.
-We expand our investigation to non-identical $\theta$s and dependency.
-Table 2 and Figure 5 give an overview of some exemplifying modelss, details are in the following sections.
-Characteristics of $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ for dependent and non-identical classifiers
-Section Independent Identical Upper 95% CI $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ $\sigma_{\operatorname{\hat{\theta}^{\,max}_{SOTA}}}$
-3.4 YES YES 0.9213 0.9173 0.0018
-4.1 YES NO 0.9177 0.9130 0.0021
-4.2 NO YES 0.9207 0.9140 0.0035
-4.3 NO NO 0.9173 0.9101 0.0036
-Table 2: The upper bound of the 95% CI, the expected value and standard deviation for the sample maximum estimator for different combinations of independency and identicality.
-The true SOTA is $\theta=0.90$, the test set size is $n=3,000$, the number of classifiers is $m=1,000$.
-The correlation coefficient and the $\Theta$-vector are found in their respective subsections.
-Bias for non-identical and dependent classifiers.
-(a) Bias as a function of $\min(\mathbf{\Theta})$, with constant $\max(\mathbf{\Theta})=0.90$
-(b) Bias as a function of correlation coefficient.
-Figure 4: Bias of the sample maximum estimator as a function of $\min(\Theta)$ for non-identical classifiers and of $\rho_{0}$ for conditional independent classifiers.
-(a) The right tip of the curve is identical classifiers, and following the curve from right to left the $\mathbf{\Theta}$-vector stretches out, and the bias decreases.
-The intersection with the solid gray vertical line corresponds to the cdf and pmf in Fig.
-5, and produces the CI, the expectation and the standard deviation in line two in Table 2.
-The intersections between the gray vertical lines correspond to the left tips of the black lines in Fig.
-4(b), where the classifiers are independent.
-(b) The red line are identical classifiers, as discussed in Section 4.2.
-The left tip of the red curve corresponds to the i.i.d.
-classifiers in Section 3.4, and the intersection with the grey vertical line corresponds to line three in Table 2.
-The black lines show bias as a function of correlation coefficient demonstrated for different $\mathbf{\Theta}$-vectors.
-The intersection between the solid black curve and the grey vertical line corresponds to line four in Table 2.
-The scale of the $y$-axis is kept the same as in Fig.
-3 for easier comparison of which parameters have the most impact.
-Overall, Table 2 demonstrates that dependency and non-identical $\theta$s reduces the bias of the $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$ estimator, and assumptions regarding dependency and identicality are crucial to include in a model.
-Figure 4 shows how the bias changes as a function of $\Theta$ and $\rho_{0}$ for these specific models.
-We use the set-up from Section 3.4 with $m=1,000$ classifiers, test set size $n=3,000$, $\theta_{SOTA}=0.90$, and $100,000$ repetitions when simulations are needed.
-We refer to classifiers with identical $\theta$s as identical classifiers.
-4.1 Non-identical, independent classifiers
-We are interested in generalising Eq.
-3.3 to the non-identical case, and we start by noting that we now have $P_{xj}=P(X\leq x|\theta_{j})$, and we seek the sum of those so that we can find the proper expression for
-$F(z|\mathbf{\Theta})=1-P(C_{z}=0|m,n,\mathbf{\Theta}).$
-The Poisson Binomial distribution, see, e.g., Wang1993, describes the sum of $m$ non-identical independent Bernoulli trials with success probabilities $p_{1},p_{2},\ldots,p_{m}$.
-The pmf for $K$ number of successes is
-$P(K=k)=\sum\limits_{{A\in G_{k}}}\prod\limits_{{i\in A}}p_{i}\prod\limits_{{j\in A^{c}}}(1-p_{j}),$ (3)
-where $G_{k}$ is all subsets of size $k$ that can be selected from {1, 2, …, m}.
-We then get the cdf for non-identical classifiers
-$F(z|\mathbf{\Theta})$ $=1-\sum\limits_{{A\in G_{0}}}\prod\limits_{{i\in A}}P_{zi}\prod\limits_{{j\in A^{c}}}(1-P_{zj})=1-\prod\limits_{j=1}^{m}\big{(}1-P_{zj}\big{)}.$ (4)
-The pmf is
-$f(z|\mathbf{\Theta})$ $=\prod\limits_{j=1}^{m}\big{(}1-P_{z\text{-}1j}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-P_{zj}\big{)}$ (5)
-$=\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z\text{-1}}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}-\prod\limits_{j=1}^{m}\big{(}1-\sum\limits_{k=0}^{z}\binom{n}{k}(1-\theta_{j})^{k}\theta_{j}^{n\text{-}k}\big{)}.$
-Analytically, it is more complex than the case of identical classifiers, but it is easy to simulate.
-We use the same parameters as in Section 3.4, replacing $\theta=0.90$ with the vector $\mathbf{\Theta}$, where $\min(\mathbf{\Theta})=0.875,\,\max(\mathbf{\Theta})=0.90$ and equally spaced elements.
-The cdf and pmf are displayed in Fig.
-5(a) and Fig.
-5(b).
-Cdf and pmf for non-identical classifiers.
-(a) Cdf $F(z)$
-(b) Pmf $f(z)$
-Figure 5: The distribution and the probability function of the sample maximum estimator for non-identical $\theta$s, where $\min(\mathbf{\Theta})=0.875,\,\max(\mathbf{\Theta})=0.90$ and equally spaced elements.
-The upper bound of the 95% CI is $0.9177$.
-The $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ is $0.9129$, with standard deviation of $0.0021$.
-Comparing this to the results in Section 3.4, we see that the bias is smaller and the variance larger.
-This is because the classifiers at the lower end does not contribute to the sample maximum, and the reduced number of contributing classifiers gives a higher variance.
-The exact reduction/increase in bias and variance depends on $\mathbf{\Theta}$, and can at its most extreme cases completely eradicate the multiplicity effect, as we shall see in Section 5.1.2.
-Figure 4(a) shows how $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})$ changes with $\min(\mathbf{\Theta})$ for equally spaced elements with $\max(\mathbf{\Theta})=0.90$.
-What we see here is that with non-identical classifiers the bias decreases rapidly at first as $\min(\mathbf{\Theta})$ moves away from $\max(\mathbf{\Theta})$, but the rate at which this happens also decreases rapidly.
-This means that taking into account that classifier performances are heterogeneous is important, and small variations from identicality can have great impact on SOTA estimation.
-In Section 5, we demonstrate for other $\mathbf{\Theta}$s, based on observations.
-4.2 Identical, dependent classifiers
-Even though the different teams work independently, their resulting classifiers are not necessarily statistically independent.
-A classifier’s ability to perform better than random guesses is due to the correlations between features and labels in the training set.
-In their training phase, different classifiers will pick up on many of the same features and correlations, regardless of which classifier they train.
-In addition, many teams will use different versions of related models, often pre-trained on the same data set.
-This scenario can be modelled by conditional independency.
-Let $Y_{0}$ be a Bernoulli variable with $P(Y_{0}=1)=\theta_{0}$.
-Let $Y_{j}$ be the Bernoulli variable with value $1$ if classifier $j$ predicts a data point correctly with $P(Y_{j}=1)=\theta_{j}$, and let $j=1,\ldots,m$ be the $m$ classifiers contributing to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$.
-Boland et al.
-(1989) described dependency in the majority systems context, where $\rho_{0j}=\text{corr}(Y_{0},Y_{j}),\,j=1,\ldots m$, and the $Y_{j}$s are independent given $Y_{0}$.
-The observations from $Y_{0}$ does not contribute directly to $\operatorname{\hat{\theta}^{\,max}_{SOTA}}$, but can be seen as, e.g., the influence that a common training set has on the classifiers.
-The binomial distribution for dependent Bernoulli trials is not an easy problem, see, e.g., Ladd (1975) and Hisakado et al.
-(2006).
-The Bernoulli trials can be simulated, see Boland et al.
-(1989) for conditional distributions.
-We give an example below for the case where $\rho_{0j}=\rho_{0}$ and $\theta_{j}=\theta$ for all classifiers.
-We set $\rho_{0}$ based on the work of Mania et al.
-(2019).
-Mania et al. (2019) calculated model similarity, defined as the probability of two classifiers giving the same output, which is equivalent to
-$(1-\theta_{j})(1-\theta_{k})+\theta_{j}\theta_{k}+2\rho_{jk}\sqrt{\theta_{j}\theta_{k}(1-\theta_{j})(1-\theta_{k})}$ (6)
-for dependent models with correlation $\rho_{jk}$.
-The average accuracy for top-performing ImageNet models, $\hat{\theta}=0.756$, and the calculated similarity of $0.85$ gives $\rho=0.36$ if we assume the same correlation for all classifiers.
-We have that $\rho_{jk}=\rho_{0j}^{2}$, and hence our proposed correlation $\rho_{0}$ is set to $0.6$.
-Note that, for $\theta=0.90$, the similarity as described by Mania et al.
-(2019) is $0.82$ for two independent classifiers.
-When the probability of correct prediction is high, there is little room for variation between classifier outcomes.
-Let $P(Y_{j}=1)=\theta=0.90$ for all classifiers, and $P(Y_{0}=1)=\theta=0.90$.
-From our simulations, we get $E(\operatorname{\hat{\theta}^{\,max}_{SOTA}})=0.9140$, a decrease in bias compared to the independent and identical classifiers.
-However, the variance has increased due to the random nature of $Y_{0}$, so the upper 95% CI is essentially the same.
-See, Table 2 for exact numbers.
-To give an idea of how the randomness of $Y_{0}$ influences the variance, we fix the $y_{0}$’s with exactly $(n-x)/n=0.90$ correct predictions.
-The expected value is the same as with random $Y_{0}$s, but the standard deviation has now decreased from $0.0035$ to $0.0015$.
-In the real world, the dependencies have both random and fixed sources.
-A random source would, e.g., be the training set and the test set, random samples from an underlying distribution.
-A fixed source can be, e.g., fixed training conditions, such as pre-trained models.
-For $\rho_{0j}=0$, we have independent classifiers, the same scenario as in Section 3.4, and for $\rho_{0j}=1$, we have $m$ classifiers producing the same output, and multiplicity is not a concern.
-For the whole range, see the red curve in Figure 4(b).
-What we see here is that the correlation coefficient value in a model has a great impact on the estimated bias if the correlation is substantial.
-This can explain that in mature fields, where most competing teams will use variants of the same model, the multiplicity will not impact the overestimation of SOTA as much as in young fields where the competing teams have more diverse models.
-This aligns with what is often observed in science: overoptimistic results in the beginning, and more sober as the field evolves.
-4.3 Non-identical, dependent classifiers
-The most complex, but also the most realistic case, is where classifiers are non-identical and dependent.
-An analytical solution is not known Hisakado et al.
-(2006), but simulations are straightforward.
-The results of Boland et al.
-(1989) can be extended, and we have that
-$P(Y_{j}|Y_{0}=1)$ $=\frac{\rho\sqrt{\theta_{j}(1-\theta_{j})\theta_{0}(1-\theta_{0})}+\theta_{j}\theta_{0}}{\theta_{0}}$ (7)
-$P(Y_{j}|Y_{0}=0)$ $=\frac{-\rho\sqrt{\theta_{j}(1-\theta_{j})\theta_{0}(1-\theta_{0})}+\theta_{j}(1-\theta_{0})}{(1-\theta_{0})}.$ (8)
-Note that, since $P\leq 1$, it follows from Eq. 8 that
-$\theta_{j}\geq\frac{\rho^{2}\frac{\theta_{0}}{1-\theta_{0}}}{1+\rho^{2}\frac{\theta_{0}}{1-\theta_{0}}},$ (9)
-and similarly for $0\leq P$ and Eq. 7.
-We get an upper bound of the $95\%$ CI of $0.91733$ and an expected value of $0.9101$, which are lower than both the dependent and the non-identical model separately.
-The standard deviation is $0.0036$, which is higher than the two other models.
-This sums of up much of Table 2, where more complex, and therefore more realistic, models reduces the bias but increases the variances.
-Figure 4(b) shows the impact of correlation for non-identical classifiers.
-What we see is that the bias decreases slowly for low correlations, but that the rate increases as $\rho_{0}$ increases.
-Importantly, Eq.
-9 restricts the discrepancy between $\min(\mathbf{\Theta})$ and $\max(\mathbf{\Theta})$, and the graphs are right-truncated.
-From Fig.
-3(a) we observe that the impact from number of classifiers flattens off, instead of restricting $\rho_{0}$ in our model, we can exclude classifiers by performance without large effect on the estimated bias.
-That is, restricting $\min(\mathbf{\Theta})$ instead of $\rho_{0}$.
-What we also observe in Fig.
-4(a) is that change in $\min(\mathbf{\Theta})$ has less impact on the lower end of the scale, so this also supports a decision of excluding teams with low performance without impacting the bias-estimation too much.
-5 Real world examples
-As we have shown above with simulated data, the sample maximum is an overoptimistic estimate for SOTA.
-These results are based on known SOTA, but faced with real world data the problem to be solved is of opposite nature: The sample $\max(\hat{\Theta})$ is known, but the true SOTA must be estimated.
-In the public competition or public data set context there are many factors that come in to play, making the bias calculations everything but straightforward.
-We here demonstrate different aspects of real world analysis through SOTA estimation of three different cases: Obesity Risk (5.1.1), Cassava Leaf (5.1.2, and Melanoma (5.2).
-Obesity Risk and Cassava Leaf have accuracy, defined as $\hat{\theta}$, as their ranking criterion, whereas Melanoma has area under the receiver operating characteristic (ROC) curve (AUC) as ranking criterion.
-To estimate multiplicity effect on the AUC scores is not as straightforward as with accuracy, but the mechanisms are the same and AUC is a more common ranking criterion than accuracy.
-In a public competition with $m$ participating teams, let $\hat{\Theta}=\{\hat{\theta}_{1},\hat{\theta}_{2},\ldots,\hat{\theta}_{m}\}$, be the estimated probability of success for each classifier based on the performance on an independent test sample of size $n$.
-The the winning team is identified by the sample maximum, $\max(\hat{\Theta})$.
-Among the different scenarios presented in Section 4, the non-identical dependent is the most realistic model.
-Whereas in Table 2 we report $E(\hat{\theta}^{\,max}_{SOTA})$ from known $\Theta$, the task is now to estimate $\theta_{SOTA}$ based on $\hat{\Theta}$.
-A straightforward approach is to find $\Theta^{\prime}=\{\theta^{\prime}_{1},\theta^{\prime}_{2},\ldots,\theta^{\prime}_{m}\}$ that gives $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$.
-That is, find a set of $\theta$’s so that the expected value of its sample maximum on a test set of size $n$ is equal to the competition sample maximum.
-The SOTA estimate is then $\max(\Theta^{\prime})$.
-We use a set-up where the simulated $\hat{\theta}^{\prime}_{j}$s are the mean values of $n$ realisations of $Y_{j}$, where $P(Y_{j})={\theta}^{\prime}_{j}$.
-The simulation is repeated $10,000$ times, and for each repetition the observed maximum value is recorded.
-$E(\max(\hat{\Theta}^{\prime}))$ is the mean value over the repetitions.
-In the studies in Section 4, the $\theta_{1},\ldots,\theta_{m}$ were chosen to be equally spaced, but this does not align with the real world sample estimates, as seen in Figures 6(a), 7(a) and 8.
-We base our suggested ${\theta}^{\prime}_{j}$’s on cropping the sample estimates, as seen, e.g., in Figure 6(c), but this is far from the only approach.
-Shrinking, which we also investigated, might give a more realistic set of ${\theta}^{\prime}_{j}$’s, but it also requires an extra parameter.
-For the examples presented here shrinking gave similar results to cropping, and are therefore not included.
-An alternative to the expected value of the sample maximum is the upper limit of the 95% CI.
-The latter is more in line with common interpretation of the uncertainty of observed results in research, but which to choose will depend on the aim of the analysis.
-We demonstrate the impact in the Obesity Risk example, reported in Table 3.
-Not every $\max(\hat{\Theta})$ is a result of multiplicity, and we demonstrate this in the Cassava Leaf example.
-5.1 Estimating $\theta_{SOTA}$
-5.1.1 Multi-Class Prediction of Obesity Risk
-Multi-Class Prediction of Obesity Risk is a competition in the 2024 Kaggle Playground Series, using a synthetic data set generated from the Obesity or CVD risk data set (Palechor and de la Hoz Manotas, 2023).
-The test sample size was $n=13,840$, and the number of participating teams were $m=3,498$ (excluding teams with accuracy lower than chance).
-The maximum sample estimate was $\max(\hat{\Theta})=0.9116$.
-Figure 6(a) shows the sample estimates, $\hat{\theta}_{j}$’s, zoomed in at the $3,081$ classifiers with sample estimates above $0.88$.
-The 95% CI for $\hat{\theta}=0.9116$ is $(0.9067,0.9162)$ and is shown as the red line in Figure 6(a), the lower bound enveloping $920$ of the sample estimates.
-This should raise suspicion that the right tail seen in Figure 6(a) can be the product of multiplicity.
-(a) Sample estimates
-(b) A simulated realisation from a)
-(c) Cropped sample estimates
-(d) A simulated realisation from c)
-Figure 6: The competition observations and simulated realisation of classifiers with corresponding probabilities of correct classification.
-The red line is the 95% CI for $\hat{\theta}^{\,\max}$ in (a).
-To demonstrate the impact of ignoring multiplicity, let ${\theta}^{\prime}_{1},\ldots,{\theta}^{\prime}_{m}$ be equal to the sample estimates and calculate the expected value of the sample maximum, $E(\max(\hat{\Theta}^{\prime}))$.
-We use the set-up from Section 4.3 with correlation coefficient $\rho_{0}=0.6$, and $P(Y_{0})=\theta_{0}=\max({\Theta}^{\prime})$.
-Note that sample estimates with low values are excluded in accordance with Eq.
-9, but their influence on the expected maximum is negligible.
-A realisation is shown in Figure 6(b).
-The upper tail, whose tip represents the maximum sample estimate, is stretched according to the variances of the $Y_{j}$s.
-The estimated expected sample maximum is $E(\max(\hat{{\Theta}}^{\prime}))=0.9159$, far above the sample maximum in the competition.
-The 95% CI is reported in Table 3.
-The lower bound exceeds $\max(\hat{\Theta})$, and we can conclude that the sample maximum is an unlikely candidate for SOTA.
-The CI is calculated by $B=1,000$ bootstrap samples of size $m$ drawn from the sample estimates, and then the simulation as described above is conducted for each bootstrap sample.
-Instead, we can find a set $\Theta^{\prime}$ that results in realisations, $\hat{\Theta}^{\prime}$, similar to the competition sample estimates in terms of sample maximum, and let $\max(\Theta^{\prime})$ be the SOTA candidate.
-To obtain a set $\Theta^{\prime}$ where $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$, we crop the competition sample estimates, as seen in Figure 6(c), and where $E(\max(\hat{\Theta}^{\prime}))$ is calculated from simulations as described above.
-Figure 6(d) shows one realisation of the cropped distribution, where $\max(\hat{\Theta}^{\prime})$ is close to $\max(\hat{\Theta})$.
-By cropping at $0.9062$ we obtain $E(\max(\hat{\Theta}^{\prime}))=\max(\hat{\Theta})$, and this serves as a new candidate for the SOTA.
-It can be argued that the new SOTA candidate is at risk of being a optimistic estimate, since the competition observations are from only one test sample.
-The upper bound of the 95% CI can mitigate such a risk, and the competition sample estimates are cropped to fit that criterion.
-The CI is calculated by bootstrapping, as described above.
-The results are shown in Table 3.
-Estimating ${\theta}_{SOTA}$
-Approach $\max(\Theta^{\prime})$ 95% CI $E(\max(\hat{\Theta}^{\prime}))$ teams
-no mult. $\max(\hat{\Theta})$ (0.9067, 0.9162) - -
-$\Theta^{\prime}=\hat{\Theta}$ $\max(\hat{\Theta})$ (0.9128, 0.9191) 0.9159 -
-cropping 0.9062 (0.9084, 0.9148) $\max(\hat{\Theta})$ 1134
-0.9026 (0.9050, $\max(\hat{\Theta})$) 0.9083 1844
-Table 3: The 95% CI and the expected value for sample maximum, $\max(\hat{\Theta}^{\prime})$, for different SOTA estimation strategies, where SOTA is defined as $\max(\Theta^{\prime})$.
-The competition sample maximum is $\max(\hat{\Theta})=0.9116$.
-The bootstrap standard deviations are in the order of magnitude $10^{\text{-}5}$-$10^{\text{-}4}$, and are therefore not displayed.
+Table 3: The expected value and standard deviation for the SOTA estimator with varying test set size, number of classifiers and probability of success.
+Row 1 is the example used in Fig.
+2.
+The subsequent rows correspond to the intersection between the gray vertical lines and the curves in Fig.
+4.
+Estimating ${\theta}_{SOTA}$ by cropping
+Approach $\max(\Theta^{\prime})$ 95% CI $\text{E}\max(\hat{\Theta}^{\prime})$ teams
+cropping 0.9063 (0.9084, 0.9148) $\operatorname{\hat{\theta}_{\max}}$ 1134
+0.9026 (0.9050, $\operatorname{\hat{\theta}_{\max}}$) 0.9083 1844
+Table 4: The 95% CI and the expected value for sample maximum for SOTA estimation by cropping.
+This corresponds to the shrinking in Table 2.
+The standard deviations are in the order of magnitude $10^{\text{-}5}$-$10^{\text{-}4}$, and are therefore not displayed.
 The rightmost column shows number of classifiers with sample performance above the estimated SOTA.
-Table 3 demonstrates that taking the effect of multiplicity into account changes how results are perceived dramatically.
-The new SOTA value is $0.9062$, and even though its reduction from best performance is only about $0.005$, the effect it has on who we consider performing at SOTA-level is dramatic: around one third of the classifiers have sample performance above the new SOTA.
-The more conservative approach using the upper bound of the 95% CI to estimate SOTA gives an even more dramatic result, and either way the impact is substantial.
-5.1.2 Cassava Leaf Disease
-In this example, the maximum sample estimate is unlikely to be the product of multiplicity.
-For The Cassava Leaf Disease Classification competition, the task was to classify images of cassava leaves.
-The training set consisted of $21,367$ labelled images of five classes, and the test set of approximately $15,000$ images.
-$3,752$ teams participated (excluding those with sample performance lower than chance).
-The three top performances appear as outliers (0.9132, 0.9043 and 0.9028) as seen in Figure 7(a).
-Several explanations for the outlying performances of the three top teams have been put forward, and the interested reader can take a deep-dive in the comment section of the competition.
-Since this is not the topic of the present work, we will not engage in the discussion here.
-The 95% CI of $\hat{\theta}=0.9132$ is $(0.9086,0.9177)$, no other sample performance is within this interval.
-By performing simulations without cropping, we get $E(\max(\hat{\theta}^{\prime}))=\max(\hat{\Theta})$, and although the upper tail is stretched, its tip does not surpass $\max(\hat{\Theta})$.
-(a) The Kaggle performances
-(b) One realisation.
-Figure 7: Cassava Leaf sample performances and one realisation.
-This analysis demonstrates that the suggested approach for SOTA estimation is robust in the sense that when multiplicity is not a likely explanation to the sample maximum value, the SOTA value is not altered by this approach.
-5.2 Estimating $AUC_{SOTA}$
-5.2.1 Melanoma Classification
-We will use the Melanoma Classification competition from 2020 as an illustrative example of $AUC_{SOTA}$ estimation because it follows many of the common structures of public challenges.
-The results are presented in Section 5.2.3.
-The SIIM-ISIC Melanoma Classification competition consisted of more than $40,000$ images of skin lesions from various sites, split into training ($33,126$), validation ($3,295$) and test set ($7,687$).
-The data set is highly imbalanced, with less than $2\%$ melanomas, which reflects the reality in many cancer detection situations.
-A full description can be found in Rotemberg et al.
-(2021).
-The size of the test set is smaller than the recommendations of, e.g., Roelofs et al.
-(2019), but reflects the reality of benchmark datasets and public challenges, as seen in, e.g., the overview of Willemink et al.
-(2020).
-We want to emphasise that the shortcomings of this competition are not unique, and that is precisely why it serves as a useful example.
-Participants were allowed to use additional data for training if the data was shared on the competition web site.
-The prize money was set to $10,000$ USD for the best-performing method, measured as AUC on the hold-out test set.
-Figure 8 shows the results.
-Participants could get intermediate results from the validation set, published on a public leaderboard.
-Figure 8: AUCs for the Melanoma Classification competition, calculated on a hold-out test set, cropped from below for better visualisation.
-The team ‘All Data Are Ext’, consisting of three Kaggle Grandmasters, won the competition, achieving an AUC of $0.9490$ on the test set, and their method and strategy is explained in Ha et al.
-(2020).
-Although we show that the AUC is an optimistic estimate for the melanoma detection SOTA, we do not question their rank in the challenge.
-Quite the contrary; the winning team chose many of the verified strategies from Section 2.
-They used additional data from previous years’ competitions, and they also did data augmentation.
-They chose ensemble learning as a strategy to avoid overfitting, a strategy implicitly supported by Talagrand (1996).
-They also trained on multiple classes, a strategy that Feldman et al.
-(2019) showed will slow down the overfitting.
-In the following sections, we introduce a strategy for estimating $AUC_{SOTA}$, using simulated examples and the Melanoma Challenge for illustration.
-5.2.2 Simulation of AUC
-In a binary classification problem, the calculation of AUC is based on scores of random variables from the positive and the negative class.
-The AUC is mathematically equivalent to the probability that a random variable from the positive class will be ranked higher, in terms of score, than a random variable from the positive class (Hanley and McNeil, 1982).
-Let $score(\cdot)$ be a function that assigns a score to a random variable.
-Let $S_{+}=score(W_{+})$, where $W_{+}$ is a random variable from the positive class, and $S_{-}=score(W_{-})$, where $W_{-}$ is a random variable from the negative class.
-Then $\text{AUC}=P(S_{-}<S_{+})=P(S_{-}-S_{+}<0)$, which is equivalent to the cdf at value 00.
-To simulate classifier scores according to a given AUC $=a$, we need to specify distributions for the variables $S_{-},S_{+}$ in such a way that $P(S_{-}-S_{+}<0)=a$.
-We use Normal distributions for illustration.
-Let $S_{+}$ and $S_{-}$ have the distributions
-$S_{+}$ $\sim\mathcal{N}(\mu_{+},\sigma^{2}_{+})$ (10)
-$S_{-}$ $\sim\mathcal{N}(\mu_{-},\sigma^{2}_{-}),$ (11)
-where $\mathcal{N}(\mu,\sigma^{2})$ is the Normal distribution with mean $\mu$ and variance $\sigma^{2}$. We then have that
-$\text{AUC}=\Phi\left(\frac{0-(\mu_{-}-\mu_{+})}{\sqrt{\sigma_{-}^{2}+\sigma_{+}^{2}}}\right),$ (12)
-where $\Phi$ is the cdf of the standard Normal distribution.
-Specifying any three of the four parameters above decides the last one for a chosen AUC $=a$.
-Figure 9(a) shows an example where AUC = 0.90.
-The distribution of classifier scores, $S$, is a mixture distribution of the $S_{+}$ and the $S_{-}$ distributions, where the weights, $\pi$ and $1-\pi$, correspond to the probability of selecting a positive or a negative random sample, and has pdf
-$\pi f_{+}(s)+(1-\pi)f_{-}(s),$ (13)
-where $f_{+}(s)$ and $f_{-}(s)$ are the pdfs of the positive and the negative class, respectively. Expected value and variance of $S$ are
-$\mu$ $=\pi\mu_{+}+(1-\pi)\mu_{-},\text{ and}$ (14)
-$\sigma^{2}$ $=\pi(\sigma_{+}^{2}+\mu_{+}^{2})+(1-\pi)(\sigma_{-}^{2}+\mu_{-}^{2})-\mu^{2}.$ (15)
-Figure 9(a) shows an example of Normally distributed negative and positive scores and an AUC of 0.90.
-The rug shows an example draw with $n=3000$ and $\pi=0.017$, corresponding to the class balance in the Melanoma challenge.
-(a) Score pdfs for a classifier with AUC= $0.90$; grey for negatives, black for positives. The rug shows an example draw from these with class balance $\pi=0.017$.
-(b) ROC curves for a classifier with AUC $=0.90$ across 100 realisations of (a).
-Figure 9: Example of what an AUC of $0.90$ might look like in practice.
-The right figure shows $100$ realizations of a ROC curves.
-The left figure shows the underlying score distributions that generated the curves in (b).
-The choice of distributions has an impact on the final SOTA estimation, as it affects the variance of the ROC curves.
-However, with high AUC, the shape has less impact, and the analyst’s choices are not suggestive of the overall conclusion.
-5.2.3 Simulation results for uncorrelated classifiers
-To demonstrate the impact that multiplicity can have in a public competition, we start with a simulated example similar to the one in Section 3.4.
-Consider a classification problem with a test set of size $n=3,000$, number of participating teams, $m=1,000$, where each classifier has a true underlying AUC of $a=0.90$, and the classifiers are independent.
-Whereas the distribution of the sample maximum for probability of success is independent of class balance as long as the probability of success is independent of class, this is not true for the sample maximum of AUC, where the variance increases with class imbalance.
-We have chosen the same class balance as in the Melanoma challenge, $\pi=0.017$, as this reflects real world examples.
-For simplicity we choose $\mu_{-}=0$ and $\sigma_{-}=\sigma_{+}=1$.
-We then solve for $\mu_{+}$ numerically.
-Now we can generate scores from the resulting Normal distribution.
-The 95% CI of a single classifier with $\hat{AUC}=0.90$ is (.$(.8558,0.9376)$, and gives a strong indication of what we can expect when $1,000$ teams apply their classifiers on a test set.
-Figure 10: Simulated distribution of the sample maximum AUC with parameters as in Figure 9. This is an analogous figure to Figure 2(b) in Section 3.4.
-Figure 10 shows the simulated distribution of the sample maximum AUC over $10,000$ repetitions of a competition where every classifier has a true AUC of $0.90$.
-The simulated expected value is $0.9562$, the standard deviation is $0.004459$, and the simulated 95% CI is $(0.9486,0.9662)$.
-Comparing this to the results in Section 3.4, we see that the bias is much larger for the sample maximum AUC.
-This is partly due to the large variance created by the class imbalance.
-Although the sample size is $n=3,000$, only $51$ of those belong to the positive class, creating large variance in the estimator.
-Again, it is clear that the observed maximum is an unlikely SOTA candidate, and to give a better estimate of the SOTA for the Melanoma challenge, we follow the same procedure as in Section 5.1: The competition sample estimates are cropped in such a manner that their realisations give an expected value of $0.9490$, the competition sample maximum.
-(a) Truncated AUC distribution based on the Melanoma challenge scores. The part of the distribution removed by truncation is shaded in dark grey.
-(b) A realization from the truncated distribution of (a).
-Figure 11: Histogram of cropped competition AUC sample estimates. Here the state of the art is an AUC of 0.8975 which leads to the expected sample maximum AUC of 0.9490.
-Figure 11(a) shows one bootstrap sample from the kaggle observations cropped at $0.8975$.
-Number of classifiers is set to $m=3,173$, which is as many as there are classifiers in the melanoma challenge with AUC $>0.5$.
-Figure 11(b) shows a corresponding realisation, where the maximum is close to $0.9490$.
-There are $2,022$ teams with performance above the new SOTA candidate.
-The discrepancy between the sample maximum and the new SOTA candidate is big, and have big implications.
-From an AUC close to $95\%$ to just under $90\%$ is a substantial drop in performance in a cancer detection setting.
-In addition, two-thirds of the classifiers have performances above the new SOTA estimate.
-As was shown in Section 4, illustrated in Figure 4(b), the bias is reduced for correlated classifiers, so the analysis with independence assumption is an overestimation of the bias.
-In this challenge, we can assume multiple sources of dependency: Access to the same training set, models pre-trained on the same data, similar models, etc.
-However, the wide CI for a single classifier tells the story of an estimator with large variance, and hence, the bias of the sample maximum will be substantial.
-5.2.4 Correlation between classifiers in the AUC simulations
-For the moment we have not extended our AUC simulations to include correlations, but it should be possible to adopt a similar setup as in Sections 4.2 and 4.3: each classifier produces scores that correlate to those of a reference classifier.
-Let $S_{0^{+}}$ and $S_{0^{-}}$ be the scores of the reference classifier, for random variables from the positive and negative class, respectively.
-Let $\rho_{+}=corr(S_{0^{+}},S_{j^{+}})$ be the correlation between the reference classifier scores $S_{0^{+}}$ and the scores of classifier $j$, $S_{j^{+}}$ for the positive class.
-If they both follow Normal distributions as outlined above, then the bivariate distribution is
-$\begin{pmatrix}S_{0^{+}}\\ S_{j^{+}}\end{pmatrix}\sim\mathcal{N}\left(\begin{pmatrix}\mu_{0^{+}}\\ \mu_{j^{+}}\end{pmatrix},\begin{pmatrix}\sigma_{0^{+}}^{2}&\rho\sigma_{0^{+}}\sigma_{j^{+}}\\ \rho\sigma_{0^{+}}\sigma_{j^{+}}&\sigma_{j^{+}}^{2}\end{pmatrix}\right),$
-and equivalently for the negative class.
-However, the joint distribution of $(S_{0},S_{j})$ is not uniquely defined by the correlation, and additional assumptions must therefore be made.
-We leave this as future work.
-6 Discussion
-We have addressed the mechanisms of multiplicity and how they affect the bias of the SOTA estimator.
-When probability of success is used as a performance measure, the exact distribution of the sample maximum estimator can be provided in the simpler cases, and known analyses methods can be engaged.
-The bias in SOTA is substantial in our examples, and possibly hinders new methods to attract interest from the scientific community.
-We have presented the simple case with an analytical solution, where the probabilities of correct prediction are known.
-We have also presented three case-studies from Kaggle competitions, where we demonstrate how a different SOTA candidate than the observed maximum can be estimated, and how it impacts the perceived results.
-This work shows that for some cases, an unbiased SOTA estimate can be calculated fairly easily through bootstrapping, and can serve as a better reference point when introducing new methods to the scientific community.
-In other cases, the SOTA estimation is more complicated.